### v1

# Dynamics of “Spontaneous” Topic Changes in Next Token Prediction with Self-Attention

Abstract
Human cognition can spontaneously shift conversation topics, often triggered by emotional or contextual signals. In contrast, self-attention-based language models depend on structured statistical cues from input tokens for next-token prediction, lacking this spontaneity. Motivated by this distinction, we investigate the factors that influence the next-token prediction to change the topic of the input sequence. We define concepts of topic continuity, ambiguous sequences, and change of topic, based on defining a topic as a set of token priority graphs (TPGs). Using a simplified single-layer self-attention architecture, we derive analytical characterizations of topic changes. Specifically, we demonstrate that (1) the model maintains the priority order of tokens related to the input topic, (2) a topic change occurs only if lower-priority tokens outnumber all higher-priority tokens of the input topic, and (3) unlike human cognition, longer context lengths and overlapping topics reduce the likelihood of spontaneous redirection. These insights highlight differences between human cognition and self-attention-based models in navigating topic changes and underscore the challenges in designing conversational AI capable of handling “spontaneous” conversations more naturally. To our knowledge, this is the first work to address these questions in such close relation to human conversation and thought.
1 Introduction
In human cognition, topic changes often occur sporadically, with little apparent structure or forethought: an spontaneous shift in focus during a conversation, a sudden leap between ideas when brainstorming, or an unexpected redirection in storytelling. These seemingly chaotic shifts are intrinsic to the human way of processing and communicating information (Christoff and Fox,, 2018; Mills et al.,, 2020). For instance, consider a casual conversation where someone is discussing their favorite book, and then they suddenly mention an unrelated memory of a family gathering. This abrupt change may be due to an emotional connection, such as recalling reading that book during a family vacation, where sensory details like the scent of the ocean or the warmth of the sun trigger a vivid memory. The hippocampus helps link these sensory cues to specific memories, resulting in an unexpected yet meaningful shift (Lisman et al.,, 2017). The prefrontal cortex then integrates these memories, emotions, and contextual information, allowing for a spontaneous redirection of conversation (Miller and Cohen,, 2001).
Self-attention mechanisms in language models rely on contextual signals in the input to perform topic shifts, with next-token prediction guided by the relevance of preceding tokens (Vaswani et al.,, 2017). Unlike humans, who transition fluidly and intuitively between topics, often driven by emotion or creativity, transformers follow a structured, statistical approach. They remain on topic unless explicit cues signal a change, reflecting a key distinction between the somehow rigid processing of language models and the dynamic transitions of human cognition (Brown et al.,, 2020; Clark et al.,, 2019; Devlin et al.,, 2019; Bisk et al.,, 2020).
It is also important to differentiate between topic changes and hallucinations. Hallucinations involve generating incorrect or fabricated information without clear contextual basis, often resulting in false or illogical content (Ji et al.,, 2023; Maynez et al.,, 2020). In contrast to topic changes, hallucinations are not deliberate or contextually driven shifts in discourse. More interestingly, some hallucinations may still appear to maintain topic consistency by staying loosely connected to the previous topic, yet they fail to introduce a meaningful or valid change. Our focus is on understanding when and why self-attention models predict next tokens from a different topic of the input, examining the conditions under which an input sequence from one topic leads to a prediction in another.
Recent advancements in the related field have substantially deepened our understanding of self-attention mechanisms. The Transformer model, introduced by Vaswani et al., (2017), revolutionized natural language processing by using self-attention to determine the importance of each token in relation to others. This innovation enables Transformers to focus on relevant context when predicting the next token, significantly enhancing capabilities natural language understanding and generation (Brown et al.,, 2020; Devlin et al.,, 2019). Efforts to improve next token prediction have prioritized optimizing the efficiency of the self-attention mechanism, achieving empirical success across tasks such as machine translation, question answering, and text summarization. Additionally, connections between self-attention and support vector machines (SVM) have offered valuable insights into optimization strategies, and mechanisms within the next token prediction (Ataee Tarzanagh et al.,, 2023; Tarzanagh et al.,, 2023; Li et al., 2024b, ). These advancements underscore the transformative role of attention-based architectures in token prediction tasks. A recent study by Li et al., (2023) highlights that in mixed topic inputs, transformers achieve higher pairwise attention between same-topic words compared to different-topic words. However, our approach aligns more closely with human cognitive processes, where the input is typically centered on a single topic, ensuring that discussions or conversations remain within a consistent context or theme.
Despite these advancements, our understanding of how LLMs manage topic changes in comparison to human cognition remains limited. Human cognition seamlessly navigates such transitions through spontaneous thought and contextual awareness (Christoff and Fox,, 2018). Investigating this distinction could provide valuable insights into the strengths and limitations of language models in mimicking human-like conversation. Bridging this gap is essential for enhancing language models, particularly in their ability to adapt to and manage dynamic topic changes–a process closely aligned with the flow of spontaneous thought in human cognition. Specifically, we address the following research questions:
  • What underlying mechanism drives topic changes during next token prediction in self-attention architectures?
  • Which factors shape the dynamics of topic changes in next token prediction?
To answer the above questions, it is essential first to explore how self-attention allocates attention to tokens for next token prediction with respect to the input topic. Understanding how attention weights are allocated offers key insights into the prioritization of tokens by self-attention mechanisms, particularly when models are trained on datasets spanning multiple topics. This fundamental understanding will pave the way for uncovering the role of attention patterns in driving “spontaneous” topic changes. To the best of our knowledge, this is the first work to explore these questions in direct relation with human conversation and thought.
Summary of findings
For simplicity, we elaborate our results using a two-topic scenario, but it is straightforward to extend them to multiple topics. Imagine an oracle that is an expert on Topic A, capable of following any conversation within that topic while staying true to its context. Now, suppose the oracle gains knowledge of Topic B and is following a conversation about Topic A. Will the oracle’s responses remain firmly within Topic A, or will the influence of the knowledge of Topic B cause the conversation to drift? This analogy encapsulates the problem we address: understanding when and why attention models might preserve a topic or change to another “spontaneously". Specifically, we make the following contributions:
  1. 1.
    Preservation of input topic priorities. Using a controlled sandbox, we demonstrate in Theorem 3 that attention models trained on mixed-topic datasets maintain the priorities of tokens associated with the original topic of an input sequence (Topic A in our analogy).
  2. 2.
    Changing topics triggered by token frequency. In Theorem 4, we show that only if a lower-priority token appears more frequently than all higher-priority tokens of Topic A, the oracle’s responses may begin to reflect Topic B, signaling a shift.
  3. 3.
    Impact of sequence length and topic overlap. Theorem 5 establishes that longer input sequences decrease the likelihood of changing topics. Furthermore, overlapping topics act as a stabilizing factor, not increasing the frequency of “spontaneous” topic changes. This finding represents a key contribution to the development of human-like AI conversational models. Unlike human dialogue, where extended discussions often encourage spontaneous topic shifts and overlapping topics promote cognitive connections that facilitate smooth transitions, our results highlight a fundamental limitation of attention-based models in replicating these human-like dynamics.
Figure 1: $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ (left) are token-priority graphs (TPGs) of the last token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively. $\mathcal{G}^{(4)}_{ab}$ (right) is the combination of TPG for $\mathbb{T}_{a}$ and TPG for $\mathbb{T}_{b}$.
Figure 2: Illustration of topic continuity, ambiguous sequence, and change of topic (see Definition 3,  4, and 5 for formal definitions). For each input sequence derived from $\mathbb{T}_{a}$, we employ a self-attention model trained on $\mathbb{T}_{a}$ and another model trained on the mixed-topics dataset $\mathbb{T}_{ab}$ to predict the next tokens $\hat{y}_{{\bf w}_{a}}$ and $\hat{y}_{{\bf w}_{ab}}$, respectively (orange). $\widehat{\mathcal{G}}^{(4)}_{ab}$ is the highest probability SCC (see Definition 1) in mixed-topics setting and $\widehat{\mathcal{G}}^{(4)}_{a}$ is the highest probability SCC in $\mathbb{T}_{a}$, The input sequence is classified as topic continuity if $\hat{y}_{{\bf w}_{ab}}\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$ (top). The input sequence is ambiguous if $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$, but $\widehat{\mathcal{G}}^{(4)}_{a}(X)\subset\widehat{\mathcal{G}}^{(4)}_{ab}(X)$ (middle). ${\bf W}_{ab}$ is considered to change of topic, if $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$ and $\widehat{\mathcal{G}}^{(4)}_{a}(X)\not\subset\widehat{\mathcal{G}}^{(4)}_{ab}(X)$ (bottom). The numeric details are provided in Figure 6 in the appendix.
Here, we outline the structural overview of our approach. Following the framework of token-priority graphs (TPGs) introduced by Li et al., 2024b , we start with defining a topic as a set of TPGs (see Sec. 3). TPGs represent the priority order of tokens conditioned on the last token of the input sequence. For simplicity, we consider two different topics, Topic A and Topic B, characterized by two different sets of TPGs. We demonstrate the preservation of input topic priorities in a model trained with mixed-topic datasets (see Sec. 4). Additionally, we introduce three concepts to categorize the prediction outputs: topic continuity, ambiguous sequence, and change of topic. We subsequently establish the conditions under which a self-attention model induces a topic change, showing that the probability of topic changes does not increase with longer input sequences or the presence of overlapping topics (see Sec. 5). To formalize this approach, we begin with the structured problem setup.
2 Problem setup
Next topic prediction with self-attention model.
In line with the approach presented by Ataee Tarzanagh et al., (2023) and Li et al., 2024b , we frame the next-token prediction task as a multi-class classification problem. Given a vocabulary of size $K$ with an embedding matrix ${\bf E}=[{\bf e}_{1}\ {\bf e}_{2}\ \cdots\ {\bf e}_{K}]^{\top}\in\mathbb{R}^{K\times d}$, we aim to predict the next token ID $y\in[K]$ based on an input sequence ${\bf X}=[{\bf x}_{1}\ {\bf x}_{2}\ \cdots\ {\bf x}_{T}]^{\top}\in\mathbb{R}^{T\times d}$ with ${\bf x}_{i}\in{\bf E}$ for all $i\in[T]$. The training dataset, denoted as
$\text{DSET}=\{({\bf X}_{i},y_{i})\in\mathbb{R}^{T_{i}\times d}\times[K]\}_{i=1}^{n},$
contains sequences of varying lengths $T_{i}$. In our notation ${\bf x}$ is the embedding vector corresponding to the token ID $x$, this is ${\bf x}={\bf e}_{x}$. For prediction, we utilize a single-layer self-attention model with a combined key-query weight matrix ${\bf W}\in\mathbb{R}^{d\times d}$ and identity value matrix as in Ataee Tarzanagh et al., (2023). The self-attention embedding output
$f_{\bf W}({\bf X})={\bf X}^{\top}\mathbb{S}({\bf X}{\bf W}\bar{{\bf x}}),$ (output)
where $\mathbb{S}(\cdot)$ is the softmax operation and $\bar{{\bf x}}:={\bf x}_{T}$, serves as a weighted representation of the tokens, allowing for context-sensitive prediction of ${\bf y}$ based on the final input token. Let $\ell:\mathbb{R}\to\mathbb{R}$ be a loss function. For the training set DSET, we aim to minimize empirical risk (ERM) with:
$L({\bf W})=\frac{1}{n}\sum_{i=1}^{n}\ell({\bf c}_{y_{i}}^{\top}{\bf X}_{i}^{\top}\mathbb{S}({\bf X}_{i}{\bf W}\bar{{\bf x}}_{i})).$ (ERM)
We assume a well pre-trained classification head matrix ${\bf C}=[{\bf c}_{1}\ {\bf c}_{2}\ \cdots\ {\bf c}_{K}]^{\top}\in\mathbb{R}^{K\times d}$. Each classification head ${\bf c}_{k}\in\mathbb{R}^{d}$ is fixed and bounded for all $k\in[K]$. Starting from ${\bf W}^{(0)}\in\mathbb{R}^{d\times d}$ with step size $\eta>0$, for $\tau\geq 0$ we optimize ${\bf W}$ with a gradient descent algorithm
${\bf W}^{(\tau+1)}={\bf W}^{(\tau)}-\eta\nabla L({\bf W}^{(\tau)}).$ (Algo-GD)
We keep the same first two assumptions from Li et al., 2024b :
Assumption 1.
$\forall y,k\in[K],k\neq y,{\bf c}_{y}^{\top}{\bf e}_{y}=1$ and ${\bf c}_{y}^{\top}{\bf e}_{k}=0$
Assumption 2.
For any $({\bf X},y)\in\text{DSET}$, the token ${\bf e}_{y}$ is contained in the input sequence ${\bf X}$.
Assumption 1 represents a variation of the weight-tying approach commonly used in language models (Press and Wolf,, 2016; Vaswani et al.,, 2017). Once training is complete, for a new input sequence ${\bf X}$, and a model characterized by $\bf W$, we predict the next token ID $\hat{y}_{\bf w}$ based on greedy decoding the probabilities from the softmax of the classification output
$\hat{y}_{\bf w}\in\arg\max_{k\in[K]}\left[\mathbb{S}\left({\bf C}f_{\bf W}({\bf X})\right)\right]_{k}.$ (1)
Token-priority graph and global convergence of the self-attention model
Li et al., 2024b defined a token-priority graph (TPG) as a directed graph with nodes representing tokens in the vocabulary. $\text{DSET}^{(k)}$ is a subset of sequences from DSET with the same last token is ${\bf e}_{k}=\bar{\bf x}$. They defined TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ such that every $\mathcal{G}^{(k)}$ is a directed graph where for every sequence $({\bf X},{\bf y})\in\text{DSET}^{(k)}$ a directed edge is added from ${\bf y}$ to every token ${\bf x}\in{\bf X}$. TPGs are further divided into strongly-connected components (SCCs), which capture subsets of tokens where priority relations are cyclic, indicating comparable token importance. For tokens within two different SCCs, strict priority orders emerge, helping the model to differentiate between tokens when learning next-token predictions. We use the same notation as Li et al., 2024b , given a directed graph $\mathcal{G}$, for $i,j\in[K]$ such that $i\neq j$:
  • $i\in\mathcal{G}$ denotes that the node $i$ belongs to $\mathcal{G}$.
  • $(i\Rightarrow j)\in\mathcal{G}$ denotes that the directed path $(i\rightarrow j)$ is present in $\mathcal{G}$ but $j\rightarrow i$ is not.
  • $(i\asymp j)\in\mathcal{G}$ means that both nodes $i$ and $j$ are in the same strongly connected component (SCC) of $\mathcal{G}$ (there exists both a path $i\rightarrow j$ and $j\rightarrow i$).
For any two distinct nodes $i,j$ in the same TPG, they either satisfy $(i\Rightarrow j)$, $(j\Rightarrow i)$ or $(i\asymp j)$. Nodes in each $\mathcal{G}^{(k)}$ represent indices in $[K]$, and SCC structure supports the self-attention mechanism’s ability to assign priority within sequences based on the conditioning last token. Theorem 2 of Li et al., 2024b proved that under Assumptions 1 and 2, the self-attention model learned through Algo-GD converges to the solution of the following Support Vector Machine (SVM) defined by the TPGs of the underlying dataset DSET
${\bf W}^{\text{svm}}=\arg\min_{\bf W}\|{\bf W}\|_{F}$ (Graph-SVM)
[…]
Here a condensed version of the theorem:
Theorem 1 (from Li et al., 2024b ).
Consider dataset DSET and suppose Assumptions 1 and 2 hold. Set loss function as $\ell(u)=-\log(u)$. Starting Algo-GD from any ${\bf W}(0)$ with constant size $\eta$, if ${\bf W}^{\text{svm}}\neq{\bf 0}$,
$\tilde{\bf W}=\lim_{\tau\to\infty}\frac{{\bf W}(\tau)}{\|{\bf W}(\tau)\|_{F}}=\frac{{\bf W}^{\text{svm}}}{\|{\bf W}^{\text{svm}}\|_{F}}$ (2)
This convergence implies that the model predicts the next token based on priorities obtained from the SCCs within the TPG relevant to the last token of the input sequence. We add here another reasonable assumption that prevents the probabilities in Equation 1 to be equal for improbable numerical reasons, and we present our first lemma.
Assumption 3.
For any $({\bf X},y)\in\text{DSET}$, $\exists i,j\in[T]\text{ and }u,v\in\mathbb{Z}$ such that $u\left[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})\right]_{i}=v\left[\mathbb{S}({\bf X}{\tilde{\bf W}}\bar{{\bf x}})\right]_{j}\iff u=v\text{ and }\left[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})\right]_{i}=\left[\mathbb{S}({\bf X}{\tilde{\bf W}}\bar{{\bf x}})\right]_{j}$.
Lemma 2.
Suppose conditions from Theorem 1 and Assumption 3 hold. Consider an input sequence ${\bf X}$ from DSET(k) and corresponding TPG $\mathcal{G}^{(k)}$, $\forall i,j\in[K]$ we have $\left[\mathbb{S}\left({\bf C}f_{\bf W}({\bf X})\right)\right]_{i}=\left[\mathbb{S}\left({\bf C}f_{\tilde{\bf W}}({\bf X})\right)\right]_{j}\iff(x_{i}\asymp x_{j})\in\mathcal{G}^{(k)}$.
This means that the tokens that maximize the probability in Equation 1 are all within the same SCC leading to the following definition:
Definition 1 (highest probability SCC).
Consider an input sequence ${\bf X}$ from DSET(k) and corresponding TPG $\mathcal{G}^{(k)}$. We define $\widehat{\mathcal{G}}^{(k)}({\bf X})\in\mathcal{G}^{(k)}$ as the highest probability SCC for ${\bf X}$ in $\mathcal{G}^{(k)}$ such that $\forall{\bf x}\in\widehat{\mathcal{G}}^{(k)}({\bf X})$ we have $\left[\mathbb{S}\left({\bf C}f_{\tilde{\bf W}}({\bf X})\right)\right]_{x}=\left\|\mathbb{S}\left({\bf C}f_{\tilde{\bf W}}({\bf X})\right)\right\|_{\infty}$.
3 Defining Topics
In order to answer our research questions regarding the dynamics of topic changes we need to define the concept of a topic. In the previous settings, a dataset DSET generates TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$, but, conversely, an existing set of TPGs can generate DSET, akin to generating data from a particular topic, which leads to the following definitions.
Definition 2 (topic).
A topic $\mathbb{T}$ is a set of TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$. Given topic $\mathbb{T}$ defined by TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$, input sequence ${\bf X}$ belongs to $\mathbb{T}$ if $\forall{\bf x}\in{\bf X},x\in\mathcal{G}^{(\bar{x})}$. A sequence $({\bf X},y)$ is within $\mathbb{T}$ if ${\bf X}$ belongs to $\mathbb{T}$ and $\forall{\bf x}\in{\bf X},\ (y\Rightarrow x)\in\mathcal{G}^{(\bar{x})}$.
We can define a topic as a set of TPGs because it inherently captures the relationships and contextual dependencies between tokens, offering a structured representation of a topic’s semantics. This means that we can use a topic $\mathbb{T}$ to generate DSET by sampling sequences $({\bf X},y)$ within $\mathbb{T}$. Given the finite number of edges, a DSET can be generated from $\mathbb{T}$ such that it can reconstruct the exact TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ that define $\mathbb{T}$, following the construction method in Li et al., 2024b . This leads to the following reasonable assumption:
Assumption 4.
A DSET generated from any topic $\mathbb{T}$ defined by $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ exactly reconstructs back the TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$.
This assumption enables the application of the results from Li et al., 2024b , with the concepts of topics and TPGs being used interchangeably.
Definition 3 (topic continuity).
Given an input sequence ${\bf X}$ that belongs to $\mathbb{T}$, a weight matrix ${\bf W}$ is said to keep topic $\mathbb{T}$ for the input sequence ${\bf X}$ if $\hat{y}_{\bf W}\in\widehat{\mathcal{G}}^{(k)}({\bf X})$.
This definition establishes the criteria for a next predicted token to preserve the topic of the input sequence. Notice that if we have two topics, $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, we can generate two different datasets DSETa and DSETb. It is clear that $\tilde{\bf W}_{a}$ [^1] Notation: The subscripts of weights and objects correspond to the associated topic. For instance $\tilde{\bf W}_{a}$ denotes the weights defined in Equation 2, obtained from DSETa, which pertains to topic $\mathbb{T}_{a}$. trained only with DSETa will always keep topic $\mathbb{T}_{a}$. But we could also obtain $\tilde{\bf W}_{ab}$ with a dataset combining DSETa and DSETb as training sets. The central question is whether $\tilde{\bf W}_{ab}$ keeps topic $\mathbb{T}_{a}$, given an input sequence ${\bf X}$ that belongs to $\mathbb{T}_{a}$, or if it instead predicts tokens that prompt a topic change. This can be understood through an analogy: consider an oracle only proficient in Topic A, capable of following conversations in the context of Topic A. If this oracle gains knowledge of Topic B, will it continue to follow the conversation within Topic A’s context when the conversation is within Topic A, or will its response begin to exhibit a shift to Topic B?
4 Attention within mixed topics
Let’s first understand the way in which attention models assign priority to tokens within mixed-topic setting. For simplicity, we elaborate our results using a two-topic scenario, but it is straightforward to extend all of our theoretical results on multiple topics. Notice the self-attention embedding output is a linear combination of $\bf X$ given by $\mathbb{S}({\bf XW\bar{x}})$. The embeddings in $\bf X$ corresponding to the highest entries in the vector $\mathbb{S}({\bf XW\bar{x}})$ will receive higher priority in their importance to predict the next token, therefore we can hypothesize that models in which $\mathbb{S}({\bf XW\bar{x}})$ are ordered in a similar way will predict similar next tokens. This idea leads to our first main result which considers this situation within a mixed topics setting:
Theorem 3.
Consider datasets DSETa and DSETb from topics $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively. Let DSETab be the union of DSETa and DSETb. Suppose Assumptions 1, 2, 3 and 4 hold. Set loss function as $\ell(u)=-\log(u)$. Starting Algo-GD from any initial point with constant size $\eta$ and if ${\bf W}_{a}^{\text{svm}}\neq{\bf 0}$ and ${\bf W}_{ab}^{\text{svm}}\neq{\bf 0}$; for a given sequence ${\bf X}$ that belongs to $\mathbb{T}_{a}$, we have that $\tilde{\bf W}_{ab}$ preserves the attention priority of $\mathbb{T}_{a}$ on input ${\bf X}$. This is $\forall i,j\in[T]$:
[…]
This implies that for an input sequence $\mathbf{X}$, a model trained in a mixed-topic setting will maintain the priority of the topic to which $\mathbf{X}$ belongs. Consequently, the attention will be allocated in the same order as if the model had been trained exclusively on the original topic of $\mathbf{X}$. The input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$, as shown in Figure 2 (top), is generated from $\mathbb{T}_{a}$. The predicted next token $\hat{y}_{{\bf w}_{ab}}$ is ${\bf e}_{5}$ and the highest probability SCC in mixed-topics is $\widehat{\mathcal{G}}^{(4)}_{ab}({\bf X})=\{{\bf e}_{5}\}$. Since $\hat{y}_{{\bf w}_{ab}}$ belongs to ${\widehat{\mathcal{G}}}^{(4)}_{ab}({\bf X})$, ${\bf W}_{ab}$ for input sequence $\bf X$ is considered as topic continuity, based on the Definition 3.
The only assumption about ${\bf X}$ on Theorem 3 is that it belongs to $\mathbb{T}_{a}$. However, if $\mathbf{X}$ belongs to $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, the priority will be preserved within both topics. Additionally, strict equality in the attention priority is maintained, but strict inequalities may not be when combining both datasets. In such cases, the union of their TPGs can form new SCCs, disrupting strict priority. As illustrated in Figure 1, we present $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ as the TPGs corresponding to last input token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively. In $\mathcal{G}^{(4)}_{a}$, the token priority is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}>{\bf e}_{4}$. Once combining $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$, the priority order in the mixed-topics is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}={\bf e}_{4}$, based on $\mathcal{G}^{(4)}_{ab}$. The strict equality ${\bf e}_{1}={\bf e}_{2}$ from $\mathcal{G}^{(4)}_{a}$ is maintained in $\mathcal{G}^{(4)}_{ab}$, but the strict inequality ${\bf e}_{2}>{\bf e}_{4}$ is replaced with ${\bf e}_{2}={\bf e}_{4}$ in mixed-topics, forming the new SCC, $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\}$, in $\mathcal{G}^{(4)}_{ab}$.
Figure 3: Convergence of attention priority similarity for $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue) and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (orange).
To further illustrate Theorem 3 we define the attention priority similarity of weights ${\bf W}^{\prime}$ relative to ${\bf W}$ for a sequence ${\bf X}$ as: $R_{{\bf W},{\bf W}^{\prime}}({\bf X})=$
$\frac{1}{T-1}\sum_{j=1}^{T-1}g\left(\left[\mathbb{S}\left({\bf X}{\bf W}^{\prime}\bar{\bf x}\right)\right]_{i_{j}}-\left[\mathbb{S}\left({\bf X}{\bf W}^{\prime}\bar{\bf x}\right)\right]_{i_{j+1}}\right),$
where $i_{1},\cdots,i_{T}$ is a permutation of $1,\cdots,T$ such that $\left[\mathbb{S}\left({\bf X}{\bf W}\bar{\bf x}\right)\right]_{i_{1}}\geq\cdots\geq\left[\mathbb{S}\left({\bf X}{\bf W}\bar{\bf x}\right)\right]_{i_{T}}$, and
$g(w)=\begin{cases}1,&\text{if $w\geq 0$,}\\ \frac{1}{e^{-w}},&\text{otherwise.}\end{cases}$
The attention priority similarity quantifies how well the weights ${\bf W}^{\prime}$ preserve the attention priority of the weights ${\bf W}$. A value of 1 indicates that the priority is fully preserved. Using this metric, we conduct experiments, with results in Figure 3. We generate embeddings with $K=10$ and $d=16$, and randomly construct TPGs for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$. Using these TPGs, we randomly generate $\text{DSET}_{a}$ and $\text{DSET}_{b}$. We compute $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$, $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ and $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ using the same procedure as Li et al., 2024b . We generate test sequences ${\bf Z}$ within $\mathbb{T}_{a}$, and we calculate the attention priority similarity of $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to both $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$. We repeat this process for multiple TPGs and input sequences (details in Appendix). Figure 3 clearly demonstrates that the similarity converges to 1 after $\tau=8000$ iterations when evaluated relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue line), but fails to converge relative to $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ (orange line). These observations align with the results of Theorem 3.
5 Explaining topic shifts
The formation of new SCCs when combining datasets suggests that the highest priority SCC for some input sequences may increase in size in this new setting. This also suggests that topic shifts may arise from ambiguity within an input sequence rather than a straightforward change in topic. In our analogy on the oracle, gaining knowledge of both Topic A and Topic B might cause a conversation to be naturally followed within Topic A or also outside Topic A. Thus, we introduce the following definition to characterize this phenomenon:
Definition 4 (ambiguous sequence).
Given DSETa and DSETb generated from two different topics $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$. Denote $\mathbb{T}_{ab}$ as the combined topic defined by a combination of DSETa and DSETb. A sequence $\bf X$ that belongs to $\mathbb{T}_{a}$ is ambiguous in $\mathbb{T}_{ab}$ with respect to $\mathbb{T}_{a}$ if $\tilde{\bf W}_{ab}$ does not keep topic $\mathbb{T}_{a}$ for ${\bf X}$, but $\widehat{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})\subset\widehat{\mathcal{G}}_{ab}^{(\bar{x})}({\bf X})$.
Definition 4 defines an ambiguous sequence as one where the highest-probability next-token predictions include tokens from both within and outside the input topic, reflecting natural ambiguity from overlapping topics. For instance, the input sequence ${\bf X}=[{\bf e}_{1},{\bf e}_{4},{\bf e}_{1},{\bf e}_{4}]^{\top}$ in Figure 2 (middle) belongs to both $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$. $\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$ is $\{{\bf e}_{1}\}$, as depicted in $\mathcal{G}_{a}^{(4)}$ from Figure 1 (left) and $\widehat{\mathcal{G}}_{ab}^{(4)}({\bf X})$ is $\{{\bf e}_{1},{\bf e}_{4}\}$, as shown in $\mathcal{G}_{ab}^{(4)}$ from Figure 1 (right). $\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$ is the subset of $\widehat{\mathcal{G}}_{ab}^{(4)}({\bf X})$, though $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$. We can argue that the next token predicted from an ambiguous sequence cannot be considered as a topic change, as it lacks the clear trigger phenomenon observed in human cognition. To address this, we propose a formal definition for a topic change:
Definition 5 (change of topic).
Given DSETa and DSETb generated from two different topics $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, and a sequence $\bf X$ that belongs to $\mathbb{T}_{a}$. The weight matrix $\tilde{\bf W}_{ab}$ changes topic $\mathbb{T}_{a}$ for sequence ${\bf X}$ if $\tilde{\bf W}_{ab}$ does not keep topic $\mathbb{T}_{a}$ for ${\bf X}$ and $\bf X$ is not ambiguous in $\mathbb{T}_{ab}$ with respect to $\mathbb{T}_{a}$.
Following the above definition, ${\bf W}_{ab}$ for the input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{4},{\bf e}_{4},{\bf e}_{4}]^{\top}$, from Figure 2 (bottom), can be considered as change of topic.
Building on the formal definitions of topic continuity, ambiguous sequences, and topic changes, we now present a necessary condition for a sequence to induce a topic change. This is achieved by introducing our final definition, grounded in the highest-priority SCC as determined by the order in the attention layer. Now, our second key result.
Definition 6 (highest priority SCC).
Consider a sequence ${\bf X}$ that belongs to $\mathbb{T}$. We define $\dot{\mathcal{G}}^{(\bar{x})}({\bf X})\in\mathcal{G}^{(\bar{x})}$ as the highest priority SCC for ${\bf X}$ in $\mathcal{G}^{(\bar{x})}$ such that $\forall{\bf x}_{i}\in\dot{\mathcal{G}}^{(\bar{x})}({\bf X})$ and ${\bf x}_{j}\in\mathcal{G}^{(\bar{x})}$ we have $(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}^{(\bar{x})}$.
Theorem 4.
Under the same settings and assumptions on datasets and training in Theorem 3, let ${\bf X}$ be a sequence that belongs to $\mathbb{T}_{a}$. If $\tilde{\bf W}_{ab}$ changes topic $\mathbb{T}_{a}$ for $\bf X$ then $\exists x_{j}\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$ such that $\forall x_{i}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$, the number of times ${\bf x}_{j}$ appears in $\bf X$ is greater than the number of times ${\bf x}_{i}$ appears in $\bf X$.
Theorem 4 implies that, for a given sequence X from $\mathbb{T}_{a}$ and its corresponding TPG, a necessary condition for a topic change is the presence of a lower-priority token that appears more frequently than any of the higher-priority tokens. This can be intuitively understood through our analogy: if the oracle is following a conversation on Topic A but the conversation contains repeated components with lower importance in Topic A, its knowledge of Topic B may steer the response toward Topic B, thereby initiating a shift away from Topic A. A natural question arises: what do these findings imply in practice? Specifically, how does the probability of change of topic behaves as the input sequence length increases or as the ambiguity between topics grows? The following theorem provides insights into these dynamics.
Theorem 5.
Under same settings and assumptions on datasets and training in Theorem 3, let $\bf X$ be a sequence that belongs to $\mathbb{T}_{a}$ with no repeated tokens, and $l$ be the number of elements in $\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})$. Let ${\bf X}^{\prime}=[{\bf x}_{1}^{\prime}\ {\bf x}_{2}^{\prime}\ \cdots\ {\bf x}_{T}^{\prime}]^{\top}$ be a random sequence of iid random tokens sampled from $\bf X$ such that for a fixed $p$, $p=\min_{x\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})}\mathbb{P}\left({\bf x}_{i}^{\prime}={\bf x}\right)$. We have:
  1. 1.
    If $p>\max_{x\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})}\mathbb{P}\left({\bf x}_{i}^{\prime}={\bf x}\right)$ then
    $\mathbb{P}(\tilde{\bf W}_{ab}$ changes topic $\mathbb{T}_{a}$ for sequence ${\bf X}^{\prime})\to 0$ as $T\to\infty.$
  2. 2.
    If $l$ increases then the probability that $\exists x_{j}^{\prime}\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$ such that $\forall x_{i}^{\prime}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$, the number of times ${\bf x}_{j}^{\prime}$ appears in $\bf X^{\prime}$ is greater than the number of times ${\bf x}_{i}^{\prime}$ appears in $\bf X^{\prime}$ does not increase.
There are two implications of this theorem. First, as the input sequence length increases sufficiently, the likelihood of topic changes vanishes. Second, increasing $l$ leads to possibly higher overlap between topics, meaning that the probability of satisfying the necessary conditions for a topic change decreases, effectively creating a bound on the frequency of topic changes. In practice, consider the oracle analogy: if the oracle is following a sufficiently long conversation on a specific topic, it becomes exceedingly unlikely to shift topics. Similarly, as topics A and B become more interconnected, this increased ambiguity does not lead to more topic changes; rather, it may reduce their occurrence. This contrasts with human cognition, where longer conversations and greater inter-connectivity of knowledge increase the likelihood of spontaneous shifts.
(a) Effect of increased sequences.
(b) Effect of increased overlap.
Figure 4: The proportion of keep topic, ambiguity, and change topic as (a) the test sequence length increases and (b) the number of edges in generated TPGs increase.
To illustrate Theorem 5, we conduct simulations following the same implementation settings as described in Section 4. We approximate $\tilde{\mathbf{W}}_{a}$ and $\tilde{\mathbf{W}}_{ab}$ as the results obtained after $\tau=8000$ iterations of Algo-GD. We quantify the proportion of test sequences in which $\tilde{{\bf W}}_{ab}$ keeps $\mathbb{T}_{a}$ (keep topic), proportion of ambiguous sequences in $\mathbb{T}_{ab}$ (ambiguity) and proportion of times in which $\tilde{{\bf W}}_{ab}$ changes topic (change topic). First, we explore the effect of longer sequences by varying the length $T$ of the test sequences $\bf Z$. We increase $T$ from 4 to 512. Figure 4 illustrates how the proportion of change topic decreases as $T$ increases. Second, we investigate the effect of overlapping topics by generating TPGs with an increased number of edges $L$. Intuitively, a higher $L$ results in greater overlap between TPGs from different topics. We vary $L$ from 4 to 18. Figure 4 demonstrates that as $L$ increases, ambiguity increases, while proportion of change topic decreases. These two findings contrast with expectations derived from human cognition but align with the result of Theorem 5. Lastly, among the 85000 test sequences generated for these experiments, 99.98% satisfy Theorem 4 (i.e., topic changes occur when a low-priority token appears more frequently than high-priority tokens). The remaining 0.02% mismatched cases were solely due to minor approximation discrepancies in the attention softmax. These results validate Theorem 4 (see specific experimental details in the Appendix).
Related work
Training and generalization of Transformer. (1) Properties of Self-attention and Softmax. The self-attention mechanism employs the softmax function to selectively emphasize different parts of the input. Gu et al., (2024), Goodfellow et al., (2016), and Deng et al., (2023) underscore the pivotal role of the softmax function in shaping attention distributions, influencing how models process and prioritize information within input sequences. Bombari and Mondelli, (2024) examined the word sensitivity of attention layers, revealing that softmax-based attention-layers are adept at capturing the significance of individual words. Despite these strengths, recent researches have highlighted limitations of the softmax function (Saratchandran et al.,, 2024; Deng et al.,, 2023). (2) Optimization in attention-based models. Additionally, recent researches interpret Transformer models as kernel machines, akin to support vector machines (SVMs), with self-attention layers performing maximum margin separation in the token space (Tarzanagh et al.,, 2023; Ataee Tarzanagh et al.,, 2023; Li et al., 2024b, ; Julistiono et al.,, 2024). (3) Chain-of-Thought and In-Context Learning. Moreover, transformers exhibit remarkable abilities in generalization through in-context learning (ICL), where models effectively learn from contextual cues during inference (Brown et al.,, 2020; Xie et al.,, 2022; Olsson et al.,, 2022). Chain-of-Thought (CoT) prompting (Wei et al.,, 2022; Zhou et al.,, 2022; Shao et al.,, 2023; Li et al., 2024a, ) enhances this by breaking down reasoning processes into intermediate steps, highlighting the emergent reasoning abilities of transformers. (4) Improvement efficiency of transformers. Recent advancements aim to improve the computational efficiency of transformers (Kitaev et al.,, 2020; Choromanski et al.,, 2021; Sukhbaatar et al.,, 2019; Wang et al.,, 2020), ensuring their viability for large-scale deployment while maintaining or enhancing their representational capabilities.
Next token prediction in LLMs. (1) Theoretical and architectural innovations. Shannon, (1951)’s foundational work laid the groundwork for estimating the predictability of natural language sequences, providing a basis for subsequent advances in language modeling. Recent studies have expanded our understanding of how LLMs anticipate future tokens from internal hidden states, offering valuable insights into the efficiency and effectiveness of transformer-based architectures (He and Su,, 2024; Pal et al.,, 2023; Shlegeris et al.,, 2024). Despite their impressive predictive capabilities, these models face fundamental limitations. For instance, Bachmann and Nagarajan, (2024) highlights the shortcomings of teacher-forced training, emphasizing how this approach can fail and suggesting strategies to improve model robustness. (2) Efficiency and Optimization. Goyal et al., (2024) introduces a novel method that incorporates a deliberate computation step before output generation, enhancing reasoning capabilities. Additionally, Gloeckle et al., (2024) advocates for multi-token prediction, which significantly improves both efficiency and speed.
Self-Attention and topic dynamics. Advancements in the field has significantly advanced our understanding of how self-attention mechanisms in transformers manage evolving semantic contexts, with works have explored various aspects of topic modeling, including dynamic topic structures (Miyamoto et al.,, 2023), hierarchical relationships (Lin et al.,, 2024), topic-aware attention (Panwar et al.,, 2021), and mechanistic analysis of topic structure (Li et al.,, 2023). While these works provide insights into managing static and hierarchical topic structures, our work focuses on the topic changes during next-token prediction within the given input sequences from a specific topic.
6 Discussion
Our findings on a simple single-layer attention model provide valuable clues about the fundamental differences in conversational dynamics between humans and machines. Unlike humans, who naturally shift topics through curiosity or contextual association, the model adheres to predictive patterns, mirroring LLMs’ tendency to stay focused on a single thread. This suggests that spontaneous topic shifts, a hallmark of human conversation, may require mechanisms beyond predictive pattern recognition. These insights raise philosophical questions about the nature of human-like interaction and highlight the need for adaptability and creativity in advancing conversational AI.
Future research will focus on how real-world LLMs perform changing topics, where the interplay of diverse and evolving contexts presents unique challenges. It would also be interesting to investigate the connection between LLM and human cognition in spontaneously changing topics. Additionally, the recurring patterns of topic retention identified in this work open new pathways for applications. Harnessing these behavioral signatures could enhance changing topics detection in conversational AI, enrich content generation by anticipating and adapting to topic transitions, and refine model interpretability. Pursuing these directions could significantly advance both the theoretical landscape and practical utility of LLMs.
Appendix A Technical Proofs
A.1 Proofs of lemmas
Proof of Lemma 2
Let ${\bf a}=\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})$.
${\bf C}f_{\tilde{\bf W}}({\bf X})$ $={\bf C}\left({\bf X}^{\top}\mathbb{S}\left({\bf X}\tilde{\bf W}\bar{{\bf x}}\right)\right)$ (3)
$={\bf C}\left({\bf X}^{\top}{\bf a}\right)$ (4)
$=\begin{bmatrix}\sum_{i=1}^{T}a_{i}\,({\bf c}_{1}^{\top}\cdot{\bf x}_{i})\\ \sum_{i=1}^{T}a_{i}\,({\bf c}_{2}^{\top}\cdot{\bf x}_{i})\\ \vdots\\ \sum_{i=1}^{T}a_{i}\,({\bf c}_{K}^{\top}\cdot{\bf x}_{i})\end{bmatrix}.$ (5)
Let $k_{i}$ be the number of times token ${\bf x}_{i}$ appears in ${\bf X}$. Then,
$\left[{\bf C}f_{\tilde{\bf W}}({\bf X})\right]_{x_{i}}=k_{i}a_{i}.$
From Assumption 3 we have that
$\left[{\bf C}f_{\tilde{\bf W}}({\bf X})\right]_{x_{i}}=\left[{\bf C}f_{\tilde{\bf W}}({\bf X})\right]_{x_{j}}$ $\iff a_{i}=a_{j}$ (7)
$\iff(x_{i}\asymp x_{j})\in\mathcal{G}^{(\bar{x})}\text{ or }x_{i}=x_{j}.$ (8)
If $x_{i}\neq x_{j}$ then $x_{i}$ and $x_{j}$ are in the same SCC.∎
Lemma 6.
For an input sequence ${\bf X}$ that belongs to $\mathbb{T}$ and $i,j\in[T]$,
  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{i}\asymp x_{j})\in\mathcal{G}^{(\bar{x})}$ or $i=j$.
  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}^{(\bar{x})}$.
  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$.
Proof.
Since $\bf X$ belongs to $\mathbb{T}$, $\forall{\bf x}\in{\bf X}$ we have $x\in\mathcal{G}^{(\bar{x})}$, therefore from the construction of TPGs by Li et al., 2024b , for every ${\bf x}_{i},{\bf x}_{j}\in{\bf X}$ we have one of the these relationships: $(x_{i}\Rightarrow x_{j})$, $(x_{j}\Rightarrow x_{i})$, $(x_{i}\asymp x_{j})$ or $x_{i}=x_{j}$. From the constraints in Algo-GD:
  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}=0\iff(x_{j}\asymp x_{i})\in\mathcal{G}^{(\bar{x})}\text{ or }i=j$.
  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}>1\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}^{(\bar{x})}$.
  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}<1\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$.
Lemma 7.
For an input sequence ${\bf X}$ that belongs to $\mathbb{T}$,
$\dot{\mathcal{G}}^{(\bar{x})}({\bf X})=\left\{x_{i}\ |\ [\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}=\|\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})\|_{\infty}\right\}.$
Proof.
Let $G=\{x_{i}\ |\ [\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}=\|\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})\|_{\infty}\}$. From Lemma 6 $\forall x_{i},x_{j}\in G$, $(x_{i}\asymp x_{j})\in\mathcal{G}^{(\bar{x})}$. Therefore all elements in $G$ belong to the same SCC. Also from Lemma 6, $\forall x_{i}\in G,x_{j}\not\in G$ we have $(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$. This means that every element in $G$ has the highest priority among tokens in $\bf X$ concluding our proof. ∎
A.2 Proof of Theorem 3
From construction, $\forall k\in[K]$, $\mathcal{G}_{a}^{(k)}\subseteq\mathcal{G}_{ab}^{(k)}$. This means that $\forall{\bf x}_{i},{\bf x}_{j}\in{\bf X}$, we have:
  • if $(x_{i}\asymp x_{j})\in\mathcal{G}_{a}^{(\bar{x})}$ then $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$
  • if $(x_{j}\Rightarrow x_{i})\in\mathcal{G}_{a}^{(\bar{x})}$ then $(x_{j}\Rightarrow x_{i})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$
  • if $(x_{i}\Rightarrow x_{j})\in\mathcal{G}_{a}^{(\bar{x})}$ then $(x_{i}\Rightarrow x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$
Combining with Lemma 6:
  • $[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{i}\asymp x_{j})\in\mathcal{G}_{a}^{(\bar{x})}$ or $i=j$, then $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $i=j\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
  • $[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}_{a}^{(\bar{x})}$ then $(x_{j}\Rightarrow x_{i})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\leq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
  • $[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}_{a}^{(\bar{x})}$ then $(x_{i}\Rightarrow x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\geq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
A.3 Proof of Theorem 4
Let ${\bf a}=\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})$ and ${\bf b}=\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})$. Without loss of generality, suppose ${\bf a}$ is in decreasing order $a_{1}\geq\cdots\geq a_{T}$. From Theorem 3, we also have $b_{1}\geq\cdots\geq b_{T}$. Let $k_{i}$ be the number of times token ${\bf x}_{i}$ appears in ${\bf X}$. Following an analogous procedure as in Lemma 2 we get
$\left[{\bf C}f_{\tilde{\bf W}_{a}(\tau)}({\bf X})\right]_{x_{i}}=k_{i}a_{i}$ (9)
$\left[{\bf C}f_{\tilde{\bf W}_{ab}(\tau)}({\bf X})\right]_{x_{i}}=k_{i}b_{i}$ (10)
We will proof the contrapositive: If $\exists x_{i}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})$ such that $k_{i}\geq k_{j}$ for all $j\in[K]$, then there is no change of topic, so $\tilde{\bf W}_{ab}$ keeps topic $\mathbb{T}_{a}$ for input sequence $\bf X$, or $\bf X$ is ambiguous in $\mathbb{T}_{ab}$ with respect to $\mathbb{T}_{a}$.
From Lemma 7, if $x_{i}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})$, we have $a_{i}\geq a_{j}$ for all $j\in[K]$. Suppose $\exists x_{i}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})$ such that $k_{i}\geq k_{j}$ for all $j\in[K]$, we have that $k_{i}a_{i}\geq k_{j}a_{j}$ for all $j\in[K]$ then $x_{i}\in\widehat{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})$. Analogously since $b_{i}\geq b_{j}$, $x_{i}\in\widehat{\mathcal{G}}_{ab}^{(\bar{x})}({\bf X})$. If $\exists x_{l}\in\widehat{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})$ with $x_{l}\neq x_{i}$ then $k_{l}a_{l}\geq k_{j}a_{j}$ for all $j\in[K]$, then $k_{l}a_{l}=k_{i}a_{i}$. Therefore from Assumption 3 and Lemma 7, $(x_{l}\asymp x_{i})\in\mathcal{G}_{a}^{(\bar{x})}$. Analogously $(x_{l}\asymp x_{i})\in\mathcal{G}_{ab}^{(\bar{x})}$. This means that if $\exists x_{i}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})$ such that $k_{i}\geq k_{j}$ for all $j\in[K]$, then $\widehat{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})\subseteq\widehat{\mathcal{G}}_{ab}^{(\bar{x})}({\bf X})$. Then $\tilde{\bf W}_{ab}$ keeps topic $\mathbb{T}_{a}$ for input sequence $\bf X$, or $\bf X$ is ambiguous in $\mathbb{T}_{ab}$ with respect to $\mathbb{T}_{a}$.∎
A.4 Proof of Theorem 5
  1. 1.
    This is a direct consequence from the law of large numbers. If $T\to\infty$ the proportion of each token will match the probability. Since $p>\max_{{\bf x}\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})}\mathbb{P}\left({\bf x}_{i}^{\prime}={\bf x}\right)$, then the probability that $\exists x_{j}^{\prime}\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$ such that $\forall x_{i}^{\prime}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$, the number of times ${\bf x}_{j}^{\prime}$ appears in $\bf X^{\prime}$ is greater than the number of times ${\bf x}_{i}^{\prime}$ appears in $\bf X^{\prime}$ will go to zero, and therefore the probability of change topics will do it also.
  2. 2.
    Without loss of generality suppose $\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})=\{x_{1},x_{2},\cdots,x_{l}\}$. Clearly if we prove the result assuming $\forall x\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X}),\ p=\mathbb{P}\left({\bf x}_{i}^{\prime}={\bf x}\right)$, we will also have it for the more general case $p=\min_{{\bf x}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})}\mathbb{P}\left({\bf x}_{i}^{\prime}={\bf x}\right)$.
    Let ${\bf X}_{l}^{\prime}=[{\bf x}_{1,l}^{\prime}\ {\bf x}_{2,l}^{\prime}\ \cdots\ {\bf x}_{T,l}^{\prime}]^{\top}$ be a random sequences generated as described in the theorem, where the size of $\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})$ is $l$. Let $k_{i,l}$ be the number of times ${\bf x}_{i}$ is selected in ${\bf X}_{l}^{\prime}$ Let $A_{l}=\max_{1\leq i\leq l}k_{i,l}$ and $B_{l}=\max_{l+1\leq i\leq K}k_{i,l}$. Let $P(l)=\mathbb{P}(B_{l}>A_{l})$. We want to prove $P(l+1)\leq P(l)$. We construct a coupling between ${\bf X}_{l}^{\prime}$ and ${\bf X}_{l+1}^{\prime}$ by performing $T$ independent trials. For each trial $i$ we generate a uniform random variable $U_{i}$ in $[0,1]$ and we choose tokens in ${\bf X}_{l}^{\prime}$ and ${\bf X}_{l+1}^{\prime}$ in this way:
    • If $U_{i}\leq pl$ both the selected tokens $x_{i,l}^{\prime}$ and $x_{i,l+1}^{\prime}$ are in $\{x_{1},x_{2},\cdots,x_{l}\}$.
    • If $pl<U_{i}\leq p(l+1)$, we select $x_{i,l}^{\prime}=x_{l+1}$ if $U_{i}\leq pl+q$ or $x_{i,l}^{\prime}=x_{l+2}$ otherwise, and we select $x_{i,l+1}^{\prime}=x_{l+1}$; where $q$ is the probability of choosing $x_{l+1}$ in ${\bf X}_{l}^{\prime}$. Since $p>q$, there is an interval where $x_{i,l}=x_{l+2}$ but $x_{i,l+1}=x_{l+1}$.
    • If $U_{i}>p(l+1)$, then both the selected tokens $x_{i,l}^{\prime}$ and $x_{i,l+1}^{\prime}$ are in $\{x_{l+2},x_{2},\cdots,x_{l}\}$. Notice that the probability of choosing $x_{i}$ in ${\bf X}_{l+1}^{\prime}$ for $i\geq l+2$ decreases because $p$ is constant.
    From the previous coupling we have that $k_{i,l}=k_{i,l+1}$ for $1\leq i\leq l$, $k_{l+1,l}\leq k_{l+1,l+1}$ for $i=l+1$, and $k_{i,l}\geq k_{i,l+1}$ for $i\geq l+2$. This means that $A_{l+1}=\max(A_{l},k_{l+1,l+1})\geq A_{l}$ and $B_{l+1}=\max_{l+2\leq i\leq K}k_{i,l+1}\leq B_{l}$. Therefore $P(l+1)=\mathbb{P}(B_{l+1}>A_{l+1})\leq\mathbb{P}(B_{l}>A_{l})=P(l)$.
Appendix B Detailed simulation studies
B.1 Simulation process
Theoretical TPGs generation. For each token $e_{k}$, $L$ edges are randomly selected to construct the theoretical TPG $\mathcal{G}^{(k)}_{theor}$ for $e_{k}$ , ensuring that $e_{k}$ is involved, as either a source or destination node. Based on these selected edges, we add additional edges from $e_{k}$ to all other tokens included in $L$ edges, thereby ensuring that all tokens in $\mathcal{G}^{k}_{theor}$ can be reached by $e_{k}$. Thus, we obtain the theoretical TPGs $\{\mathcal{G}^{(k)}_{a,theor}\}^{K}_{k=1}$ for Topic A . This process is repeated to generate another group of theoretical TPGs $\{\mathcal{G}^{(k)}_{b,theor}\}^{K}_{k=1}$ for the Topic B. Let $\mathcal{G}^{(k)}_{a,theor}$ and $\mathcal{G}^{(k)}_{b,theor}$ combine for each $k$, we obtain the theoretical TPGs for topics combinations $\{\mathcal{G}^{(k)}_{ab,theor}\}^{K}_{k=1}$.
Training Dataset Generation. Generate training datasets DSETa and DSETb based on $\{\mathcal{G}^{(k)}_{a,theor}\}^{K}_{k=1}$ and $\{\mathcal{G}^{(k)}_{b,theor}\}^{K}_{k=1}$, respectively. For each input sequence in DSET, the sequence length $T_{train}$ is 4, which means ${\bf X}=[{\bf x}_{1}\ {\bf x}_{2}\ \cdots\ {\bf x}_{T_{train}}]^{\top}\in\mathbb{R}^{T_{train}\times d}$ with ${\bf x}_{i}$ from ${\bf E}=[{\bf e}_{1},{\bf e}_{2},...{\bf e}_{K}]^{\top}$. ${\bf e}_{k}$ is randomly selected as the last token and other tokens (other input tokens and the next predicted token) are chosen based on $\mathcal{G}^{(k)}_{theor}$. Specifically, the next token ${\bf e}_{T_{train}+1}$ is determined by sampling with the weighted probability in $\mathcal{G}^{(k)}_{theor}$, where the weight for each token corresponds to the number of outcoming edges. Given Assumption 2, we randomly choose the position of the next token in the input sequence. Then, the remaining input tokens are randomly selected from tokens connected by incoming edges from ${\bf e}_{k}$ (i.e., ${\bf e}_{k}\rightarrow{\bf e}_{i}$) and placed in the random position within the input sequence. This process is repeated $n$ times to generate training data for each topic respectively. Empirical TPGs $\{\mathcal{G}^{(k)}_{a,empir}\}^{K}_{k=1}$ and $\{\mathcal{G}^{(k)}_{b,empir}\}^{K}_{k=1}$ are derived from the training datasets DSETa and DSETb. According to Assumption 4, the empirical TPGs $\{\mathcal{G}^{(k)}_{empir}\}^{K}_{k=1}$ are expected to be identical to the theoretical TPGs $\{\mathcal{G}^{(k)}_{theor}\}^{K}_{k=1}$ for each topic. The experiments are conducted with $5000$ instances, with each parameter setting evaluated over 50 epochs, consisting of 100 sequences per epoch.
Trained attention weights. We employ a single-layer attention mechanism implemented in PyTorch. The model is trained using the SGD optimizer with a learning rate $\eta=0.01$ for $8000$ iterations. The training of attention weights is divided into two stages for each instance: (1) computing ${\bf W}^{\text{svm}}$ for each topic  [^2] Note: ${\bf W}^{\text{svm}}={\bf 0}$ means the number of SCCs is $1$ for $\mathcal{G}^{k},\forall k\in[K]$. During the simulation, we proceed to the next instance when ${\bf W}^{\text{svm}}={\bf 0}$ until reaching a total of $100$ instances.; (2) get $\bf{W}(\tau)$ at each iteration for each topic. In Stage (1), prior to using the CVXPY package to get ${\bf W}^{\text{svm}}$, SCCs are identified for each TPG derived from the using Tarjan’s algorithm. Afterward, ${\bf W}^{\text{svm}}$ is normalized to ensure consistency in subsequence computations. In Stage (2), the $MLayerAttn$ function encapsulates the architecture of a single-layer attention-based model. The training function is then used to optimize the attention weights by minimizing the loss defined in ERM. Finally, the correlation between ${\bf W}^{\text{svm}}$ and $\bf{W}(\tau)$ is calculated using the dot product.
Next token prediction. To differentiate the input sequence length of the testing data from that of the training data, we introduce $T_{test}$. TPGs based on the training dataset DSETa are utilized to generate test datasets consisting of 100 sequences from $\mathbb{T}_{a}$ per epoch. Specifically, the last token ${\bf x}_{T_{test}}$ of the test input sequence is randomly selected from $K$ tokens (i.e. ${\bf x}_{T_{test}}={\bf e}_{k}$) and the remaining input tokens are randomly chosen based on the SCCs of $\mathcal{G}^{k}_{a}$, where tokens with higher priority are assigned greater weights. For instance, in $\mathcal{G}^{4}_{a}$, tokens ${\bf e}_{1},{\bf e}_{2},{\bf e}_{4}$ are captured with the priority order ${\bf e}_{1}={\bf e}_{2}>{\bf e}_{4}$. The weights assigned to input tokens ${\bf e}_{1},{\bf e}_{2}$, and ${\bf e}_{4}$ are $0.4,0.4$, and $0.2$, respectively. It reflects that ${\bf e}_{1}$ and ${\bf e}_{2}$ are in the same higher-priority SCC, thus having greater weights compared to ${\bf e}_{4}$. Intuitively, tokens within the same SCC are more likely to co-occur than those from different SCCs. This approach enables the generated test input sequences to mimic real word relationships and reflect their contextual groupings. Following the generation of the test dataset from $\mathbb{T}_{a}$, the next tokens $\hat{y}_{\bf w_{a}}$ and $\hat{y}_{\bf w_{ab}}$ are predicted by Equation 1, with ${\bf W}_{a}$ and ${\bf W}_{ab}$ obtained from the last iteration. To reduce the potential numerical issues in the outputs, $\mathbb{S}({\bf X}{\bf W}\bar{{\bf x}})$ is rounded to three decimals, ensuring that tokens within the same SCC yield consistent softmax outputs.
B.2 Experiments results
In this section, we provide empirical evidence to support our findings outlined in Theorem 3, Theorem 4, and Theorem 5, offering detailed insights into the behavior of self-attention mechanisms for changing topics.
B.2.1 Simulation in Section 4
For Figure 3 we used $K=10,L=4,d=16,\text{and }T_{test}=T_{train}=4$, we also showed point-wise confidence intervals of rank similarities on all 5000 instances for each value of $\tau$.
B.2.2 Simulation in Section 5
For Figure 4 (left), we predict next tokens for 5000 test sequences from $\mathbb{T}_{a}$ with $T_{test}=\{4,8,16,24,32,64,128,256,512\}$, while fixing $L=4$, $d=16$, $T_{train}=4$, and $K=10$. The proportion of each scenario with varying $T$ is illustrated in Table 1. For Figure 4 (right), we predict next tokens for 5000 test sequences (the sequence length is $T_{test}=20$) using models trained with $L=\{4,6,8,10,12,14,16,18\}$, $d=16$, $K=10$, and $T_{train}=4$. The proportion of each scenario with varying $L$ is illustrated in Table 2.
Table 1: The proportion of keep topic, ambiguous, and change of topic with varying $T_{test}=\{4,8,16,24,32,64,128,256,512\}$.
$T_{test}$ Keep(%) Ambiguous(%) Change(%)
4 98.60 $\pm$ 1.54 1.40 $\pm$ 1.54 0.00 $\pm$ 0.00
8 98.50 $\pm$ 1.47 0.96 $\pm$ 1.11 0.54 $\pm$ 0.76
16 98.06 $\pm$ 1.33 0.54 $\pm$ 0.76 1.40 $\pm$ 1.07
24 98.48 $\pm$ 1.31 0.26 $\pm$ 0.44 1.26 $\pm$ 1.10
32 98.84 $\pm$ 1.15 0.12 $\pm$ 0.33 1.04 $\pm$ 1.03
64 99.10 $\pm$ 1.07 0.04 $\pm$ 0.20 0.86 $\pm$ 1.05
128 99.64 $\pm$ 0.53 0.02 $\pm$ 0.14 0.34 $\pm$ 0.52
256 99.98 $\pm$ 0.14 0.00 $\pm$ 0.00 0.02 $\pm$ 0.14
512 100.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
Table 2: The proportion of keep topic, ambiguous, and change of topic with varying $L=\{4,6,8,10,12,14,16,18\}$.
L Keep(%) Ambiguous(%) Change(%)
4 98.22 $\pm$ 1.43 0.26 $\pm$ 0.60 1.52 $\pm$ 1.31
6 98.30 $\pm$ 1.37 0.50 $\pm$ 0.68 1.20 $\pm$ 1.11
8 98.18 $\pm$ 1.49 0.68 $\pm$ 0.68 1.14 $\pm$ 1.23
10 98.42 $\pm$ 1.25 0.76 $\pm$ 0.85 0.82 $\pm$ 1.02
12 98.14 $\pm$ 1.32 0.82 $\pm$ 0.92 1.04 $\pm$ 0.97
14 97.96 $\pm$ 1.44 1.24 $\pm$ 1.06 0.80 $\pm$ 0.86
16 98.28 $\pm$ 1.33 0.98 $\pm$ 0.91 0.74 $\pm$ 0.85
18 98.02 $\pm$ 1.58 1.26 $\pm$ 1.14 0.72 $\pm$ 0.86
B.2.3 Additional experiments
(a) $K=6,L=4$
(b) $K=10,L=4$
(c) $K=14,L=4$
(d) $K=10,L=8$
(e) $K=10,L=12$
(f) $K=10,L=16$
Figure 5: Convergence of $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$(blue), $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$(orange), and $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$(green) for varying $K$ and $L$, with fixed $T_{train}=4$ and $d=16$.
Building upon the convergence experiments in Li et al., 2024b , our work demonstrates that the correlation coefficients $\langle{\bf W}_{ab}(\tau),{\bf W}^{\text{svm}}_{ab}\rangle/\langle\|{\bf W}_{ab}(\tau)\|_{F},\|{\bf W}^{\text{svm}}_{ab}\|_{F}\rangle$ (green lines) in Figure 5, measured with varying $K=\{6,10,14\}$ and $L=\{8,12,16\}$, approach to 1. These results indicate that Theorem 1 extends beyond individual topics to also capture the convergence dynamics of mixed-topics scenarios, albeit with relatively slower convergence. In these experiments, we fix $T_{train}=4$ and $d=16$. Each point represents the average over 5000 randomly generated instances, trained with 8000 iterations. The shaded area around each line represents the $95\%$ confidence interval, computed over 50 epochs.
B.3 Numerical analysis of each scenario in Figure 2
(a) Topic continuity.
(b) Ambiguous sequence.
(c) Change of topic.
Figure 6: Numeric details for each scenario: (a) topic continuity, (b) ambiguous sequence, and (c) change of topic.
Figure 6 provides a numerical breakdown for each scenario in Figure 2. In Figure 6, each distinct color corresponds to a unique token within the input sequence $\bf X$, which consists of 4 tokens. ${\bf e}_{4}$ is the last token across all three input sequences. For each input sequence $\bf X$, we apply ${\bf W}_{a}(\tau)$ and ${\bf W}_{ab}(\tau)$ with $\tau=8000$ to predict the next token, yielding $\hat{y}_{{\bf W}_{a}}$ and $\hat{y}_{{\bf W}_{ab}}$, respectively.
Let $[\mathbb{S}({\bf X}{\bf W}_{a}(\tau){\bar{\bf x}})]_{i}=a_{i}$ and $[\mathbb{S}({\bf X}{\bf W}_{ab}(\tau){\bar{\bf x}})]_{i}=b_{i}$, for $i\in[T]$. Following Equation 9 and Equation 10, we compute $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{x_{i}}$ and $[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{x_{i}}$ to get the highest probability SCC and predict the next token for each input sequence.
Topic continuity. In Fig. 6(a), input sequence $\bf X$ consists of four unique tokens: ${\bf e}_{5}$, ${\bf e}_{1}$, ${\bf e}_{3}$, and ${\bf e}_{4}$. Based on $\mathcal{G}^{(4)}_{a}$ in Figure 1, the priority order of these tokens is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}>{\bf e}_{4}$, with corresponding $a_{i}$ values: $0.45>0.25>0.20>0.1$. Since $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.45$ is the largest, $\widehat{\mathcal{G}}_{a}^{(4)}=\{{\bf e}_{5}\}$ and $\hat{y}_{{\bf W}_{a}}={\bf e}_{5}$. In the mixed-topics scenario, ${\bf W}_{ab}$ preserves the attention priority but ${\bf e}_{4}$ and ${\bf e}_{1}$ have the same priority: ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{4}$, with corresponding $b_{i}$ values: $0.40>0.30>0.15=0.15$. Token ${\bf e}_{5}$ is still with the highest probability to be chosen, as $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.40$. Following the Definition 3, ${\bf W}_{ab}$ keeps topic for the the input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$.
Ambiguous sequence. Input sequence $\bf X$ in Fig. 6(b) has two unique tokens: ${\bf e}_{1}$ and ${\bf e}_{4}$. The priority order is ${\bf e}_{1}>{\bf e}_{4}$, following $\mathcal{G}^{(4)}_{a}$ in Figure 1. The corresponding values are $a_{1}=a_{3}=0.3$ and $a_{2}=a_{4}=0.2$. Then $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{1}}=2\times 0.30$ and $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.20$. Thus, $\widehat{\mathcal{G}}_{a}^{(4)}$ is $\{{\bf e}_{5}\}$ with the highest probability. ${\bf W}_{ab}$ makes ${\bf e}_{4}$ and ${\bf e}_{1}$ with the same priority, as indicated by $\mathcal{G}^{(4)}_{ab}$ in Figure 1. Both ${\bf e}_{1}$ and ${\bf e}_{4}$ are within the highest probability SCC, $\widehat{\mathcal{G}}_{ab}^{(4)}$, due to $[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{1}}=[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.25$. Although $\hat{y}_{{\bf W}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}$, $\widehat{\mathcal{G}}_{a}^{(4)}\in\widehat{\mathcal{G}}_{ab}^{(4)}$. Therefore, the sequence ${\bf X}=[{\bf e}_{1},{\bf e}_{4},{\bf e}_{1},{\bf e}_{4}]^{\top}$ is ambiguous, based on the Definition 4.
Change of topic. For the input sequence $\bf X$ in Fig. 6(c), the only two unique tokens, ${\bf e}_{5}$ and ${\bf e}_{4}$, are with the same priority order in both $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{ab}$ from Figure 1: ${\bf e}_{5}>{\bf e}_{1}$. With ${\bf W}_{a}(\tau)$ trained in $\mathbb{T}_{a}$, the token ${\bf e}_{5}$ has $a_{1}=0.70$ and the token ${\bf e}_{4}$ has $a_{2}=a_{3}=a_{4}=0.10$. Obviously, $1\times 0.70=[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}>[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{4}}=3\times 0.10$. Thus, $\widehat{\mathcal{G}}_{a}^{(4)}$ consists of ${\bf e}_{5}$. However, $\widehat{\mathcal{G}}_{ab}^{(4)}$ consists of ${\bf e}_{4}$ instead of ${\bf e}_{5}$, due to $1\times 0.40=[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{5}}<[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{4}}=3\times 0.20$. Since $\hat{y}_{{\bf W}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}$ and $\widehat{\mathcal{G}}_{a}^{(4)}\not\subset\widehat{\mathcal{G}}_{ab}^{(4)}$, ${\bf W}_{ab}$ changes topic for the input sequence $\bf X$. Moreover, we have $({\bf e}_{5}\Rightarrow{\bf e}_{i})\in\mathcal{G}^{(4)}_{ab}$ for $i\in[4]$, as shown in Figure 1. Thus, the highest priority SCC (Definition 1) in $\mathbb{T}_{ab}$ is $\dot{\mathcal{G}}^{(4)}_{ab}({\bf X})=\{{\bf e}_{5}\}$. In the input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{4},{\bf e}_{4},{\bf e}_{4}]^{\top}$, the lower-priority token ${\bf e}_{4}\not\in\dot{\mathcal{G}}^{(4)}_{ab}({\bf X})$ appears more frequently than the higher-priority token ${\bf e}_{5}\in\dot{\mathcal{G}}^{(4)}_{ab}({\bf X})$, illustrating our Theorem 4.

### v2

@@
 Using a simplified single-layer self-attention architecture, we derive analytical characterizations of topic changes.
-Specifically, we demonstrate that (1) the model maintains the priority order of tokens related to the input topic, (2) a topic change occurs only if lower-priority tokens outnumber all higher-priority tokens of the input topic, and (3) unlike human cognition, longer context lengths and overlapping topics reduce the likelihood of spontaneous redirection.
+Specifically, we demonstrate that (1) the model maintains the priority order of tokens related to the input topic, (2) a topic change can occur only if lower-priority tokens outnumber all higher-priority tokens of the input topic, and (3) unlike human cognition, longer context lengths and overlapping topics reduce the likelihood of spontaneous redirection.
 These insights highlight differences between human cognition and self-attention-based models in navigating topic changes and underscore the challenges in designing conversational AI capable of handling “spontaneous” conversations more naturally.
-To our knowledge, this is the first work to address these questions in such close relation to human conversation and thought.
+To the best of our knowledge, no prior work has explored these questions with a focus as closely aligned to human conversation and thought.
 1 Introduction
@@
 These seemingly chaotic shifts are intrinsic to the human way of processing and communicating information (Christoff and Fox,, 2018; Mills et al.,, 2020).
-For instance, consider a casual conversation where someone is discussing their favorite book, and then they suddenly mention an unrelated memory of a family gathering.
+For instance, someone discussing their favorite book might suddenly mention a family gathering.
 This abrupt change may be due to an emotional connection, such as recalling reading that book during a family vacation, where sensory details like the scent of the ocean or the warmth of the sun trigger a vivid memory.
-The hippocampus helps link these sensory cues to specific memories, resulting in an unexpected yet meaningful shift (Lisman et al.,, 2017).
-The prefrontal cortex then integrates these memories, emotions, and contextual information, allowing for a spontaneous redirection of conversation (Miller and Cohen,, 2001).
-Self-attention mechanisms in language models rely on contextual signals in the input to perform topic shifts, with next-token prediction guided by the relevance of preceding tokens (Vaswani et al.,, 2017).
-Unlike humans, who transition fluidly and intuitively between topics, often driven by emotion or creativity, transformers follow a structured, statistical approach.
-They remain on topic unless explicit cues signal a change, reflecting a key distinction between the somehow rigid processing of language models and the dynamic transitions of human cognition (Brown et al.,, 2020; Clark et al.,, 2019; Devlin et al.,, 2019; Bisk et al.,, 2020).
+The hippocampus helps link these sensory cues to memories, enabling a meaningful shift (Lisman et al.,, 2017).
+The prefrontal cortex then integrates these memories, emotions, and context to redirect conversations spontaneously (Miller and Cohen,, 2001).
+However, large language models (LLMs) rely on contextual signals in the input to perform topic shifts.
+Unlike humans, who transition intuitively between topics, often driven by emotion or creativity, LLMs follow a structured, statistical approach, remaining on topic unless explicit cues signal a change.
+Figure 2 illustrates this distinction using the first sentence of the book “One hundred years of solitude” (García Márquez,, 1967).
+Additionally, related studies on conversational agents have recognized the practical challenges of spontaneous topic changes and proposed various practical approaches (Lim et al.,, 2010; Hwang et al.,, 2024; Xie et al.,, 2021; Soni et al.,, 2022; Ni et al.,, 2022; Lin et al.,, 2023).
 It is also important to differentiate between topic changes and hallucinations.
@@
 Our focus is on understanding when and why self-attention models predict next tokens from a different topic of the input, examining the conditions under which an input sequence from one topic leads to a prediction in another.
-Recent advancements in the related field have substantially deepened our understanding of self-attention mechanisms.
+Recent advancements in the related field have substantially deepened our understanding of self-attention mechanisms in LLMs.
 The Transformer model, introduced by Vaswani et al., (2017), revolutionized natural language processing by using self-attention to determine the importance of each token in relation to others.
-This innovation enables Transformers to focus on relevant context when predicting the next token, significantly enhancing capabilities natural language understanding and generation (Brown et al.,, 2020; Devlin et al.,, 2019).
-Efforts to improve next token prediction have prioritized optimizing the efficiency of the self-attention mechanism, achieving empirical success across tasks such as machine translation, question answering, and text summarization.
-Additionally, connections between self-attention and support vector machines (SVM) have offered valuable insights into optimization strategies, and mechanisms within the next token prediction (Ataee Tarzanagh et al.,, 2023; Tarzanagh et al.,, 2023; Li et al., 2024b, ).
+This innovation enables Transformers to focus on relevant context when predicting the next token, significantly enhancing capabilities natural language understanding and generation(Brown et al.,, 2020; Clark et al.,, 2019; Devlin et al.,, 2019; Bisk et al.,, 2020).
+Additionally, connections between self-attention and support vector machines (SVMs) have offered valuable insights into optimization strategies, and mechanisms within the next token prediction (Tarzanagh et al., 2023b, ; Tarzanagh et al., 2023a, ; Li et al., 2024b, ).
 These advancements underscore the transformative role of attention-based architectures in token prediction tasks.
@@
 However, our approach aligns more closely with human cognitive processes, where the input is typically centered on a single topic, ensuring that discussions or conversations remain within a consistent context or theme.
-Despite these advancements, our understanding of how LLMs manage topic changes in comparison to human cognition remains limited.
+Figure 1: Illustration of the difference between human cognition and AI contextual completion.
+The original fragment of “One hundred years of solitude" (García Márquez,, 1967) (top) has a clear “spontaneous" thought, but the GPT-2’s completion [^2] Just to illustrate, we use the prompt Please continue this short sentence, forgetting about “One hundred Years of Solitude”, since on a real conversation the LLM would be blind to the final output.
+(bottom), demonstrates continuity.
+Despite these advancements, our understanding of how LLMs manage topic changes compared to human cognition remains limited.
 Human cognition seamlessly navigates such transitions through spontaneous thought and contextual awareness (Christoff and Fox,, 2018).
@@
 This fundamental understanding will pave the way for uncovering the role of attention patterns in driving “spontaneous” topic changes.
-To the best of our knowledge, this is the first work to explore these questions in direct relation with human conversation and thought.
+To the best of our knowledge, there are no other studies that have investigated these dynamics so closely in relation to human conversation and thought processes.
 Summary of Findings
@@
     Changing topics triggered by token frequency.
-In Theorem 4, we show that only if a lower-priority token appears more frequently than all higher-priority tokens of Topic A, the oracle’s responses may begin to reflect Topic B, signaling a shift.
+In Theorem 4, we show that only if a lower-priority token appears more frequently than all higher-priority tokens of Topic A, the oracle’s responses may reflect a change of topic.
   3. 3.
@@
 Unlike human dialogue, where extended discussions often encourage spontaneous topic shifts and overlapping topics promote cognitive connections that facilitate smooth transitions, our results highlight a fundamental limitation of attention-based models in replicating these human-like dynamics.
-Figure 1: $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ (left) are token-priority graphs (TPGs) of the last token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively.
-$\mathcal{G}^{(4)}_{ab}$ (right) is the combination of TPG for $\mathbb{T}_{a}$ and TPG for $\mathbb{T}_{b}$.
-Figure 2: Illustration of topic continuity, ambiguous sequence, and change of topic (see Definition 3,  4, and 5 for formal definitions).
-For each input sequence derived from $\mathbb{T}_{a}$, we employ a self-attention model trained on $\mathbb{T}_{a}$ and another model trained on the mixed-topics dataset $\mathbb{T}_{ab}$ to predict the next tokens $\hat{y}_{{\bf w}_{a}}$ and $\hat{y}_{{\bf w}_{ab}}$, respectively (orange).
-$\widehat{\mathcal{G}}^{(4)}_{ab}$ is the highest probability SCC (see Definition 1) in mixed-topics setting and $\widehat{\mathcal{G}}^{(4)}_{a}$ is the highest probability SCC in $\mathbb{T}_{a}$, The input sequence is classified as topic continuity if $\hat{y}_{{\bf w}_{ab}}\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$ (top).
-The input sequence is ambiguous if $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$, but $\widehat{\mathcal{G}}^{(4)}_{a}(X)\subset\widehat{\mathcal{G}}^{(4)}_{ab}(X)$ (middle).
-${\bf W}_{ab}$ is considered to change of topic, if $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$ and $\widehat{\mathcal{G}}^{(4)}_{a}(X)\not\subset\widehat{\mathcal{G}}^{(4)}_{ab}(X)$ (bottom).
-The numeric details are provided in Figure 6 in the appendix.
-Here, we outline the structural overview of our approach.
-Following the framework of token-priority graphs (TPGs) introduced by Li et al., 2024b , we start with defining a topic as a set of TPGs (see Sec.
-3).
-TPGs represent the priority order of tokens conditioned on the last token of the input sequence.
+Figure 2: Overview of our approach.
+We define a topic by a set of TPGs $\{\mathcal{G}^{(k)}\}^{K}_{k=1}$ (Def.
+2) and generate a dataset for each topic such that the dataset can reconstruct the corresponding TPGs exactly (Assumption 4).
+The combination of dataset A and dataset B becomes the dataset for the mixed-topic model.
+We train self-attention models independently on each dataset.
+Then, we generate a new input sequence from topic A and predict the next token with two models, self-attention model A and self-attention model (mixed-topic).
+The next-token prediction with the mixed-topic model is categorized into three outcomes: keeps topic A (topic continuity from Def.
+3), ambiguous sequence (from Def.
+4), changes topic A (change of topic from Def.
+5).
+Further details for each category are in Figure 4.
+Following the framework of token-priority graphs (TPGs) introduced by Li et al., 2024b , we start with defining a topic as a set of TPGs (see Sec 3).
 For simplicity, we consider two different topics, Topic A and Topic B, characterized by two different sets of TPGs.
-We demonstrate the preservation of input topic priorities in a model trained with mixed-topic datasets (see Sec.
-4).
+We demonstrate the preservation of input topic priorities in a model trained with mixed-topic datasets (see Sec 4).
 Additionally, we introduce three concepts to categorize the prediction outputs: topic continuity, ambiguous sequence, and change of topic.
-We subsequently establish the conditions under which a self-attention model induces a topic change, showing that the probability of topic changes does not increase with longer input sequences or the presence of overlapping topics (see Sec.
-5).
+Figure 2 outlines the structural overview of our approach.
+We subsequently establish the conditions under which a self-attention model induces a topic change, showing that the probability of topic changes does not increase with longer input sequences or the presence of overlapping topics (see Sec 5).
+All proofs are refer to Appendix A.
 To formalize this approach, we begin with the structured problem setup.
 2 Problem Setup
-Next topic prediction with self-attention model.
-In line with the approach presented by Ataee Tarzanagh et al., (2023) and Li et al., 2024b , we frame the next-token prediction task as a multi-class classification problem.
+Next Topic Prediction with Self-attention Model
+In line with the approach presented by Tarzanagh et al., 2023b and Li et al., 2024b , we frame the next-token prediction task as a multi-class classification problem.
 Given a vocabulary of size $K$ with an embedding matrix ${\bf E}=[{\bf e}_{1}\ {\bf e}_{2}\ \cdots\ {\bf e}_{K}]^{\top}\in\mathbb{R}^{K\times d}$, we aim to predict the next token ID $y\in[K]$ based on an input sequence ${\bf X}=[{\bf x}_{1}\ {\bf x}_{2}\ \cdots\ {\bf x}_{T}]^{\top}\in\mathbb{R}^{T\times d}$ with ${\bf x}_{i}\in{\bf E}$ for all $i\in[T]$.
@@
 In our notation ${\bf x}$ is the embedding vector corresponding to the token ID $x$, this is ${\bf x}={\bf e}_{x}$.
-For prediction, we utilize a single-layer self-attention model with a combined key-query weight matrix ${\bf W}\in\mathbb{R}^{d\times d}$ and identity value matrix as in Ataee Tarzanagh et al., (2023).
+For prediction, we utilize a single-layer self-attention model with a combined key-query weight matrix ${\bf W}\in\mathbb{R}^{d\times d}$ and identity value matrix as in Tarzanagh et al., 2023b .
 The self-attention embedding output
 $f_{\bf W}({\bf X})={\bf X}^{\top}\mathbb{S}({\bf X}{\bf W}\bar{{\bf x}}),$ (output)
-where $\mathbb{S}(\cdot)$ is the softmax operation and $\bar{{\bf x}}:={\bf x}_{T}$, serves as a weighted representation of the tokens, allowing for context-sensitive prediction of ${\bf y}$ based on the final input token.
+where $\mathbb{S}(\cdot)$ is the softmax operation and $\bar{{\bf x}}:={\bf x}_{T}$, serves as a weighted representation of the tokens, allowing for context-sensitive prediction of $y$ based on the final input token.
 Let $\ell:\mathbb{R}\to\mathbb{R}$ be a loss function.
-For the training set DSET, we aim to minimize empirical risk (ERM) with:
+For the training dataset DSET, we consider the empirical risk minimization (ERM) with:
 $L({\bf W})=\frac{1}{n}\sum_{i=1}^{n}\ell({\bf c}_{y_{i}}^{\top}{\bf X}_{i}^{\top}\mathbb{S}({\bf X}_{i}{\bf W}\bar{{\bf x}}_{i})).$ (ERM)
@@
 ${\bf W}^{(\tau+1)}={\bf W}^{(\tau)}-\eta\nabla L({\bf W}^{(\tau)}).$ (Algo-GD)
-We keep the same first two assumptions from Li et al., 2024b :
+We keep the first two assumptions from Li et al., 2024b :
 Assumption 1.
@@
 For any $({\bf X},y)\in\text{DSET}$, the token ${\bf e}_{y}$ is contained in the input sequence ${\bf X}$.
-Assumption 1 represents a variation of the weight-tying approach commonly used in language models (Press and Wolf,, 2016; Vaswani et al.,, 2017).
+Assumption 1 represents a variation of the weight-tying approach commonly used in language models (Press and Wolf,, 2017; Vaswani et al.,, 2017).
 Once training is complete, for a new input sequence ${\bf X}$, and a model characterized by $\bf W$, we predict the next token ID $\hat{y}_{\bf w}$ based on greedy decoding the probabilities from the softmax of the classification output
@@
 $\text{DSET}^{(k)}$ is a subset of sequences from DSET with the same last token is ${\bf e}_{k}=\bar{\bf x}$.
-They defined TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ such that every $\mathcal{G}^{(k)}$ is a directed graph where for every sequence $({\bf X},{\bf y})\in\text{DSET}^{(k)}$ a directed edge is added from ${\bf y}$ to every token ${\bf x}\in{\bf X}$.
+They defined TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ such that every $\mathcal{G}^{(k)}$ is a directed graph where for every sequence $({\bf X},y)\in\text{DSET}^{(k)}$ a directed edge is added from ${\bf e}_{y}$ to every token ${\bf x}\in{\bf X}$.
 TPGs are further divided into strongly-connected components (SCCs), which capture subsets of tokens where priority relations are cyclic, indicating comparable token importance.
@@
 […]
-Here a condensed version of the theorem:
-Theorem 1 (from Li et al., 2024b ).
+Here is a condensed version of the theorem:
+Theorem 1 (Li et al., 2024b ).
 Consider dataset DSET and suppose Assumptions 1 and 2 hold.
@@
 This convergence implies that the model predicts the next token based on priorities obtained from the SCCs within the TPG relevant to the last token of the input sequence.
+Unlike the work in Li et al., 2024b , which considers multiple loss functions in subsequent results, we focus exclusively on a log-loss function in this work, leaving the exploration of other loss functions for future research.
 We add here another reasonable assumption that prevents the probabilities in Equation 1 to be equal for improbable numerical reasons, and we present our first lemma.
 Assumption 3.
-For any $({\bf X},y)\in\text{DSET}$, $\exists i,j\in[T]\text{ and }u,v\in\mathbb{Z}$ such that $u\left[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})\right]_{i}=v\left[\mathbb{S}({\bf X}{\tilde{\bf W}}\bar{{\bf x}})\right]_{j}\iff u=v\text{ and }\left[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})\right]_{i}=\left[\mathbb{S}({\bf X}{\tilde{\bf W}}\bar{{\bf x}})\right]_{j}$.
+For any $({\bf X},y)\in\text{DSET}$, $\exists i,j\in[T]\text{ and }u,v\in\mathbb{Z}$ such that $u\left[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})\right]_{i}=v\left[\mathbb{S}({\bf X}{\tilde{\bf W}}\bar{{\bf x}})\right]_{j}\text{ if and only if }u=v\text{ and }\left[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})\right]_{i}=\left[\mathbb{S}({\bf X}{\tilde{\bf W}}\bar{{\bf x}})\right]_{j}$.
 Lemma 2.
 Suppose conditions from Theorem 1 and Assumption 3 hold.
-Consider an input sequence ${\bf X}$ from DSET(k) and corresponding TPG $\mathcal{G}^{(k)}$, $\forall i,j\in[K]$ we have $\left[\mathbb{S}\left({\bf C}f_{\bf W}({\bf X})\right)\right]_{i}=\left[\mathbb{S}\left({\bf C}f_{\tilde{\bf W}}({\bf X})\right)\right]_{j}\iff(x_{i}\asymp x_{j})\in\mathcal{G}^{(k)}$.
+Consider an input sequence ${\bf X}$ from DSET(k) and corresponding TPG $\mathcal{G}^{(k)}$, $\forall i,j\in[K]$ we have $\left[\mathbb{S}\left({\bf C}f_{\bf W}({\bf X})\right)\right]_{i}=\left[\mathbb{S}\left({\bf C}f_{\tilde{\bf W}}({\bf X})\right)\right]_{j}\text{ iff }(x_{i}\asymp x_{j})\in\mathcal{G}^{(k)}$.
 This means that the tokens that maximize the probability in Equation 1 are all within the same SCC leading to the following definition:
@@
 A DSET generated from any topic $\mathbb{T}$ defined by $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ exactly reconstructs back the TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$.
-This assumption enables the application of the results from Li et al., 2024b , with the concepts of topics and TPGs being used interchangeably.
+Detailed explanation is provided in Appendix B. This assumption enables the application of the results from Li et al., 2024b , with the concepts of topics and TPGs being used interchangeably.
 Definition 3 (topic continuity).
@@
 […]
+Figure 3: Illustration of topic combination.
+$\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ are token-priority graphs (TPGs) of the last token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively.
+$\mathcal{G}^{(4)}_{ab}$ combines the TPG for $\mathbb{T}_{a}$ and the TPG for $\mathbb{T}_{b}$.
 This implies that for an input sequence $\mathbf{X}$, a model trained in a mixed-topic setting will maintain the priority of the topic to which $\mathbf{X}$ belongs.
@@
 The strict equality ${\bf e}_{1}={\bf e}_{2}$ from $\mathcal{G}^{(4)}_{a}$ is maintained in $\mathcal{G}^{(4)}_{ab}$, but the strict inequality ${\bf e}_{2}>{\bf e}_{4}$ is replaced with ${\bf e}_{2}={\bf e}_{4}$ in mixed-topics, forming the new SCC, $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\}$, in $\mathcal{G}^{(4)}_{ab}$.
-Figure 3: Convergence of attention priority similarity for $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue) and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (orange).
+Figure 4: Illustration of topic continuity, ambiguous sequence, and change of topic (see Definition 3,  4, and 5 for formal definitions).
+For each input sequence derived from $\mathbb{T}_{a}$, we employ a self-attention model trained on $\mathbb{T}_{a}$ and another model trained on the mixed-topics dataset $\mathbb{T}_{ab}$ to predict the next tokens $\hat{y}_{{\bf w}_{a}}$ and $\hat{y}_{{\bf w}_{ab}}$, respectively (orange).
+$\widehat{\mathcal{G}}^{(4)}_{ab}$ is the highest probability SCC (see Definition 1) in mixed-topics setting and $\widehat{\mathcal{G}}^{(4)}_{a}$ is the highest probability SCC in $\mathbb{T}_{a}$.
+Top: The input sequence is classified as topic continuity if $\hat{y}_{{\bf w}_{ab}}\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$.
+Middle: The input sequence is ambiguous if $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$, but $\widehat{\mathcal{G}}^{(4)}_{a}(X)\subset\widehat{\mathcal{G}}^{(4)}_{ab}(X)$.
+Bottom: ${\bf W}_{ab}$ is considered to change of topic, if $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$ and $\widehat{\mathcal{G}}^{(4)}_{a}(X)\not\subset\widehat{\mathcal{G}}^{(4)}_{ab}(X)$.
+The numeric details are provided in Appendix C.3.
 To further illustrate Theorem 3 we define the attention priority similarity of weights ${\bf W}^{\prime}$ relative to ${\bf W}$ for a sequence ${\bf X}$ as: $R_{{\bf W},{\bf W}^{\prime}}({\bf X})=$
@@
 We generate test sequences ${\bf Z}$ within $\mathbb{T}_{a}$, and we calculate the attention priority similarity of $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to both $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$.
-We repeat this process for multiple TPGs and input sequences (details in Appendix).
+We repeat this process for multiple TPGs and input sequences (simulation details in Appendix C).
 Figure 5 clearly demonstrates that the similarity converges to 1 after $\tau=8000$ iterations when evaluated relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue line), but fails to converge relative to $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ (orange line).
 These observations align with the results of Theorem 3.
+Figure 5: Convergence of attention priority similarity for $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue) and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (orange).
 5 Explaining Topic Shifts
@@
 In our analogy on the oracle, gaining knowledge of both Topic A and Topic B might cause a conversation to be naturally followed within Topic A or also outside Topic A.
-Thus, we introduce the following definition to characterize this phenomenon:
+We introduce the following definition to characterize this phenomenon:
 Definition 4 (ambiguous sequence).
@@
 Definition 5 (change of topic).
-Given DSETa and DSETb generated from two different topics $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, and a sequence $\bf X$ that belongs to $\mathbb{T}_{a}$.
+Given DSETa and DSETb generated from two topics $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, and a sequence $\bf X$ that belongs to $\mathbb{T}_{a}$.
 The weight matrix $\tilde{\bf W}_{ab}$ changes topic $\mathbb{T}_{a}$ for sequence ${\bf X}$ if $\tilde{\bf W}_{ab}$ does not keep topic $\mathbb{T}_{a}$ for ${\bf X}$ and $\bf X$ is not ambiguous in $\mathbb{T}_{ab}$ with respect to $\mathbb{T}_{a}$.
-Following the above definition, ${\bf W}_{ab}$ for the input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{4},{\bf e}_{4},{\bf e}_{4}]^{\top}$, from Figure 2 (bottom), can be considered as change of topic.
+In Figure 4, ${\bf W}_{ab}$ for the last input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{4},{\bf e}_{4},{\bf e}_{4}]^{\top}$ can be considered as change of topic.
 Building on the formal definitions of topic continuity, ambiguous sequences, and topic changes, we now present a necessary condition for a sequence to induce a topic change.
@@
 Consider a sequence ${\bf X}$ that belongs to $\mathbb{T}$.
-We define $\dot{\mathcal{G}}^{(\bar{x})}({\bf X})\in\mathcal{G}^{(\bar{x})}$ as the highest priority SCC for ${\bf X}$ in $\mathcal{G}^{(\bar{x})}$ such that $\forall{\bf x}_{i}\in\dot{\mathcal{G}}^{(\bar{x})}({\bf X})$ and ${\bf x}_{j}\in\mathcal{G}^{(\bar{x})}$ we have $(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}^{(\bar{x})}$.
+We define $\dot{\mathcal{G}}^{(\bar{x})}({\bf X})\subseteq\mathcal{G}^{(\bar{x})}$ as the highest priority SCC for ${\bf X}$ in $\mathcal{G}^{(\bar{x})}$ such that $\forall{x}_{i}\in\dot{\mathcal{G}}^{(\bar{x})}({\bf X})$ and ${x}_{j}\in\mathcal{G}^{(\bar{x})}$ we have $(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}^{(\bar{x})}$.
 Theorem 4.
@@
 First, as the input sequence length increases sufficiently, the likelihood of topic changes vanishes.
-Second, increasing $l$ leads to possibly higher overlap between topics, meaning that the probability of satisfying the necessary conditions for a topic change decreases, effectively creating a bound on the frequency of topic changes.
+Second, notice that increasing $l$ increases the probability of overlap between topics, and it also decreases the probability of satisfying the necessary conditions for a topic change, effectively creating a bound on the frequency of topic changes.
 In practice, consider the oracle analogy: if the oracle is following a sufficiently long conversation on a specific topic, it becomes exceedingly unlikely to shift topics.
@@
 This contrasts with human cognition, where longer conversations and greater inter-connectivity of knowledge increase the likelihood of spontaneous shifts.
-(a) Effect of increased sequences.
-(b) Effect of increased overlap.
-Figure 4: The proportion of keep topic, ambiguity, and change topic as (a) the test sequence length increases and (b) the number of edges in generated TPGs increase.
+(a) Length of sequences.
+(b) Overlap between topics.
+Figure 6: The proportion of keep topic (topic continuity), ambiguity (ambiguous sequence), and change topic (change of topic) as (a) the input sequence length increases and (b) the overlap between topics increases.
 To illustrate Theorem 5, we conduct simulations following the same implementation settings as described in Section 4.
@@
 We increase $T$ from 4 to 512.
-Figure 4 illustrates how the proportion of change topic decreases as $T$ increases.
+Figure 6(a) illustrates how the proportion of change topic decreases as $T$ increases.
 Second, we investigate the effect of overlapping topics by generating TPGs with an increased number of edges $L$.
-Intuitively, a higher $L$ results in greater overlap between TPGs from different topics.
+Intuitively, a higher $L$ results in an increase $l$ and a greater overlap between TPGs of different topics.
 We vary $L$ from 4 to 18.
-Figure 4 demonstrates that as $L$ increases, ambiguity increases, while proportion of change topic decreases.
+Figure 6(b) demonstrates that as $L$ increases, ambiguity increases, while proportion of change topic decreases.
 These two findings contrast with expectations derived from human cognition but align with the result of Theorem 5.
@@
 The remaining 0.02% mismatched cases were solely due to minor approximation discrepancies in the attention softmax.
-These results validate Theorem 4 (see specific experimental details in the Appendix).
+These results validate Theorem 4 (simulation details in Appendix C).
 Related Work
@@
 (2) Optimization in attention-based models.
-Additionally, recent researches interpret Transformer models as kernel machines, akin to support vector machines (SVMs), with self-attention layers performing maximum margin separation in the token space (Tarzanagh et al.,, 2023; Ataee Tarzanagh et al.,, 2023; Li et al., 2024b, ; Julistiono et al.,, 2024).
+Additionally, recent researches interpret Transformer models as kernel machines, akin to support vector machines (SVMs), with self-attention layers performing maximum margin separation in the token space (Tarzanagh et al., 2023a, ; Tarzanagh et al., 2023b, ; Li et al., 2024b, ; Julistiono et al.,, 2024).
 (3) Chain-of-Thought and In-Context Learning.
 Moreover, transformers exhibit remarkable abilities in generalization through in-context learning (ICL), where models effectively learn from contextual cues during inference (Brown et al.,, 2020; Xie et al.,, 2022; Olsson et al.,, 2022).
-Chain-of-Thought (CoT) prompting (Wei et al.,, 2022; Zhou et al.,, 2022; Shao et al.,, 2023; Li et al., 2024a, ) enhances this by breaking down reasoning processes into intermediate steps, highlighting the emergent reasoning abilities of transformers.
+Chain-of-Thought (CoT) prompting (Wei et al.,, 2022; Zhou et al.,, 2023; Shao et al.,, 2023; Li et al., 2024a, ) enhances this by breaking down reasoning processes into intermediate steps, highlighting the emergent reasoning abilities of transformers.
 (4) Improvement efficiency of transformers.
@@
 6 Discussion
-Our findings on a simple single-layer attention model provide valuable clues about the fundamental differences in conversational dynamics between humans and machines.
-Unlike humans, who naturally shift topics through curiosity or contextual association, the model adheres to predictive patterns, mirroring LLMs’ tendency to stay focused on a single thread.
-This suggests that spontaneous topic shifts, a hallmark of human conversation, may require mechanisms beyond predictive pattern recognition.
-These insights raise philosophical questions about the nature of human-like interaction and highlight the need for adaptability and creativity in advancing conversational AI.
-Future research will focus on how real-world LLMs perform changing topics, where the interplay of diverse and evolving contexts presents unique challenges.
-It would also be interesting to investigate the connection between LLM and human cognition in spontaneously changing topics.
-Additionally, the recurring patterns of topic retention identified in this work open new pathways for applications.
-Harnessing these behavioral signatures could enhance changing topics detection in conversational AI, enrich content generation by anticipating and adapting to topic transitions, and refine model interpretability.
-Pursuing these directions could significantly advance both the theoretical landscape and practical utility of LLMs.
+Our exploration with a simple single-layer attention model reveals fundamental clues on the distinctions between human and AI conversational dynamics.
+While humans naturally shift topics through curiosity or contextual triggers, the model remains anchored to the context.
+These observations suggest that human-like spontaneity in topic changes demands mechanisms beyond pattern recognition, raising deeper questions about the nature of interaction and highlighting the importance of creativity in advancing conversational AI.
+However, our analysis relies on the simplified single-layer self-attention setting from Li et al., 2024b .
+This does not directly extend to complex modern LLMs with deeper architectures and varied training goals.
+Future work will explore how these findings translate to more intricate models, examining their ability to change topics.
+Moreover, the recurring theme of topic continuity characterized in this work opens new avenues for applications, such as automatic change topic detection in conversational AI based on discrepancies between model predictions and speaker outputs.
 Appendix A Technical Proofs
 A.1 Proofs of Lemmas
-Proof of Lemma 2
+A.1.1 Proof of Lemma 2
 Let ${\bf a}=\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})$.
@@
 If $x_{i}\neq x_{j}$ then $x_{i}$ and $x_{j}$ are in the same SCC.∎
+A.1.2 Proof of Lemma 6
 Lemma 6.
@@
   • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}<1\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$.
+A.1.3 Proof of Lemma 7
 Lemma 7.
@@
 Therefore $P(l+1)=\mathbb{P}(B_{l+1}>A_{l+1})\leq\mathbb{P}(B_{l}>A_{l})=P(l)$.
-Appendix B Detailed simulation studies
-B.1 Simulation process
+Appendix B Detailed Explanation of Assumption 4
+As illustrated in Figure 7, the dataset for $\mathbb{T}_{a}$ and the dataset for $\mathbb{T}_{b}$ demonstrate interchangeability with $\mathcal{G}_{a}^{(4)}$ and $\mathcal{G}_{b}^{(4)}$, respectively.
+Figure 7: Illustration of Assumption 4.
+Here are two datasets related to the TPGs, $\mathcal{G}_{a}^{(4)}$ and $\mathcal{G}_{b}^{(4)}$, from Figure 3.
+From the directed arrows in $\mathcal{G}_{a}^{(4)}$, we can generate a dataset with the last token $\bf{e}_{4}$ for $\mathbb{T}_{a}$, which can reconstruct back the $\mathcal{G}_{a}^{(4)}$.
+A similar process applies for the $\mathcal{G}_{b}^{(4)}$.
+Appendix C Detailed Simulation Studies
+C.1 Simulation Process
 Theoretical TPGs generation.
@@
 To reduce the potential numerical issues in the outputs, $\mathbb{S}({\bf X}{\bf W}\bar{{\bf x}})$ is rounded to three decimals, ensuring that tokens within the same SCC yield consistent softmax outputs.
-B.2 Experiments results
+Our code for all simulations is available in ( ).
+C.2 Experiments Results
 In this section, we provide empirical evidence to support our findings outlined in Theorem 3, Theorem 4, and Theorem 5, offering detailed insights into the behavior of self-attention mechanisms for changing topics.
-B.2.1 Simulation in Section 4
+C.2.1 Simulation in Section 4
 For Figure 5 we used $K=10,L=4,d=16,\text{and }T_{test}=T_{train}=4$, we also showed point-wise confidence intervals of rank similarities on all 5000 instances for each value of $\tau$.
-B.2.2 Simulation in Section 5
-For Figure 4 (left), we predict next tokens for 5000 test sequences from $\mathbb{T}_{a}$ with $T_{test}=\{4,8,16,24,32,64,128,256,512\}$, while fixing $L=4$, $d=16$, $T_{train}=4$, and $K=10$.
-The proportion of each scenario with varying $T$ is illustrated in Table 1.
-For Figure 4 (right), we predict next tokens for 5000 test sequences (the sequence length is $T_{test}=20$) using models trained with $L=\{4,6,8,10,12,14,16,18\}$, $d=16$, $K=10$, and $T_{train}=4$.
-The proportion of each scenario with varying $L$ is illustrated in Table 2.
+C.2.2 Simulation in Section 5
 Table 1: The proportion of keep topic, ambiguous, and change of topic with varying $T_{test}=\{4,8,16,24,32,64,128,256,512\}$.
@@
 18 98.02 $\pm$ 1.58 1.26 $\pm$ 1.14 0.72 $\pm$ 0.86
-B.2.3 Additional experiments
+In Figure 6(a), we predict next tokens for 5000 test sequences from $\mathbb{T}_{a}$ with $T_{test}=\{4,8,16,24,32,64,128,256,512\}$, while fixing $L=4$, $d=16$, $T_{train}=4$, and $K=10$.
+The proportion of each scenario with varying $T$ is illustrated in Table 1.
+For Figure 6(b), we predict next tokens for 5000 test sequences (the sequence length is $T_{test}=20$) using models trained with $L=\{4,6,8,10,12,14,16,18\}$, $d=16$, $K=10$, and $T_{train}=4$.
+The proportion of each scenario with varying $L$ is illustrated in Table 2.
+C.2.3 Additional experiments
 (a) $K=6,L=4$
@@
 The shaded area around each line represents the $95\%$ confidence interval, computed over 50 epochs.
-B.3 Numerical analysis of each scenario in Figure 2
+C.3 Numerical Analysis of each Scenario in Figure 4
 (a) Topic continuity.
@@
 In Fig.
-6(a), input sequence $\bf X$ consists of four unique tokens: ${\bf e}_{5}$, ${\bf e}_{1}$, ${\bf e}_{3}$, and ${\bf e}_{4}$.
+9(a), input sequence $\bf X$ consists of four unique tokens: ${\bf e}_{5}$, ${\bf e}_{1}$, ${\bf e}_{3}$, and ${\bf e}_{4}$.
 Based on $\mathcal{G}^{(4)}_{a}$ in Figure 3, the priority order of these tokens is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}>{\bf e}_{4}$, with corresponding $a_{i}$ values: $0.45>0.25>0.20>0.1$.
@@
 Input sequence $\bf X$ in Fig.
-6(b) has two unique tokens: ${\bf e}_{1}$ and ${\bf e}_{4}$.
+9(b) has two unique tokens: ${\bf e}_{1}$ and ${\bf e}_{4}$.
 The priority order is ${\bf e}_{1}>{\bf e}_{4}$, following $\mathcal{G}^{(4)}_{a}$ in Figure 3.
@@
 For the input sequence $\bf X$ in Fig.
-6(c), the only two unique tokens, ${\bf e}_{5}$ and ${\bf e}_{4}$, are with the same priority order in both $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{ab}$ from Figure 1: ${\bf e}_{5}>{\bf e}_{1}$.
+9(c), the only two unique tokens, ${\bf e}_{5}$ and ${\bf e}_{4}$, are with the same priority order in both $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{ab}$ from Figure 3: ${\bf e}_{5}>{\bf e}_{1}$.
 With ${\bf W}_{a}(\tau)$ trained in $\mathbb{T}_{a}$, the token ${\bf e}_{5}$ has $a_{1}=0.70$ and the token ${\bf e}_{4}$ has $a_{2}=a_{3}=a_{4}=0.10$.

### v3

@@
-Dynamics of “Spontaneous” Topic Changes in Next Token Prediction with Self-Attention
+Dynamics of Spontaneous Topic Changes in Next Token Prediction with Self-Attention
 Abstract
-Human cognition can spontaneously shift conversation topics, often triggered by emotional or contextual signals.
-In contrast, self-attention-based language models depend on structured statistical cues from input tokens for next-token prediction, lacking this spontaneity.
-Motivated by this distinction, we investigate the factors that influence the next-token prediction to change the topic of the input sequence.
-We define concepts of topic continuity, ambiguous sequences, and change of topic, based on defining a topic as a set of token priority graphs (TPGs).
-Using a simplified single-layer self-attention architecture, we derive analytical characterizations of topic changes.
-Specifically, we demonstrate that (1) the model maintains the priority order of tokens related to the input topic, (2) a topic change can occur only if lower-priority tokens outnumber all higher-priority tokens of the input topic, and (3) unlike human cognition, longer context lengths and overlapping topics reduce the likelihood of spontaneous redirection.
-These insights highlight differences between human cognition and self-attention-based models in navigating topic changes and underscore the challenges in designing conversational AI capable of handling “spontaneous” conversations more naturally.
-To the best of our knowledge, no prior work has explored these questions with a focus as closely aligned to human conversation and thought.
+Human cognition is punctuated by abrupt, spontaneous shifts between topics—driven by emotional, contextual, or associative cues—a phenomenon known as spontaneous thought in neuroscience.
+In contrast, self-attention based models depend on structured patterns over their inputs to predict each next token, lacking spontaneity.
+Motivated by this distinction, we characterize spontaneous topic changes in self-attention architectures, revealing both their similarities and their divergences from spontaneous human thought.
+First, we establish theoretical results under a simplified, single-layer self-attention model with suitable conditions by defining the topic as a set of Token Priority Graphs (TPGs).
+Specifically, we demonstrate that (1) the model maintains the priority order of tokens related to the input topic, (2) a spontaneous topic change can occur only if lower-priority tokens outnumber all higher-priority tokens of the input topic, and (3) unlike human cognition, the longer context length or the more ambiguous input topic reduces the likelihood of spontaneous change.
+Second, we empirically validate that these dynamics persist in modern, state-of-the-art LLMs, underscoring a fundamental disparity between human cognition and AI behaviour in the context of spontaneous topic changes.
+To the best of our knowledge, no prior work has explored these questions with a focus as closely aligned to human thought.
 1 Introduction
-In human cognition, topic changes often occur sporadically, with little apparent structure or forethought: an spontaneous shift in focus during a conversation, a sudden leap between ideas when brainstorming, or an unexpected redirection in storytelling.
-These seemingly chaotic shifts are intrinsic to the human way of processing and communicating information (Christoff and Fox,, 2018; Mills et al.,, 2020).
-For instance, someone discussing their favorite book might suddenly mention a family gathering.
-This abrupt change may be due to an emotional connection, such as recalling reading that book during a family vacation, where sensory details like the scent of the ocean or the warmth of the sun trigger a vivid memory.
-The hippocampus helps link these sensory cues to memories, enabling a meaningful shift (Lisman et al.,, 2017).
-The prefrontal cortex then integrates these memories, emotions, and context to redirect conversations spontaneously (Miller and Cohen,, 2001).
-However, large language models (LLMs) rely on contextual signals in the input to perform topic shifts.
-Unlike humans, who transition intuitively between topics, often driven by emotion or creativity, LLMs follow a structured, statistical approach, remaining on topic unless explicit cues signal a change.
+Human cognition is punctuated by abrupt, apparently unstructured topic changes, the hallmark of human spontaneous thought, a phenomenon that has become a central topic in cognitive neuroscience (Bellana et al.,, 2022; Christoff and Fox,, 2018; Christoff et al.,, 2011; Kucyi et al.,, 2023; Mildner and Tamir,, 2019, 2024; Mills et al.,, 2020).
+For example, a spontaneous shift in focus during a conversation, a sudden leap between ideas when brainstorming, or an unexpected redirection in storytelling.
+These abrupt changes may be due to an emotional connection, such as recalling reading that book during a family vacation, where sensory details like the scent of the ocean or the warmth of the sun trigger a vivid memory.
+However, LLMs rely on contextual signals in the input to perform topic shifts, instead of spontaneous topic changes.
+They follow a structured, statistical approach, remaining on topic unless explicit cues signal a change.
 Figure 2 illustrates this distinction using the first sentence of the book “One hundred years of solitude” (García Márquez,, 1967).
-Additionally, related studies on conversational agents have recognized the practical challenges of spontaneous topic changes and proposed various practical approaches (Lim et al.,, 2010; Hwang et al.,, 2024; Xie et al.,, 2021; Soni et al.,, 2022; Ni et al.,, 2022; Lin et al.,, 2023).
-It is also important to differentiate between topic changes and hallucinations.
-Hallucinations involve generating incorrect or fabricated information without clear contextual basis, often resulting in false or illogical content (Ji et al.,, 2023; Maynez et al.,, 2020).
-In contrast to topic changes, hallucinations are not deliberate or contextually driven shifts in discourse.
-More interestingly, some hallucinations may still appear to maintain topic consistency by staying loosely connected to the previous topic, yet they fail to introduce a meaningful or valid change.
-Our focus is on understanding when and why self-attention models predict next tokens from a different topic of the input, examining the conditions under which an input sequence from one topic leads to a prediction in another.
-Recent advancements in the related field have substantially deepened our understanding of self-attention mechanisms in LLMs.
-The Transformer model, introduced by Vaswani et al., (2017), revolutionized natural language processing by using self-attention to determine the importance of each token in relation to others.
-This innovation enables Transformers to focus on relevant context when predicting the next token, significantly enhancing capabilities natural language understanding and generation(Brown et al.,, 2020; Clark et al.,, 2019; Devlin et al.,, 2019; Bisk et al.,, 2020).
-Additionally, connections between self-attention and support vector machines (SVMs) have offered valuable insights into optimization strategies, and mechanisms within the next token prediction (Tarzanagh et al., 2023b, ; Tarzanagh et al., 2023a, ; Li et al., 2024b, ).
-These advancements underscore the transformative role of attention-based architectures in token prediction tasks.
-A recent study by Li et al., (2023) highlights that in mixed topic inputs, transformers achieve higher pairwise attention between same-topic words compared to different-topic words.
-However, our approach aligns more closely with human cognitive processes, where the input is typically centered on a single topic, ensuring that discussions or conversations remain within a consistent context or theme.
-Figure 1: Illustration of the difference between human cognition and AI contextual completion.
-The original fragment of “One hundred years of solitude" (García Márquez,, 1967) (top) has a clear “spontaneous" thought, but the GPT-2’s completion [^2] Just to illustrate, we use the prompt Please continue this short sentence, forgetting about “One hundred Years of Solitude”, since on a real conversation the LLM would be blind to the final output.
-(bottom), demonstrates continuity.
-Despite these advancements, our understanding of how LLMs manage topic changes compared to human cognition remains limited.
-Human cognition seamlessly navigates such transitions through spontaneous thought and contextual awareness (Christoff and Fox,, 2018).
-Investigating this distinction could provide valuable insights into the strengths and limitations of language models in mimicking human-like conversation.
-Bridging this gap is essential for enhancing language models, particularly in their ability to adapt to and manage dynamic topic changes–a process closely aligned with the flow of spontaneous thought in human cognition.
-Specifically, we address the following research questions:
-  • What underlying mechanism drives topic changes during next token prediction in self-attention architectures?
-  • Which factors shape the dynamics of topic changes in next token prediction?
-To answer the above questions, it is essential first to explore how self-attention allocates attention to tokens for next token prediction with respect to the input topic.
-Understanding how attention weights are allocated offers key insights into the prioritization of tokens by self-attention mechanisms, particularly when models are trained on datasets spanning multiple topics.
-This fundamental understanding will pave the way for uncovering the role of attention patterns in driving “spontaneous” topic changes.
-To the best of our knowledge, there are no other studies that have investigated these dynamics so closely in relation to human conversation and thought processes.
+Our work takes initial steps toward formalizing the dynamics of spontaneous topic changes in LLMs and analyzing how they relate to or diverge from human spontaneous thought.
+To this end, we ground our theoretical analysis in a single-layer self-attention model and empirically extend it to modern LLMs, laying groundwork for drawing parallels between AI models and human cognition.
+Recent advancements in the related field have substantially deepened our understanding of self-attention architectures.
+(Li et al., 2024b, ; Tarzanagh et al., 2023a, ; Tarzanagh et al., 2023b, ) have linked the self-attention to support vector machines (SVMs), offering optimization strategies for next-token prediction.
+A study by Li et al., (2023) highlights that in mixed-topics inputs, transformers achieve higher pairwise attention between same-topic words compared to different-topic words.
+In parallel, prior studies have recognized the practical challenges of spontaneous topic changes in LLMs and proposed various approaches to address them (Hwang et al.,, 2024; Lim et al.,, 2010; Lin et al.,, 2023; Ni et al.,, 2022; Soni et al.,, 2022; Xie et al.,, 2021).
+Notably, spontaneous topic changes must be differentiated from hallucinations, generating incorrect or fabricated information without clear contextual basis (Ji et al.,, 2023; Maynez et al.,, 2020).
+Despite these advancements, our understanding of the dynamics of spontaneous topic changes in LLMs remains limited.
+Investigating the parallels and differences between spontaneous topic changes in self-attention models and human spontaneous thought could provide valuable insights for improving the efficiency of language models with human-like thought.
+Since modern LLMs rely on self-attention architectures, we begin by theoretically characterizing spontaneous topic shifts in a simplified setting.
+We then extend these findings through experiments on more complex, state-of-the-art models.
+To the best of our knowledge, no prior studies have investigated these dynamics so closely in relation to human thought.
+Figure 2 outlines our theoretical framework.
+To make the mathematical analysis tractable, we follow the same single-layer self-attention framework with log-loss objective function governed by Assumptions 14 from Li et al., 2024b .
+Inspired by token-priority graphs (TPGs) (Li et al., 2024b, ) and building on attribution graphs from Ameisen et al., (2025) for exposing an LLM’s internal computation, we define a topic as a set of TPGs.
+This graph-based formulation aligns naturally with recent advances in structured representations for LLMs (Sen et al.,, 2023; Wang et al.,, 2025).
+Furthermore, this mirrors neuroscience models of human spontaneous thought, in which concepts serve as nodes connected by associative edges (Mildner and Tamir,, 2019).
+Despite relying on these specific settings, our experiments extend our findings to modern LLMs, empirically confirming that relaxing these assumptions does not seem to undermine our core insights.
+Figure 1: Illustration of the difference between human cognition and LLMs.
+The original fragment of “One hundred years of solitude" (García Márquez,, 1967) (top) has a clear spontaneous thought, but the GPT-2’s completion (bottom), demonstrates continuity.
+[^2] Just to illustrate, we use the prompt Please continue this short sentence, forgetting about “One hundred Years of Solitude”, since on a real conversation the LLM would be blind to the final output.
+Figure 2: Overview of our theoretical framework.
+We define a topic by a set of TPGs $\{\mathcal{G}^{(k)}\}^{K}_{k=1}$ (Def.
+2) and generate a dataset for each topic.
+The combination of dataset A and dataset B becomes the dataset for the mixed-topic model.
+We train self-attention models independently on each dataset.
+Then, we generate a new input sequence from topic A and predict the next token with two models, self-attention model A and self-attention model (mixed-topic).
+The next-token prediction with the mixed-topic model is categorized into three outcomes: keeps topic A (topic continuity from Def.
+3), ambiguous sequence (from Def.
+4), changes topic A (change of topic from Def.
+5).
+Further details for each category are in Figure 3.
 Summary of Findings
-For simplicity, we elaborate our results using a two-topic scenario, but it is straightforward to extend them to multiple topics.
 Imagine an oracle that is an expert on Topic A, capable of following any conversation within that topic while staying true to its context.
 Now, suppose the oracle gains knowledge of Topic B and is following a conversation about Topic A.
-Will the oracle’s responses remain firmly within Topic A, or will the influence of the knowledge of Topic B cause the conversation to drift?
-This analogy encapsulates the problem we address: understanding when and why attention models might preserve a topic or change to another “spontaneously".
+Will the oracle’s responses remain within Topic A, or will the influence of the knowledge of Topic B cause the conversation to drift?
+This analogy encapsulates the problem we address: understanding when and why attention models might preserve a topic or change to another spontaneously.
 Specifically, we make the following contributions:
@@
   3. 3.
-    Impact of sequence length and topic overlap.
+    Impact of sequence length and topic ambiguity.
 Theorem 5 establishes that longer input sequences decrease the likelihood of changing topics.
-Furthermore, overlapping topics act as a stabilizing factor, not increasing the frequency of “spontaneous” topic changes.
-This finding represents a key contribution to the development of human-like AI conversational models.
-Unlike human dialogue, where extended discussions often encourage spontaneous topic shifts and overlapping topics promote cognitive connections that facilitate smooth transitions, our results highlight a fundamental limitation of attention-based models in replicating these human-like dynamics.
-Figure 2: Overview of our approach.
-We define a topic by a set of TPGs $\{\mathcal{G}^{(k)}\}^{K}_{k=1}$ (Def.
-2) and generate a dataset for each topic such that the dataset can reconstruct the corresponding TPGs exactly (Assumption 4).
-The combination of dataset A and dataset B becomes the dataset for the mixed-topic model.
-We train self-attention models independently on each dataset.
-Then, we generate a new input sequence from topic A and predict the next token with two models, self-attention model A and self-attention model (mixed-topic).
-The next-token prediction with the mixed-topic model is categorized into three outcomes: keeps topic A (topic continuity from Def.
-3), ambiguous sequence (from Def.
-4), changes topic A (change of topic from Def.
-5).
-Further details for each category are in Figure 4.
-Following the framework of token-priority graphs (TPGs) introduced by Li et al., 2024b , we start with defining a topic as a set of TPGs (see Sec 3).
-For simplicity, we consider two different topics, Topic A and Topic B, characterized by two different sets of TPGs.
-We demonstrate the preservation of input topic priorities in a model trained with mixed-topic datasets (see Sec 4).
-Additionally, we introduce three concepts to categorize the prediction outputs: topic continuity, ambiguous sequence, and change of topic.
-Figure 2 outlines the structural overview of our approach.
-We subsequently establish the conditions under which a self-attention model induces a topic change, showing that the probability of topic changes does not increase with longer input sequences or the presence of overlapping topics (see Sec 5).
-All proofs are refer to Appendix A.
-To formalize this approach, we begin with the structured problem setup.
+Furthermore, topic ambiguity acts as a stabilizing factor, not increasing the frequency of spontaneous topic changes.
+  4. 4.
+    Difference between LLMs and human cognition.
+In Section 6 we empirically extend Theorem 5 to modern deeper LLMs.
+Unlike human cognition, where extended discussions often encourage spontaneous thoughts and topic ambiguity promotes cognitive connections, our results highlight an opposite behavior in LLMs: neither longer prompts nor greater topic ambiguity appreciably raise the likelihood of an spontaneous topic change.
+Overview of the paper structure.
+We begin with the problem setup in Sec 2.
+Sec 3 introduces the definition of topic, and Sec 4 examines how self-attention models allocate the token priorities within the mixed-topics.
+In Sec 5, we establish the conditions under which a self-attention model induces spontaneous topic changes and show the dynamics of topic changes with longer input sequences or the presence of topic ambiguity.
+We then extend our analysis to frontier LLMs in Sec 6.
+Related work and discussion are provided in Secs 7 and 8, respectively.
+All proofs are provided in Appendix A.
 2 Problem Setup
-Next Topic Prediction with Self-attention Model
+2.1 Next Topic Prediction with Self-attention Model
 In line with the approach presented by Tarzanagh et al., 2023b and Li et al., 2024b , we frame the next-token prediction task as a multi-class classification problem.
@@
 Assumption 1.
-$\forall y,k\in[K],k\neq y,{\bf c}_{y}^{\top}{\bf e}_{y}=1$ and ${\bf c}_{y}^{\top}{\bf e}_{k}=0$
+$\forall y,k\in[K],k\neq y,{\bf c}_{y}^{\top}{\bf e}_{y}=1$ and ${\bf c}_{y}^{\top}{\bf e}_{k}=0$.
 Assumption 2.
@@
 $\hat{y}_{\bf w}\in\arg\max_{k\in[K]}\left[\mathbb{S}\left({\bf C}f_{\bf W}({\bf X})\right)\right]_{k}.$ (1)
-Token-priority Graph and Global Convergence of the Self-attention Model
+2.2 Token-priority Graph and Global Convergence of the Self-attention Model
 Li et al., 2024b defined a token-priority graph (TPG) as a directed graph with nodes representing tokens in the vocabulary.
@@
 They defined TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ such that every $\mathcal{G}^{(k)}$ is a directed graph where for every sequence $({\bf X},y)\in\text{DSET}^{(k)}$ a directed edge is added from ${\bf e}_{y}$ to every token ${\bf x}\in{\bf X}$.
-TPGs are further divided into strongly-connected components (SCCs), which capture subsets of tokens where priority relations are cyclic, indicating comparable token importance.
+TPGs are further divided into strongly-connected components (SCCs), which capture subsets of tokens with equal priority.
 For tokens within two different SCCs, strict priority orders emerge, helping the model to differentiate between tokens when learning next-token predictions.
@@
   • $i\in\mathcal{G}$ denotes that the node $i$ belongs to $\mathcal{G}$.
-  • $(i\Rightarrow j)\in\mathcal{G}$ denotes that the directed path $(i\rightarrow j)$ is present in $\mathcal{G}$ but $j\rightarrow i$ is not.
+  • $(i\Rightarrow j)\in\mathcal{G}$ denotes that the directed path $(i\rightarrow j)$ is presented in $\mathcal{G}$ but $j\rightarrow i$ is not.
   • $(i\asymp j)\in\mathcal{G}$ means that both nodes $i$ and $j$ are in the same strongly connected component (SCC) of $\mathcal{G}$ (there exists both a path $i\rightarrow j$ and $j\rightarrow i$).
@@
 In order to answer our research questions regarding the dynamics of topic changes we need to define the concept of a topic.
-In the previous settings, a dataset DSET generates TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$, but, conversely, an existing set of TPGs can generate DSET, akin to generating data from a particular topic, which leads to the following definitions.
+In the previous settings, a dataset DSET generates TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$, but, conversely, an existing set of TPGs can generate DSET.
+Therefore, inspired by Ameisen et al., (2025) that introduces attribution graphs to reveal the LLMs’ internal computational structure, we define a topic as a set of TPGs:
 Definition 2 (topic).
@@
 A sequence $({\bf X},y)$ is within $\mathbb{T}$ if ${\bf X}$ belongs to $\mathbb{T}$ and $\forall{\bf x}\in{\bf X},\ (y\Rightarrow x)\in\mathcal{G}^{(\bar{x})}$.
-We can define a topic as a set of TPGs because it inherently captures the relationships and contextual dependencies between tokens, offering a structured representation of a topic’s semantics.
-This means that we can use a topic $\mathbb{T}$ to generate DSET by sampling sequences $({\bf X},y)$ within $\mathbb{T}$.
+Our graph-based formulation aligns with recent advances in structured representations of LLMs (Sen et al.,, 2023; Wang et al.,, 2025).
 Given the finite number of edges, a DSET can be generated from $\mathbb{T}$ such that it can reconstruct the exact TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ that define $\mathbb{T}$, following the construction method in Li et al., 2024b .
@@
 Given an input sequence ${\bf X}$ that belongs to $\mathbb{T}$, a weight matrix ${\bf W}$ is said to keep topic $\mathbb{T}$ for the input sequence ${\bf X}$ if $\hat{y}_{\bf W}\in\widehat{\mathcal{G}}^{(k)}({\bf X})$.
-This definition establishes the criteria for a next predicted token to preserve the topic of the input sequence.
-Notice that if we have two topics, $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, we can generate two different datasets DSETa and DSETb.
-It is clear that $\tilde{\bf W}_{a}$ [^3] Notation: The subscripts of weights and objects correspond to the associated topic.
+Remark.
+If we have two topics, $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, we can generate two different datasets DSETa and DSETb.
+The union of $\{\mathcal{G}^{(k)}_{a}\}_{k=1}^{K}$ for $\mathbb{T}_{a}$ and $\{\mathcal{G}^{(k)}_{b}\}_{k=1}^{K}$ for $\mathbb{T}_{b}$ forms $\{\mathcal{G}^{(k)}_{ab}\}_{k=1}^{K}$ for the mixed-topics, denoted as the TPGs for the combined DSETa and DSETb.
+It is clear that $\tilde{\bf W}_{a}$ trained only with DSETa will always keep topic $\mathbb{T}_{a}$.
+[^3] Notation: The subscripts of weights and objects correspond to the associated topic.
 For instance $\tilde{\bf W}_{a}$ denotes the weights defined in Equation 2, obtained from DSETa, which pertains to topic $\mathbb{T}_{a}$.
-trained only with DSETa will always keep topic $\mathbb{T}_{a}$.
 But we could also obtain $\tilde{\bf W}_{ab}$ with a dataset combining DSETa and DSETb as training sets.
 The central question is whether $\tilde{\bf W}_{ab}$ keeps topic $\mathbb{T}_{a}$, given an input sequence ${\bf X}$ that belongs to $\mathbb{T}_{a}$, or if it instead predicts tokens that prompt a topic change.
-This can be understood through an analogy: consider an oracle only proficient in Topic A, capable of following conversations in the context of Topic A.
-If this oracle gains knowledge of Topic B, will it continue to follow the conversation within Topic A’s context when the conversation is within Topic A, or will its response begin to exhibit a shift to Topic B?
 4 Attention within Mixed Topics
 Let’s first understand the way in which attention models assign priority to tokens within mixed-topic setting.
-For simplicity, we elaborate our results using a two-topic scenario, but it is straightforward to extend all of our theoretical results on multiple topics.
+For simplicity, we elaborate our results using a two-topic scenario, but it is straightforward to extend the results on multiple topics.
 Notice the self-attention embedding output is a linear combination of $\bf X$ given by $\mathbb{S}({\bf XW\bar{x}})$.
-The embeddings in $\bf X$ corresponding to the highest entries in the vector $\mathbb{S}({\bf XW\bar{x}})$ will receive higher priority in their importance to predict the next token, therefore we can hypothesize that models in which $\mathbb{S}({\bf XW\bar{x}})$ are ordered in a similar way will predict similar next tokens.
+The embeddings in $\bf X$ corresponding to the highest entries in $\mathbb{S}({\bf XW\bar{x}})$ will receive higher priority to predict the next token, therefore we can hypothesize that models in which $\mathbb{S}({\bf XW\bar{x}})$ are ordered in a similar way will predict similar next tokens.
 This idea leads to our first main result which considers this situation within a mixed topics setting:
@@
 This is $\forall i,j\in[T]$:
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Figure 3: Illustration of topic combination.
-$\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ are token-priority graphs (TPGs) of the last token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively.
-$\mathcal{G}^{(4)}_{ab}$ combines the TPG for $\mathbb{T}_{a}$ and the TPG for $\mathbb{T}_{b}$.
+  • $\text{if }\quad[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\text{, then}\quad[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
+  • $\text{if }\quad[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\text{, then}\quad[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\geq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
+  • $\text{if }\quad[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\text{, then}\quad[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\leq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
 This implies that for an input sequence $\mathbf{X}$, a model trained in a mixed-topic setting will maintain the priority of the topic to which $\mathbf{X}$ belongs.
 Consequently, the attention will be allocated in the same order as if the model had been trained exclusively on the original topic of $\mathbf{X}$.
-The input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$, as shown in Figure 4 (top), is generated from $\mathbb{T}_{a}$.
-The predicted next token $\hat{y}_{{\bf w}_{ab}}$ is ${\bf e}_{5}$ and the highest probability SCC in mixed-topics is $\widehat{\mathcal{G}}^{(4)}_{ab}({\bf X})=\{{\bf e}_{5}\}$.
+For the first input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$ from $\mathbb{T}_{a}$, as shown in Figure 3 (right), the predicted next token $\hat{y}_{{\bf w}_{ab}}$ is ${\bf e}_{5}$ and the highest probability SCC in mixed-topics is $\widehat{\mathcal{G}}^{(4)}_{ab}({\bf X})=\{{\bf e}_{5}\}$.
 Since $\hat{y}_{{\bf w}_{ab}}$ belongs to ${\widehat{\mathcal{G}}}^{(4)}_{ab}({\bf X})$, ${\bf W}_{ab}$ for input sequence $\bf X$ is considered as topic continuity, based on the Definition 3.
@@
 However, if $\mathbf{X}$ belongs to $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, the priority will be preserved within both topics.
-Additionally, strict equality in the attention priority is maintained, but strict inequalities may not be when combining both datasets.
-In such cases, the union of their TPGs can form new SCCs, disrupting strict priority.
-As illustrated in Figure 3, we present $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ as the TPGs corresponding to last input token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively.
-In $\mathcal{G}^{(4)}_{a}$, the token priority is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}>{\bf e}_{4}$.
-Once combining $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$, the priority order in the mixed-topics is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}={\bf e}_{4}$, based on $\mathcal{G}^{(4)}_{ab}$.
-The strict equality ${\bf e}_{1}={\bf e}_{2}$ from $\mathcal{G}^{(4)}_{a}$ is maintained in $\mathcal{G}^{(4)}_{ab}$, but the strict inequality ${\bf e}_{2}>{\bf e}_{4}$ is replaced with ${\bf e}_{2}={\bf e}_{4}$ in mixed-topics, forming the new SCC, $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\}$, in $\mathcal{G}^{(4)}_{ab}$.
-Figure 4: Illustration of topic continuity, ambiguous sequence, and change of topic (see Definition 3,  4, and 5 for formal definitions).
-For each input sequence derived from $\mathbb{T}_{a}$, we employ a self-attention model trained on $\mathbb{T}_{a}$ and another model trained on the mixed-topics dataset $\mathbb{T}_{ab}$ to predict the next tokens $\hat{y}_{{\bf w}_{a}}$ and $\hat{y}_{{\bf w}_{ab}}$, respectively (orange).
-$\widehat{\mathcal{G}}^{(4)}_{ab}$ is the highest probability SCC (see Definition 1) in mixed-topics setting and $\widehat{\mathcal{G}}^{(4)}_{a}$ is the highest probability SCC in $\mathbb{T}_{a}$.
-Top: The input sequence is classified as topic continuity if $\hat{y}_{{\bf w}_{ab}}\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$.
-Middle: The input sequence is ambiguous if $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$, but $\widehat{\mathcal{G}}^{(4)}_{a}(X)\subset\widehat{\mathcal{G}}^{(4)}_{ab}(X)$.
-Bottom: ${\bf W}_{ab}$ is considered to change of topic, if $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}^{(4)}_{a}(X)$ and $\widehat{\mathcal{G}}^{(4)}_{a}(X)\not\subset\widehat{\mathcal{G}}^{(4)}_{ab}(X)$.
-The numeric details are provided in Appendix C.3.
-To further illustrate Theorem 3 we define the attention priority similarity of weights ${\bf W}^{\prime}$ relative to ${\bf W}$ for a sequence ${\bf X}$ as: $R_{{\bf W},{\bf W}^{\prime}}({\bf X})=$
-$\frac{1}{T-1}\sum_{j=1}^{T-1}g\left(\left[\mathbb{S}\left({\bf X}{\bf W}^{\prime}\bar{\bf x}\right)\right]_{i_{j}}-\left[\mathbb{S}\left({\bf X}{\bf W}^{\prime}\bar{\bf x}\right)\right]_{i_{j+1}}\right),$
-where $i_{1},\cdots,i_{T}$ is a permutation of $1,\cdots,T$ such that $\left[\mathbb{S}\left({\bf X}{\bf W}\bar{\bf x}\right)\right]_{i_{1}}\geq\cdots\geq\left[\mathbb{S}\left({\bf X}{\bf W}\bar{\bf x}\right)\right]_{i_{T}}$, and
-$g(w)=\begin{cases}1,&\text{if $w\geq 0$,}\\ \frac{1}{e^{-w}},&\text{otherwise.}\end{cases}$
-The attention priority similarity quantifies how well the weights ${\bf W}^{\prime}$ preserve the attention priority of the weights ${\bf W}$.
-A value of 1 indicates that the priority is fully preserved.
-Using this metric, we conduct experiments, with results in Figure 5.
-We generate embeddings with $K=10$ and $d=16$, and randomly construct TPGs for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$.
-Using these TPGs, we randomly generate $\text{DSET}_{a}$ and $\text{DSET}_{b}$.
-We compute $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$, $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ and $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ using the same procedure as Li et al., 2024b .
-We generate test sequences ${\bf Z}$ within $\mathbb{T}_{a}$, and we calculate the attention priority similarity of $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to both $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$.
-We repeat this process for multiple TPGs and input sequences (simulation details in Appendix C).
-Figure 5 clearly demonstrates that the similarity converges to 1 after $\tau=8000$ iterations when evaluated relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue line), but fails to converge relative to $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ (orange line).
-These observations align with the results of Theorem 3.
-Figure 5: Convergence of attention priority similarity for $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue) and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (orange).
+Additionally, strict equality in the attention priority holds, but strict inequalities may not, as the union of their TPGs can form new SCCs, like $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\}$ in $\mathcal{G}^{(4)}_{ab}$ (left side of Figure 3), disrupting strict priority (see details in Appendix C).
+Figure 3: Depiction of each scenario in next token prediction.
+Left: Taking the last token ${\bf e}_{4}$ as an example, $\mathcal{G}^{(4)}_{ab}$ for $\mathbb{T}_{ab}$ is formed by the union of $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$.
+Right: For each input sequence belonging to $\mathbb{T}_{a}$, we employ a self-attention model trained on $\mathbb{T}_{a}$ and another model trained on the mixed-topics dataset $\mathbb{T}_{ab}$ to predict the next tokens, denoted as $\hat{y}_{{\bf w}_{a}}$ and $\hat{y}_{{\bf w}_{ab}}$ (orange).
+$\widehat{\mathcal{G}}^{(4)}_{ab}$ and $\widehat{\mathcal{G}}^{(4)}_{a}$ represent the highest probability SCCs (Definition 1) in mixed-topics setting and in $\mathbb{T}_{a}$, respectively.
+There are three scenarios, topic continuity (Definition 3), ambiguous sequence (Definition 4), and change of topic (Definition 5).
+The numeric details for each scenario are provided in Appendix D.5.
 5 Explaining Topic Shifts
@@
 Definition 4 defines an ambiguous sequence as one where the highest-probability next-token predictions include tokens from both within and outside the input topic, reflecting natural ambiguity from overlapping topics.
-For instance, the input sequence ${\bf X}=[{\bf e}_{1},{\bf e}_{4},{\bf e}_{1},{\bf e}_{4}]^{\top}$ in Figure 4 (middle) belongs to both $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$.
-$\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$ is $\{{\bf e}_{1}\}$, as depicted in $\mathcal{G}_{a}^{(4)}$ from Figure 3 (left) and $\widehat{\mathcal{G}}_{ab}^{(4)}({\bf X})$ is $\{{\bf e}_{1},{\bf e}_{4}\}$, as shown in $\mathcal{G}_{ab}^{(4)}$ from Figure 3 (right).
-$\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$ is the subset of $\widehat{\mathcal{G}}_{ab}^{(4)}({\bf X})$, though $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$.
+Take the second input sequence ${\bf X}=[{\bf e}_{1},{\bf e}_{4},{\bf e}_{1},{\bf e}_{4}]^{\top}$ in Figure 3 (right) as an example.
+$\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$ is $\{{\bf e}_{1}\}$, as depicted in $\mathcal{G}_{a}^{(4)}$ from Figure 3 (left) and $\widehat{\mathcal{G}}_{ab}^{(4)}({\bf X})$ is $\{{\bf e}_{1},{\bf e}_{4}\}$, as shown in $\mathcal{G}_{ab}^{(4)}$ from Figure 3 (left).
+$\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$ is a subset of $\widehat{\mathcal{G}}_{ab}^{(4)}({\bf X})$, although $\hat{y}_{{\bf w}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}({\bf X})$.
 We can argue that the next token predicted from an ambiguous sequence cannot be considered as a topic change, as it lacks the clear trigger phenomenon observed in human cognition.
@@
 The weight matrix $\tilde{\bf W}_{ab}$ changes topic $\mathbb{T}_{a}$ for sequence ${\bf X}$ if $\tilde{\bf W}_{ab}$ does not keep topic $\mathbb{T}_{a}$ for ${\bf X}$ and $\bf X$ is not ambiguous in $\mathbb{T}_{ab}$ with respect to $\mathbb{T}_{a}$.
-In Figure 4, ${\bf W}_{ab}$ for the last input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{4},{\bf e}_{4},{\bf e}_{4}]^{\top}$ can be considered as change of topic.
+In Figure 3 (right), ${\bf W}_{ab}$ changes topic for the last input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{4},{\bf e}_{4},{\bf e}_{4}]^{\top}$, following the Definition 5.
 Building on the formal definitions of topic continuity, ambiguous sequences, and topic changes, we now present a necessary condition for a sequence to induce a topic change.
 This is achieved by introducing our final definition, grounded in the highest-priority SCC as determined by the order in the attention layer.
-Now, our second key result.
 Definition 6 (highest priority SCC).
@@
 Theorem 4.
-Under the same settings and assumptions on datasets and training in Theorem 3, let ${\bf X}$ be a sequence that belongs to $\mathbb{T}_{a}$.
+Under the same settings and assumptions in Theorem 3, let ${\bf X}$ be a sequence that belongs to $\mathbb{T}_{a}$.
 If $\tilde{\bf W}_{ab}$ changes topic $\mathbb{T}_{a}$ for $\bf X$ then $\exists x_{j}\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$ such that $\forall x_{i}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$, the number of times ${\bf x}_{j}$ appears in $\bf X$ is greater than the number of times ${\bf x}_{i}$ appears in $\bf X$.
@@
 A natural question arises: what do these findings imply in practice?
-Specifically, how does the probability of change of topic behaves as the input sequence length increases or as the ambiguity between topics grows?
-The following theorem provides insights into these dynamics.
+Specifically, how does the probability of change of topic behave as the input sequence length or the topic ambiguity increases?
+The following theorem sheds light on these dynamics.
 Theorem 5.
@@
   1. 1.
-    If $p>\max_{x\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})}\mathbb{P}\left({\bf x}_{i}^{\prime}={\bf x}\right)$ then
-    $\mathbb{P}(\tilde{\bf W}_{ab}$ changes topic $\mathbb{T}_{a}$ for sequence ${\bf X}^{\prime})\to 0$ as $T\to\infty.$
+    If $p>\max_{x\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})}\mathbb{P}\left({\bf x}_{i}^{\prime}={\bf x}\right)$,
+    then $\lim_{T\to\infty}{\mathbb{P}(\tilde{\bf W}_{ab}\text{ changes topic }\mathbb{T}_{a}\text{ for }{\bf X}^{\prime})}=0$.
   2. 2.
-    If $l$ increases then the probability that $\exists x_{j}^{\prime}\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$ such that $\forall x_{i}^{\prime}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$, the number of times ${\bf x}_{j}^{\prime}$ appears in $\bf X^{\prime}$ is greater than the number of times ${\bf x}_{i}^{\prime}$ appears in $\bf X^{\prime}$ does not increase.
+    If $l$ increases then the probability that $\exists x_{j}^{\prime}\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$ such that $\forall x_{i}^{\prime}\in\dot{\mathcal{G}}_{a}^{(\bar{x})}(\bf X)$, ${\bf x}_{j}^{\prime}$ outnumbers ${\bf x}_{i}^{\prime}$ in $\bf X^{\prime}$ does not increase.
 There are two implications of this theorem.
@@
 This contrasts with human cognition, where longer conversations and greater inter-connectivity of knowledge increase the likelihood of spontaneous shifts.
-(a) Length of sequences.
-(b) Overlap between topics.
-Figure 6: The proportion of keep topic (topic continuity), ambiguity (ambiguous sequence), and change topic (change of topic) as (a) the input sequence length increases and (b) the overlap between topics increases.
-To illustrate Theorem 5, we conduct simulations following the same implementation settings as described in Section 4.
+To illustrate Theorem 5 with simulations, we generate embeddings with $K=10$ and $d=16$.
 We approximate $\tilde{\mathbf{W}}_{a}$ and $\tilde{\mathbf{W}}_{ab}$ as the results obtained after $\tau=8000$ iterations of Algo-GD.
@@
 We increase $T$ from 4 to 512.
-Figure 6(a) illustrates how the proportion of change topic decreases as $T$ increases.
+Figure 5 illustrates how the proportion of change topic decreases as $T$ increases.
 Second, we investigate the effect of overlapping topics by generating TPGs with an increased number of edges $L$.
@@
 We vary $L$ from 4 to 18.
-Figure 6(b) demonstrates that as $L$ increases, ambiguity increases, while proportion of change topic decreases.
+Figure 5 demonstrates that as $L$ increases, ambiguity increases, while proportion of change topic doesn’t increase.
 These two findings contrast with expectations derived from human cognition but align with the result of Theorem 5.
@@
 The remaining 0.02% mismatched cases were solely due to minor approximation discrepancies in the attention softmax.
-These results validate Theorem 4 (simulation details in Appendix C).
-Related Work
+These results validate Theorem 4 (See simulation details in Appendix D).
+(a) Input length.
+(b) Topic ambiguity.
+Figure 4: The proportion of topic continuity, ambiguity, and topic change as (a) input length and (b) topic ambiguity increase.
+(a) Input length.
+(b) Topic ambiguity.
+Figure 5: Cosine similarity between continuations generated with single-topic and mixed-topic knowledge as (a) input length and (b) topic ambiguity vary.
+6 Experiments in Frontier LLMs
+To prove Theorem 5 we work within the simplified, single-layer self-attention model of Li et al., 2024b .
+Although this abstraction omits many hallmarks of contemporary LLMs (deep stacks of attention blocks, alternative cost functions, and other training heuristics), it offers a mathematically tractable setting that lets us derive interesting mathematical results.
+These results, in turn, can be used to understand how cutting-edge LLMs behave in terms of spontaneous topic changes.
+We empirically investigate such behavior on four frontier models: GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3.
+Real dataset.
+We randomly select 100 arXiv papers published in March 2025 since the publicly disclosed knowledge cutoff dates for our study LLMs fall at the end of 2024 or earlier.
+This ensures that these models have not been trained on these data.
+We consider each paper as a different “topic”.
+Experimental setup.
+For two distinct papers A and B, and an input prompt ($\bf X)$ from paper A, we consider a measure of topic continuity as the cosine similarity between the embeddings of the texts generated when the LLM has contextual knowledge solely from paper A ($\hat{y}_{\bf{W}_{a}}$) and when the LLM has contextual knowledge from both paper A and B ($\hat{y}_{\bf{W}_{ab}}$).
+We consider this cosine similarity as an empirical proxy for our definition of topic continuity (Definition 3): therefore the larger the similarity, the smaller the chance that the model has lead to a change of topic.
+This proxy suggests two testable consequences which become the empirical counterpart of our Theorem 5: Cosine similarity is expected to increase with the length of the input prompt, and is not expected to decrease when increasing ambiguity between paper A and paper B.
+To more closely align with our theoretical framework, where a model gains knowledge of topic A and incrementally gains knowledge of topic B, we implement a Retrieval-Augmented Generation (RAG) approach, retrieving information exclusively from paper A or jointly from papers A and B (Von Oswald et al.,, 2023).
+Based on input prompt, we retrieve the top 3 most relevant excerpts from paper A or B to form the contextual knowledge set A or set B.
+The combined contextual knowledge set is simply the union of sets A and B.
+We add set A to the input prompt to obtain the generated text with sole knowledge of paper A ($\hat{y}_{\bf{W}_{a}}$), and we add the combined set to the input prompt to obtain the generated text with combined knowledge of paper A and B ($\hat{y}_{\bf{W}_{ab}}$).
+To closely follow our greedy decoding approach in our theoretical framework, we set the temperature parameter to 0 for all LLMs.
+We designate each paper as paper A, and we randomly select 5 different papers from the remaining 99 papers as distinct paper B.
+For each input segment, we calculate the average cosine similarity between $\hat{y}_{\bf{W}a}$ and $\hat{y}_{\bf{W}{ab}}$ across these five pairs of paper A and paper B, using each LLM.
+The results for each LLM are averaged over the 100 papers (See more experimental details in Appendix E).
+Experiment 1: Impact of the input length.
+We use the first $10,30,\ldots,150$ words from each paper A’s abstract as the input prompt.
+Figure 5 plots, for each LLM, the average cosine similarity as a function of the input length; the shaded bands indicate 95% confidence intervals.
+Across all models, similarity tends to increase with prompt length, aligning with the behavior predicted by Theorem 5.
+Experiment 2: Impact of the topic ambiguity.
+We fix the input prompt length to the first $80$ words for each paper A’s abstract.
+We quantify paper ambiguity by the average similarity among each paper’s keywords: higher keyword similarity signifies lower paper ambiguity because topic is more specific.
+We partitioned the papers into six equal-width bins along this ambiguity spectrum.
+Figure 5 summarises the results: each boxplot shows the distribution of cosine similarities within an ambiguity bin (x-axis ordered from least to most ambiguous).
+Across all LLMs the median similarity does not seem to decrease, in agreement with the prediction of Theorem 5.
+Taken together, the two experiments provide preliminary empirical support for Theorem 5, showing that its prediction, derived for a single-layer self-attention toy model, can be extended to today’s deep, multi-layer LLMs.
+Crucially, an important divergence between machine and human cognition persists in these frontier models: neither longer prompts nor greater topic ambiguity appreciably raise the likelihood of an spontaneous topic change.
+7 Related Work
 Training and generalization of Transformer.
-(1) Properties of Self-attention and Softmax.
+(1) Properties of Softmax.
 The self-attention mechanism employs the softmax function to selectively emphasize different parts of the input.
@@
 Bombari and Mondelli, (2024) examined the word sensitivity of attention layers, revealing that softmax-based attention-layers are adept at capturing the significance of individual words.
-Despite these strengths, recent researches have highlighted limitations of the softmax function (Saratchandran et al.,, 2024; Deng et al.,, 2023).
+However, recent work has also pointed out limitations of the softmax function (Saratchandran et al.,, 2024; Deng et al.,, 2023).
 (2) Optimization in attention-based models.
 Additionally, recent researches interpret Transformer models as kernel machines, akin to support vector machines (SVMs), with self-attention layers performing maximum margin separation in the token space (Tarzanagh et al., 2023a, ; Tarzanagh et al., 2023b, ; Li et al., 2024b, ; Julistiono et al.,, 2024).
-(3) Chain-of-Thought and In-Context Learning.
-Moreover, transformers exhibit remarkable abilities in generalization through in-context learning (ICL), where models effectively learn from contextual cues during inference (Brown et al.,, 2020; Xie et al.,, 2022; Olsson et al.,, 2022).
-Chain-of-Thought (CoT) prompting (Wei et al.,, 2022; Zhou et al.,, 2023; Shao et al.,, 2023; Li et al., 2024a, ) enhances this by breaking down reasoning processes into intermediate steps, highlighting the emergent reasoning abilities of transformers.
+(3) Chain-of-Thought (CoT) and In-Context Learning (ICL).
+Moreover, transformers exhibit remarkable abilities in generalization through ICL, where models effectively learn from contextual cues during inference (Brown et al.,, 2020; Xie et al.,, 2022; Olsson et al.,, 2022).
+CoT prompting (Wei et al.,, 2022; Zhou et al.,, 2023; Shao et al.,, 2023; Li et al., 2024a, ) enhances this by breaking down reasoning processes into intermediate steps, highlighting the emergent reasoning abilities of transformers.
 (4) Improvement efficiency of transformers.
@@
 Self-Attention and topic dynamics.
-Advancements in the field has significantly advanced our understanding of how self-attention mechanisms in transformers manage evolving semantic contexts, with works have explored various aspects of topic modeling, including dynamic topic structures (Miyamoto et al.,, 2023), hierarchical relationships (Lin et al.,, 2024), topic-aware attention (Panwar et al.,, 2021), and mechanistic analysis of topic structure (Li et al.,, 2023).
-While these works provide insights into managing static and hierarchical topic structures, our work focuses on the topic changes during next-token prediction within the given input sequences from a specific topic.
+Advancements in self-attention research have deepened our understanding of how transformers handle evolving semantic contexts.
+Prior work has explored diverse aspects of topic modeling, such as dynamic topic structures (Miyamoto et al.,, 2023), hierarchical relationships (Lin et al.,, 2024), topic-aware attention mechanisms (Panwar et al.,, 2021), and the mechanistic underpinnings of topic representation (Li et al.,, 2023).
+While these studies provide insights into managing static and hierarchical topic structures, our work focuses on the topic changes with the given input sequences from a specific topic.
 8 Discussion
-Our exploration with a simple single-layer attention model reveals fundamental clues on the distinctions between human and AI conversational dynamics.
-While humans naturally shift topics through curiosity or contextual triggers, the model remains anchored to the context.
-These observations suggest that human-like spontaneity in topic changes demands mechanisms beyond pattern recognition, raising deeper questions about the nature of interaction and highlighting the importance of creativity in advancing conversational AI.
-However, our analysis relies on the simplified single-layer self-attention setting from Li et al., 2024b .
-This does not directly extend to complex modern LLMs with deeper architectures and varied training goals.
-Future work will explore how these findings translate to more intricate models, examining their ability to change topics.
-Moreover, the recurring theme of topic continuity characterized in this work opens new avenues for applications, such as automatic change topic detection in conversational AI based on discrepancies between model predictions and speaker outputs.
+Our theoretical analysis on self-attention models and empirical investigations on modern LLMs reveal fundamental clues regarding the distinctions between model-based spontaneous topic changes and human spontaneous thought, a phenomenon that is critical for comparing conversational dynamics across humans and AI.
+In an era of growing concern about AI’s cognitive resemblance to humans, our framework provides preliminary results differentiating these phenomena, thereby opening pathways for future interdisciplinary research at the interface of artificial and human cognition.
+Limitations.
+Our theoretical framework builds on the same simplified single-layer self-attention model with a log-loss objective from Li et al., 2024b and defining topics as TPGs.
+These abstractions do not fully capture the complexities of contemporary LLMs, including deep attention architectures, alternative loss functions, and diverse training objectives.
+Despite loosening these assumptions, our experiments suggest that the essence of our core theoretical conclusions holds across modern LLMs within our framework of study.
+Future work will investigate how broadly these theoretical insights generalize to complex LLM architectures.
 Appendix A Technical Proofs
-A.1 Proofs of Lemmas
-A.1.1 Proof of Lemma 2
+A.1 Proof of Lemma 2
 Let ${\bf a}=\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})$.
@@
 If $x_{i}\neq x_{j}$ then $x_{i}$ and $x_{j}$ are in the same SCC.∎
-A.1.2 Proof of Lemma 6
+A.2 Proof of Lemma 6
 Lemma 6.
@@
   • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}<1\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$.
-A.1.3 Proof of Lemma 7
+A.3 Proof of Lemma 7
 Lemma 7.
@@
 ∎
-A.2 Proof of Theorem 3
+A.4 Proof of Theorem 3
 From construction, $\forall k\in[K]$, $\mathcal{G}_{a}^{(k)}\subseteq\mathcal{G}_{ab}^{(k)}$. This means that $\forall{\bf x}_{i},{\bf x}_{j}\in{\bf X}$, we have:
@@
   • $[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}_{a}^{(\bar{x})}$ then $(x_{i}\Rightarrow x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\geq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
-A.3 Proof of Theorem 4
+A.5 Proof of Theorem 4
 Let ${\bf a}=\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})$ and ${\bf b}=\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})$.
@@
 Then $\tilde{\bf W}_{ab}$ keeps topic $\mathbb{T}_{a}$ for input sequence $\bf X$, or $\bf X$ is ambiguous in $\mathbb{T}_{ab}$ with respect to $\mathbb{T}_{a}$.∎
-A.4 Proof of Theorem 5
+A.6 Proof of Theorem 5
   1. 1.
@@
 Figure 6: Illustration of Assumption 4.
-Here are two datasets related to the TPGs, $\mathcal{G}_{a}^{(4)}$ and $\mathcal{G}_{b}^{(4)}$, from Figure 3.
+Here are two datasets related to the TPGs, $\mathcal{G}_{a}^{(4)}$ and $\mathcal{G}_{b}^{(4)}$, from Figure 3 (left).
 From the directed arrows in $\mathcal{G}_{a}^{(4)}$, we can generate a dataset with the last token $\bf{e}_{4}$ for $\mathbb{T}_{a}$, which can reconstruct back the $\mathcal{G}_{a}^{(4)}$.
 A similar process applies for the $\mathcal{G}_{b}^{(4)}$.
-Appendix C Detailed Simulation Studies
-C.1 Simulation Process
+Appendix C Explanation of TPGs Combination
+Figure 3 presents $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ as the TPGs corresponding to the last input token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively.
+In $\mathcal{G}^{(4)}_{a}$, the token priority is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}>{\bf e}_{4}$.
+In $\mathcal{G}^{(4)}_{ab}$ for the mixed-topics, the priority order is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}={\bf e}_{4}$.
+The strict equality ${\bf e}_{1}={\bf e}_{2}$ from $\mathcal{G}^{(4)}_{a}$ is maintained in $\mathcal{G}^{(4)}_{ab}$, but the strict inequality ${\bf e}_{2}>{\bf e}_{4}$ is replaced with ${\bf e}_{2}={\bf e}_{4}$ in mixed-topics, forming the new SCC, $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\}$, in $\mathcal{G}^{(4)}_{ab}$.
+Appendix D Detailed Simulation Studies with Single-layer Self-attention
+D.1 Simulation Process
 Theoretical TPGs generation.
@@
 Our code for all simulations is available in ( ).
-C.2 Experiments Results
-In this section, we provide empirical evidence to support our findings outlined in Theorem 3, Theorem 4, and Theorem 5, offering detailed insights into the behavior of self-attention mechanisms for changing topics.
-C.2.1 Simulation in Section 4
-For Figure 5 we used $K=10,L=4,d=16,\text{and }T_{test}=T_{train}=4$, we also showed point-wise confidence intervals of rank similarities on all 5000 instances for each value of $\tau$.
-C.2.2 Simulation in Section 5
-Table 1: The proportion of keep topic, ambiguous, and change of topic with varying $T_{test}=\{4,8,16,24,32,64,128,256,512\}$.
+D.2 Additional Experiments to Support Theorem 3
+To further illustrate Theorem 3 we define the attention priority similarity of weights ${\bf W}^{\prime}$ relative to ${\bf W}$ for a sequence ${\bf X}$ as: $R_{{\bf W},{\bf W}^{\prime}}({\bf X})=$
+$\frac{1}{T-1}\sum_{j=1}^{T-1}g\left(\left[\mathbb{S}\left({\bf X}{\bf W}^{\prime}\bar{\bf x}\right)\right]_{i_{j}}-\left[\mathbb{S}\left({\bf X}{\bf W}^{\prime}\bar{\bf x}\right)\right]_{i_{j+1}}\right),$
+where $i_{1},\cdots,i_{T}$ is a permutation of $1,\cdots,T$ such that $\left[\mathbb{S}\left({\bf X}{\bf W}\bar{\bf x}\right)\right]_{i_{1}}\geq\cdots\geq\left[\mathbb{S}\left({\bf X}{\bf W}\bar{\bf x}\right)\right]_{i_{T}}$, and
+$g(w)=\begin{cases}1,&\text{if $w\geq 0$,}\\ \frac{1}{e^{-w}},&\text{otherwise.}\end{cases}$
+The attention priority similarity quantifies how well the weights ${\bf W}^{\prime}$ preserve the attention priority of the weights ${\bf W}$.
+A value of 1 indicates that the priority is fully preserved.
+Using this metric, we conduct experiments, with results in Figure 7.
+We generate embeddings with $K=10$ and $d=16$, and randomly construct TPGs for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$.
+Using these TPGs, we randomly generate $\text{DSET}_{a}$ and $\text{DSET}_{b}$.
+We compute $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$, $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ and $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ using the same procedure as Li et al., 2024b .
+We generate test sequences ${\bf Z}$ within $\mathbb{T}_{a}$, and we calculate the attention priority similarity of $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to both $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$.
+We repeat this process for multiple TPGs and input sequences (simulation details in Appendix D).
+Figure 7 clearly demonstrates that the similarity converges to 1 after $\tau=8000$ iterations when evaluated relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue line), but fails to converge relative to $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ (orange line).
+These observations align with the results of Theorem 3.
+Figure 7: Convergence of attention priority similarity for $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue) and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ (orange).
+D.3 Simulation in Section 5
+In Figure 5(a), we predict next tokens for 5000 test sequences from $\mathbb{T}_{a}$ with $T_{test}=\{4,8,16,24,32,64,128,256,512\}$, while fixing $L=4$, $d=16$, $T_{train}=4$, and $K=10$.
+The proportion of each scenario with varying $T$ is illustrated in Table 1.
+For Figure 5(b), we predict next tokens for 5000 test sequences (the sequence length is $T_{test}=20$) using models trained with $L=\{4,6,8,10,12,14,16,18\}$, $d=16$, $K=10$, and $T_{train}=4$.
+The proportion of each scenario with varying $L$ is illustrated in Table 2.
+Table 1: Proportion of keep topic, ambiguous, and change of topic with varying $T_{test}=\{4,8,16,24,32,64,128,256,512\}$.
 $T_{test}$ Keep(%) Ambiguous(%) Change(%)
@@
 512 100.00 $\pm$ 0.00 0.00 $\pm$ 0.00 0.00 $\pm$ 0.00
-Table 2: The proportion of keep topic, ambiguous, and change of topic with varying $L=\{4,6,8,10,12,14,16,18\}$.
+Table 2: Proportion of keep topic, ambiguous, and change of topic with varying $L=\{4,6,8,10,12,14,16,18\}$.
 $L$ Keep(%) Ambiguous(%) Change(%)
@@
 18 98.02 $\pm$ 1.58 1.26 $\pm$ 1.14 0.72 $\pm$ 0.86
-In Figure 6(a), we predict next tokens for 5000 test sequences from $\mathbb{T}_{a}$ with $T_{test}=\{4,8,16,24,32,64,128,256,512\}$, while fixing $L=4$, $d=16$, $T_{train}=4$, and $K=10$.
-The proportion of each scenario with varying $T$ is illustrated in Table 1.
-For Figure 6(b), we predict next tokens for 5000 test sequences (the sequence length is $T_{test}=20$) using models trained with $L=\{4,6,8,10,12,14,16,18\}$, $d=16$, $K=10$, and $T_{train}=4$.
-The proportion of each scenario with varying $L$ is illustrated in Table 2.
-C.2.3 Additional experiments
+D.4 Additional Experiments for Convergence in Mixed-topics
 (a) $K=6,L=4$
@@
 The shaded area around each line represents the $95\%$ confidence interval, computed over 50 epochs.
-C.3 Numerical Analysis of each Scenario in Figure 4
-(a) Topic continuity.
-(b) Ambiguous sequence.
-(c) Change of topic.
-Figure 9: Numeric details for each scenario: (a) topic continuity, (b) ambiguous sequence, and (c) change of topic.
+D.5 Numerical Analysis for Each Scenario in Figure 3
 Figure 9 provides a numerical breakdown for each scenario in Figure 3.
@@
 Following Equation 9 and Equation 10, we compute $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{x_{i}}$ and $[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{x_{i}}$ to get the highest probability SCC and predict the next token for each input sequence.
+(a) Topic continuity.
+(b) Ambiguous sequence.
+(c) Change of topic.
+Figure 9: Numeric details for each scenario: (a) topic continuity, (b) ambiguous sequence, and (c) change of topic.
 Topic continuity.
@@
 9(a), input sequence $\bf X$ consists of four unique tokens: ${\bf e}_{5}$, ${\bf e}_{1}$, ${\bf e}_{3}$, and ${\bf e}_{4}$.
-Based on $\mathcal{G}^{(4)}_{a}$ in Figure 3, the priority order of these tokens is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}>{\bf e}_{4}$, with corresponding $a_{i}$ values: $0.45>0.25>0.20>0.1$.
+Based on $\mathcal{G}^{(4)}_{a}$ in Figure 3 (left), the priority order of these tokens is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}>{\bf e}_{4}$, with corresponding $a_{i}$ values: $0.45>0.25>0.20>0.1$.
 Since $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.45$ is the largest, $\widehat{\mathcal{G}}_{a}^{(4)}=\{{\bf e}_{5}\}$ and $\hat{y}_{{\bf W}_{a}}={\bf e}_{5}$.
@@
 9(b) has two unique tokens: ${\bf e}_{1}$ and ${\bf e}_{4}$.
-The priority order is ${\bf e}_{1}>{\bf e}_{4}$, following $\mathcal{G}^{(4)}_{a}$ in Figure 3.
+The priority order is ${\bf e}_{1}>{\bf e}_{4}$, following $\mathcal{G}^{(4)}_{a}$ in Figure 3 (left).
 The corresponding values are $a_{1}=a_{3}=0.3$ and $a_{2}=a_{4}=0.2$.
@@
 Thus, $\widehat{\mathcal{G}}_{a}^{(4)}$ is $\{{\bf e}_{5}\}$ with the highest probability.
-${\bf W}_{ab}$ makes ${\bf e}_{4}$ and ${\bf e}_{1}$ with the same priority, as indicated by $\mathcal{G}^{(4)}_{ab}$ in Figure 3.
+${\bf W}_{ab}$ makes ${\bf e}_{4}$ and ${\bf e}_{1}$ with the same priority, as indicated by $\mathcal{G}^{(4)}_{ab}$ in Figure 3 (left).
 Both ${\bf e}_{1}$ and ${\bf e}_{4}$ are within the highest probability SCC, $\widehat{\mathcal{G}}_{ab}^{(4)}$, due to $[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{1}}=[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.25$.
@@
 For the input sequence $\bf X$ in Fig.
-9(c), the only two unique tokens, ${\bf e}_{5}$ and ${\bf e}_{4}$, are with the same priority order in both $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{ab}$ from Figure 3: ${\bf e}_{5}>{\bf e}_{1}$.
+9(c), the only two unique tokens, ${\bf e}_{5}$ and ${\bf e}_{4}$, are with the same priority order in both $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{ab}$ from Figure 3 (left): ${\bf e}_{5}>{\bf e}_{1}$.
 With ${\bf W}_{a}(\tau)$ trained in $\mathbb{T}_{a}$, the token ${\bf e}_{5}$ has $a_{1}=0.70$ and the token ${\bf e}_{4}$ has $a_{2}=a_{3}=a_{4}=0.10$.
@@
 Since $\hat{y}_{{\bf W}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}$ and $\widehat{\mathcal{G}}_{a}^{(4)}\not\subset\widehat{\mathcal{G}}_{ab}^{(4)}$, ${\bf W}_{ab}$ changes topic for the input sequence $\bf X$.
-Moreover, we have $({\bf e}_{5}\Rightarrow{\bf e}_{i})\in\mathcal{G}^{(4)}_{ab}$ for $i\in[4]$, as shown in Figure 3.
+Moreover, we have $({\bf e}_{5}\Rightarrow{\bf e}_{i})\in\mathcal{G}^{(4)}_{ab}$ for $i\in[4]$, as shown in Figure 3 (left).
 Thus, the highest priority SCC (Definition 1) in $\mathbb{T}_{ab}$ is $\dot{\mathcal{G}}^{(4)}_{ab}({\bf X})=\{{\bf e}_{5}\}$.
 In the input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{4},{\bf e}_{4},{\bf e}_{4}]^{\top}$, the lower-priority token ${\bf e}_{4}\not\in\dot{\mathcal{G}}^{(4)}_{ab}({\bf X})$ appears more frequently than the higher-priority token ${\bf e}_{5}\in\dot{\mathcal{G}}^{(4)}_{ab}({\bf X})$, illustrating our Theorem 4.
+Appendix E Experimental Details in Section 6
+In this section, we provide the experimental details in four LLMs: GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3.
+Here, we outline the general procedure used in each model, under identical parameter settings, to generate continuations for each segment of the abstract as follows:
+  1. 1.
+    Extract the first $T$ words from paper $A$’s abstract as the input segment $\bf{X}$ from Topic $A$.
+  2. 2.
+    Randomly select 5 different papers from paper $A$, as papers in $\{B_{i}\}_{i=1}^{5}$.
+  3. 3.
+    For the input segment $\bf{X}$, apply RAG to extract top 3 relevant excerpts (chunks) from paper $A$ as the the knowledge $A$, denoted as $\text{Ref}_{A}$. Each chunk is with 800 tokens.
+  4. 4.
+    Similarly, retrieve top 3 relevant excerpts from paper $B_{i}$ as the knowledge $B_{i}$, denoted as $\text{Ref}_{B_{i}}$.
+  5. 5.
+    Combine the knowledge from Topic $A$ and from Topic $B_{i}$ as the knowledge $AB_{i}$ for mixed-Topic, denoted as $\text{Ref}_{AB_{i}}$.
+  6. 6.
+    For the input segment $\bf{X}$, each LLM with the follow prompts, $\text{Prompt}_{A}$ and $\text{Prompt}_{AB_{i}}$, to generate the continuations as $\hat{y}_{\bf{W}_{a}}$ and $\hat{y}_{\bf{W}_{ab}}$, respectively.
+Notably, the only difference between $\text{Prompt}_{A}$ and $\text{Prompt}_{AB_{i}}$ is the reference excerpts provided $\text{Ref}_{A}$ or $\text{Ref}_{AB_{i}}$.
+All LLMs were queried with a temperature of 0 to match the greedy decoding in our theoretical framework.
+The maximum completion length was set to 1000 tokens to ensure that the generated continuations could complete the abstract.
+    1. (a)
+      $\text{Prompt}_{A}$:
+      Here are some relevant excerpts from research paper(s) as reference:$\text{Ref}_{A}$.
+Below is the 1st fragment of an abstract from arXiv paper $A$: $\bf{X}$.
+Please continue the 2nd fragment of the abstract based on the relevant excerpts without including the given content in the output.
+    2. (b)
+      $\text{Prompt}_{AB_{i}}$:
+      Here are some relevant excerpts from research paper(s) as reference:$\text{Ref}_{AB_{i}}$.
+Below is the 1st fragment of an abstract from arXiv paper $A$: $\bf{X}$.
+Please continue the 2nd fragment of the abstract based on the relevant excerpts without including the given content in the output.
+  7. 7.
+    Calculate the average cosine similarity between $\hat{y}_{\bf{W}a}$ and $\hat{y}_{\bf{W}{ab}}$ across five pairs of paper $A$ and paper $B_{i}$.
+E.1 Impact of the Input Length
+To investigate the impact of the input length, we vary $T=\{10,30,50,70,90,110,130,150\}$ for every paper as Topic $A$, increasing the length of the input segment $\bf X$, as shown on the x-axis from Figure 5.
+E.2 Impact of the Topic Ambiguity
+We quantify topic (paper) ambiguity by the average similarity among each paper’s keywords.
+As arXiv papers do not provide keywords, we use Llama-3.3 to generate four keywords for each paper prior to generating continuations with the LLMs.
+To investigate the topic ambiguity, we fix the input length with $T=80$ for every paper as Topic $A$ and order papers by the average keywords similarity, as shown on the x-axis of Figure 5.
+Higher keywords similarity corresponds to lower topic ambiguity.
+Appendix F Computational Resources for Experiments
+In our simulations based on the single-layer self-attention model, each group of parameter setting requires 7 hours to train two models separately, one for single input topic and one for mixed-topic, followed by 2 additional hours for next-token prediction.
+In our experiments on LLMs, we query GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3 through API calls.
+All experiments were conducted on a standard laptop without specialized hardware.
+For each LLM, the full process, including selecting relevant excerpts using RAG and generating continuations, requires approximately 20 hours of runtime, with a total of 13 million input tokens and 2.5 million output tokens.
+The total API usage cost for the experiments is approximately 200 USD.
+Appendix G Impact Statement
+Our investigation highlights fundamental differences between spontaneous topic changes in LLMs and spontaneous human thought, informing the development of more natural and flexible AI systems in domains such as customer service and mental health support.
+However, improving such capabilities can raise ethical considerations, including inadvertent manipulation of user focus, especially in persuasive or sensitive contexts.
+Our work, while largely theoretical, emphasizes the importance of fairness, privacy, and user autonomy as developers refine these systems to serve users’ interests, respect contextual boundaries, and remain accountable.
+This research has the potential to advance both Machine Learning and Human-Computer Interaction by informing new architectures that mimic human-like topic shifts; nevertheless, any real-world application of these findings should be accompanied by vigilant oversight to mitigate risks of misuse—such as deceptive or manipulative dialogue shifting.
+There are many other potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v4

@@
 Human cognition is punctuated by abrupt, spontaneous shifts between topics—driven by emotional, contextual, or associative cues—a phenomenon known as spontaneous thought in neuroscience.
-In contrast, self-attention based models depend on structured patterns over their inputs to predict each next token, lacking spontaneity.
-Motivated by this distinction, we characterize spontaneous topic changes in self-attention architectures, revealing both their similarities and their divergences from spontaneous human thought.
-First, we establish theoretical results under a simplified, single-layer self-attention model with suitable conditions by defining the topic as a set of Token Priority Graphs (TPGs).
-Specifically, we demonstrate that (1) the model maintains the priority order of tokens related to the input topic, (2) a spontaneous topic change can occur only if lower-priority tokens outnumber all higher-priority tokens of the input topic, and (3) unlike human cognition, the longer context length or the more ambiguous input topic reduces the likelihood of spontaneous change.
-Second, we empirically validate that these dynamics persist in modern, state-of-the-art LLMs, underscoring a fundamental disparity between human cognition and AI behaviour in the context of spontaneous topic changes.
-To the best of our knowledge, no prior work has explored these questions with a focus as closely aligned to human thought.
+In contrast, self-attention-based models rely on structured patterns over their inputs to predict each next token, lacking spontaneity.
+Motivated by this distinction, we characterize spontaneous topic changes in self-attention architectures and reveal divergences from spontaneous human thought.
+First, we establish theoretical results under a simplified, single-layer self-attention model with suitable conditions by defining a topic as a set of Token Priority Graphs (TPGs).
+Specifically, we demonstrate that (1) the model maintains the priority order of tokens related to the input topic, (2) a spontaneous topic change can occur only if lower-priority tokens outnumber all higher-priority tokens of the input topic, and (3) unlike human cognition, the longer context length or the more ambiguous input topic does not increase the likelihood of spontaneous change.
+Second, we empirically validate that the effect of input length or topic ambiguity persists in modern, state-of-the-art LLMs, underscoring a fundamental disparity between human cognition and AI behavior in the context of spontaneous topic changes.
+To the best of our knowledge, no prior work has explored these questions with a focus so closely aligned to human thought.
 1 Introduction
-Human cognition is punctuated by abrupt, apparently unstructured topic changes, the hallmark of human spontaneous thought, a phenomenon that has become a central topic in cognitive neuroscience (Bellana et al.,, 2022; Christoff and Fox,, 2018; Christoff et al.,, 2011; Kucyi et al.,, 2023; Mildner and Tamir,, 2019, 2024; Mills et al.,, 2020).
+Human cognition is punctuated by abrupt, apparently unstructured topic changes, the hallmark of spontaneous human thought, a phenomenon that has become a central topic in cognitive neuroscience [Bellana2022, spontaneousthought, Christoff2011, Kucyi2023, MILDNER2019763, judith2024why, outofblue].
 For example, a spontaneous shift in focus during a conversation, a sudden leap between ideas when brainstorming, or an unexpected redirection in storytelling.
-These abrupt changes may be due to an emotional connection, such as recalling reading that book during a family vacation, where sensory details like the scent of the ocean or the warmth of the sun trigger a vivid memory.
-However, LLMs rely on contextual signals in the input to perform topic shifts, instead of spontaneous topic changes.
+These abrupt changes may be due to an emotional connection, such as recalling reading a book during a family vacation, where sensory details like the scent of the ocean or the warmth of the sun trigger a vivid memory.
+However, LLMs shift topics in response to contextual cues in the input, rather than initiating spontaneous topic changes on their own.
 They follow a structured, statistical approach, remaining on topic unless explicit cues signal a change.
-Figure 2 illustrates this distinction using the first sentence of the book “One hundred years of solitude” (García Márquez,, 1967).
-Our work takes initial steps toward formalizing the dynamics of spontaneous topic changes in LLMs and analyzing how they relate to or diverge from human spontaneous thought.
-To this end, we ground our theoretical analysis in a single-layer self-attention model and empirically extend it to modern LLMs, laying groundwork for drawing parallels between AI models and human cognition.
+Figure 2 illustrates this distinction using the first sentence of the book “One hundred years of solitude” [marquez1967].
+Our work takes initial steps toward formalizing the dynamics of spontaneous topic changes in LLMs and analyzing how they relate to or diverge from spontaneous human thought.
+To this end, we ground our theoretical analysis in a single-layer self-attention model and empirically extend it to modern LLMs, laying the groundwork for drawing comparisons between AI models and human cognition.
 Recent advancements in the related field have substantially deepened our understanding of self-attention architectures.
-(Li et al., 2024b, ; Tarzanagh et al., 2023a, ; Tarzanagh et al., 2023b, ) have linked the self-attention to support vector machines (SVMs), offering optimization strategies for next-token prediction.
-A study by Li et al., (2023) highlights that in mixed-topics inputs, transformers achieve higher pairwise attention between same-topic words compared to different-topic words.
-In parallel, prior studies have recognized the practical challenges of spontaneous topic changes in LLMs and proposed various approaches to address them (Hwang et al.,, 2024; Lim et al.,, 2010; Lin et al.,, 2023; Ni et al.,, 2022; Soni et al.,, 2022; Xie et al.,, 2021).
-Notably, spontaneous topic changes must be differentiated from hallucinations, generating incorrect or fabricated information without clear contextual basis (Ji et al.,, 2023; Maynez et al.,, 2020).
+li2024mechanics, tarzanagh2023transformers, ataee2023max have linked the self-attention to support vector machines (SVMs), offering optimization strategies for next-token prediction.
+li2023transformers highlight that in mixed-topic inputs, transformers achieve higher pairwise attention between same-topic words compared to different-topic words.
+In parallel, prior studies have recognized the practical challenges of spontaneous topic changes in LLMs and proposed approaches to address them [hwang-etal-2024-mp2d, Lim2010AST, multi-granularity2023, Ni2021RecentAI, soni-etal-2022-empirical, xie-etal-2021-tiage-benchmark].
+Notably, spontaneous topic changes must be differentiated from hallucinations, generating incorrect or fabricated information without a clear contextual basis [ji2023survey, maynez-etal-2020-faithfulness].
 Despite these advancements, our understanding of the dynamics of spontaneous topic changes in LLMs remains limited.
-Investigating the parallels and differences between spontaneous topic changes in self-attention models and human spontaneous thought could provide valuable insights for improving the efficiency of language models with human-like thought.
-Since modern LLMs rely on self-attention architectures, we begin by theoretically characterizing spontaneous topic shifts in a simplified setting.
+Investigating the relationship between spontaneous topic changes in self-attention models and spontaneous human thought can provide valuable insights into the cognitive discrepancies of current language models compared with humans.
+Since modern LLMs rely on self-attention architectures, we begin by theoretically characterizing spontaneous topic changes in a simplified setting.
 We then extend these findings through experiments on more complex, state-of-the-art models.
@@
 Figure 2 outlines our theoretical framework.
-To make the mathematical analysis tractable, we follow the same single-layer self-attention framework with log-loss objective function governed by Assumptions 14 from Li et al., 2024b .
-Inspired by token-priority graphs (TPGs) (Li et al., 2024b, ) and building on attribution graphs from Ameisen et al., (2025) for exposing an LLM’s internal computation, we define a topic as a set of TPGs.
-This graph-based formulation aligns naturally with recent advances in structured representations for LLMs (Sen et al.,, 2023; Wang et al.,, 2025).
-Furthermore, this mirrors neuroscience models of human spontaneous thought, in which concepts serve as nodes connected by associative edges (Mildner and Tamir,, 2019).
+To make the mathematical analysis tractable, we follow the same single-layer self-attention framework with log-loss objective function governed by Assumptions 14 from li2024mechanics.
+Inspired by token-priority graphs (TPGs) [li2024mechanics] and building on attribution graphs from ameisen2025circuit for exposing an LLM’s internal computation, we define a topic as a set of TPGs.
+This graph-based formulation aligns naturally with recent advances in structured representations for LLMs [sen-etal-2023-knowledge, wang2025model].
+Furthermore, this mirrors neuroscience models of spontaneous human thought, in which concepts serve as nodes connected by associative edges [MILDNER2019763].
 Despite relying on these specific settings, our experiments extend our findings to modern LLMs, empirically confirming that relaxing these assumptions does not seem to undermine our core insights.
 Figure 1: Illustration of the difference between human cognition and LLMs.
-The original fragment of “One hundred years of solitude" (García Márquez,, 1967) (top) has a clear spontaneous thought, but the GPT-2’s completion (bottom), demonstrates continuity.
+The original fragment of “One hundred years of solitude" [marquez1967] (top) has a clear spontaneous thought, but the GPT-2’s completion (bottom), demonstrates continuity.
 [^2] Just to illustrate, we use the prompt Please continue this short sentence, forgetting about “One hundred Years of Solitude”, since on a real conversation the LLM would be blind to the final output.
 Figure 2: Overview of our theoretical framework.
-We define a topic by a set of TPGs $\{\mathcal{G}^{(k)}\}^{K}_{k=1}$ (Def.
+We define a topic as a set of TPGs $\{\mathcal{G}^{(k)}\}^{K}_{k=1}$ (Def.
 2) and generate a dataset for each topic.
@@
 The next-token prediction with the mixed-topic model is categorized into three outcomes: keeps topic A (topic continuity from Def.
-3), ambiguous sequence (from Def.
-4), changes topic A (change of topic from Def.
+3); ambiguous sequence (from Def.
+4); or changes topic A (change of topic from Def.
 5).
-Further details for each category are in Figure 3.
-Summary of Findings
+Further details for each category are shown in Figure 3.
+1.1 Summary of findings
 Imagine an oracle that is an expert on Topic A, capable of following any conversation within that topic while staying true to its context.
@@
     Preservation of input topic priorities.
-Using a controlled sandbox, we demonstrate in Theorem 3 that attention models trained on mixed-topic datasets maintain the priorities of tokens associated with the original topic of an input sequence (Topic A in our analogy).
+Using a controlled sandbox, we demonstrate in Theorem 2 that self-attention models trained on mixed-topic datasets maintain the priorities of tokens associated with the original topic of an input sequence (Topic A in our analogy).
   2. 2.
     Changing topics triggered by token frequency.
-In Theorem 4, we show that only if a lower-priority token appears more frequently than all higher-priority tokens of Topic A, the oracle’s responses may reflect a change of topic.
+In Theorem 3, we show that the oracle’s responses may reflect a change of topic only if a lower-priority token appears more frequently than all higher-priority tokens of Topic A.
   3. 3.
-    Impact of sequence length and topic ambiguity.
+    Impact of input length and topic ambiguity.
 Theorem 4 establishes that longer input sequences decrease the likelihood of changing topics.
-Furthermore, topic ambiguity acts as a stabilizing factor, not increasing the frequency of spontaneous topic changes.
+Furthermore, input topic ambiguity acts as a stabilizing factor, not increasing the frequency of spontaneous topic changes.
   4. 4.
     Difference between LLMs and human cognition.
-In Section 6 we empirically extend Theorem 5 to modern deeper LLMs.
-Unlike human cognition, where extended discussions often encourage spontaneous thoughts and topic ambiguity promotes cognitive connections, our results highlight an opposite behavior in LLMs: neither longer prompts nor greater topic ambiguity appreciably raise the likelihood of an spontaneous topic change.
+In Section 6 we empirically extend Theorem 4 to modern, deeper LLMs.
+Unlike human cognition, where extended discussions often encourage spontaneous thoughts and topic ambiguity promotes cognitive connections, our results highlight the opposite behavior in LLMs: neither longer prompts nor greater topic ambiguity appreciably increases the likelihood of a spontaneous topic change.
 Overview of the paper structure.
 We begin with the problem setup in Sec 2.
-Sec 3 introduces the definition of topic, and Sec 4 examines how self-attention models allocate the token priorities within the mixed-topics.
+Sec 3 introduces the definition of topic, and Sec 4 examines how self-attention models allocate the token priorities within the mixed topics.
 In Sec 5, we establish the conditions under which a self-attention model induces spontaneous topic changes and show the dynamics of topic changes with longer input sequences or the presence of topic ambiguity.
@@
 2.1 Next topic prediction with self-attention model
-In line with the approach presented by Tarzanagh et al., 2023b and Li et al., 2024b , we frame the next-token prediction task as a multi-class classification problem.
+In line with the approach presented by ataee2023max and li2024mechanics, we frame the next-token prediction task as a multi-class classification problem.
 Given a vocabulary of size $K$ with an embedding matrix ${\bf E}=[{\bf e}_{1}\ {\bf e}_{2}\ \cdots\ {\bf e}_{K}]^{\top}\in\mathbb{R}^{K\times d}$, we aim to predict the next token ID $y\in[K]$ based on an input sequence ${\bf X}=[{\bf x}_{1}\ {\bf x}_{2}\ \cdots\ {\bf x}_{T}]^{\top}\in\mathbb{R}^{T\times d}$ with ${\bf x}_{i}\in{\bf E}$ for all $i\in[T]$.
@@
 In our notation $\bf x$} is the embedding vector corresponding to the token ID $x$, this is ${\bf x}={\bf e}_{x}$.
-For prediction, we utilize a single-layer self-attention model with a combined key-query weight matrix ${\bf W}\in\mathbb{R}^{d\times d}$ and identity value matrix as in Tarzanagh et al., 2023b .
+For prediction, we utilize a single-layer self-attention model with a combined key-query weight matrix ${\bf W}\in\mathbb{R}^{d\times d}$ and identity value matrix as in ataee2023max.
 The self-attention embedding output
@@
 ${\bf W}^{(\tau+1)}={\bf W}^{(\tau)}-\eta\nabla L({\bf W}^{(\tau)}).$ (Algo-GD)
-We keep the first two assumptions from Li et al., 2024b :
+We keep the first two assumptions from li2024mechanics:
 Assumption 1.
@@
 For any $({\bf X},y)\in\text{DSET}$, the token ${\bf e}_{y}$ is contained in the input sequence $\bf X$}.
-Assumption 1 represents a variation of the weight-tying approach commonly used in language models (Press and Wolf,, 2017; Vaswani et al.,, 2017).
+Assumption 1 represents a variation of the weight-tying approach commonly used in language models [press2016using, vaswani2017attention].
 Once training is complete, for a new input sequence $\bf X$}, and a model characterized by $\bf W$, we predict the next token ID $\hat{y}_{\bf w}$ based on greedy decoding the probabilities from the softmax of the classification output
@@
 2.2 Token-priority graph and global convergence of the self-attention model
-Li et al., 2024b defined a token-priority graph (TPG) as a directed graph with nodes representing tokens in the vocabulary.
+li2024mechanics defined a token-priority graph (TPG) as a directed graph with nodes representing tokens in the vocabulary.
 $\text{DSET}^{(k)}$ is a subset of sequences from DSET with the same last token is ${\bf e}_{k}=\bar{\bf x}$.
@@
 For tokens within two different SCCs, strict priority orders emerge, helping the model to differentiate between tokens when learning next-token predictions.
-We use the same notation as Li et al., 2024b , given a directed graph $\mathcal{G}$, for $i,j\in[K]$ such that $i\neq j$:
+We use the same notation as li2024mechanics, given a directed graph $\mathcal{G}$, for $i,j\in[K]$ such that $i\neq j$:
   • $i\in\mathcal{G}$ denotes that the node $i$ belongs to $\mathcal{G}$.
@@
 Nodes in each $\mathcal{G}^{(k)}$ represent indices in $[K]$, and SCC structure supports the self-attention mechanism’s ability to assign priority within sequences based on the conditioning last token.
-Theorem 2 of Li et al., 2024b proved that under Assumptions 1 and 2, the self-attention model learned through Algo-GD converges to the solution of the following Support Vector Machine (SVM) defined by the TPGs of the underlying dataset DSET
+Theorem 2 of li2024mechanics proved that under Assumptions 1 and 2, the self-attention model learned through Algo-GD converges to the solution of the following Support Vector Machine (SVM) defined by the TPGs of the underlying dataset DSET
 ${\bf W}^{\text{svm}}=\arg\min_{\bf W}\|{\bf W}\|_{F}$ (Graph-SVM)
-[…]
+$\text{s.t.}\quad({\bf e}_{i}-{\bf e}_{j})^{\top}{\bf W}{\bf e}_{k}\begin{cases}=0,&\forall(i\asymp j)\in\mathcal{G}^{(k)}\\ \geq 1,&\forall(i\Rightarrow j)\in\mathcal{G}^{(k)}\end{cases}\forall k\in[K].$
 Here is a condensed version of the theorem:
-Theorem 1 (Li et al., 2024b ).
+Theorem 1 (li2024mechanics).
 Consider dataset DSET and suppose Assumptions 1 and 2 hold.
@@
 This convergence implies that the model predicts the next token based on priorities obtained from the SCCs within the TPG relevant to the last token of the input sequence.
-Unlike the work in Li et al., 2024b , which considers multiple loss functions in subsequent results, we focus exclusively on a log-loss function in this work, leaving the exploration of other loss functions for future research.
-We add here another reasonable assumption that prevents the probabilities in Equation 1 to be equal for improbable numerical reasons, and we present our first lemma.
+Unlike the work in li2024mechanics, which considers both hard retrieval and soft composition components and examines multiple loss functions in subsequent results, we focus exclusively on a log-loss function in this work, leaving the exploration of other loss functions for future research.
+Since the soft composition component is not required for our subsequent definitions and theoretical results, we concentrate solely on the hard retrieval component.
+We add here another reasonable assumption that prevents the probabilities in Equation 1 from being equal due to improbable numerical reasons, and we present our first lemma.
 Assumption 3.
@@
 Suppose conditions from Theorem 1 and Assumption 3 hold.
-Consider an input sequence ${\bf X}$ from DSET(k) and corresponding TPG $\mathcal{G}^{(k)}$, $\forall i,j\in[K]$ we have $\left[\mathbb{S}\left({\bf C}f_{\bf W}({\bf X})\right)\right]_{i}=\left[\mathbb{S}\left({\bf C}f_{\tilde{\bf W}}({\bf X})\right)\right]_{j}\text{ iff }(x_{i}\asymp x_{j})\in\mathcal{G}^{(k)}$.
-This means that the tokens that maximize the probability in Equation 1 are all within the same SCC leading to the following definition:
+Consider an input sequence $\bf X$} from DSET(k) and corresponding TPG $\mathcal{G}^{(k)}$, $\forall i,j\in[K]$ we have $\left[\mathbb{S}\left({\bf C}f_{\tilde{\bf W}}({\bf X})\right)\right]_{i}=\left[\mathbb{S}\left({\bf C}f_{\tilde{\bf W}}({\bf X})\right)\right]_{j}\text{ iff }(x_{i}\asymp x_{j})\in\mathcal{G}^{(k)}$.
+This means that the tokens that maximize the probability for weights $\tilde{\bf W}$ in Equation 1 are all within the same SCC leading to the following definition:
 Definition 1 (highest probability SCC).
@@
 In the previous settings, a dataset DSET generates TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$, but, conversely, an existing set of TPGs can generate DSET.
-Therefore, inspired by Ameisen et al., (2025) that introduces attribution graphs to reveal the LLMs’ internal computational structure, we define a topic as a set of TPGs:
+Therefore, inspired by ameisen2025circuit that introduces attribution graphs to reveal the LLMs’ internal computational structure, we define a topic as a set of TPGs:
 Definition 2 (topic).
@@
 A sequence $({\bf X},y)$ is within $\mathbb{T}$ if $\bf X$} belongs to $\mathbb{T}$ and $\forall{\bf x}\in{\bf X},\ (y\Rightarrow x)\in\mathcal{G}^{(\bar{x})}$.
-Our graph-based formulation aligns with recent advances in structured representations of LLMs (Sen et al.,, 2023; Wang et al.,, 2025).
-Given the finite number of edges, a DSET can be generated from $\mathbb{T}$ such that it can reconstruct the exact TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ that define $\mathbb{T}$, following the construction method in Li et al., 2024b .
+Our graph-based formulation aligns with recent advances in structured representations of LLMs [sen-etal-2023-knowledge, wang2025model].
+Given the finite number of edges, a DSET can be generated from $\mathbb{T}$ such that it can reconstruct the exact TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ that define $\mathbb{T}$, following the construction method in li2024mechanics.
 This leads to the following reasonable assumption:
@@
 A DSET generated from any topic $\mathbb{T}$ defined by $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$ exactly reconstructs back the TPGs $\{\mathcal{G}^{(k)}\}_{k=1}^{K}$.
-Detailed explanation is provided in Appendix B. This assumption enables the application of the results from Li et al., 2024b , with the concepts of topics and TPGs being used interchangeably.
+Detailed explanation is provided in Appendix B. This assumption enables the application of the results from li2024mechanics, with the concepts of topics and TPGs being used interchangeably.
 Definition 3 (topic continuity).
@@
 Remark.
-If we have two topics, $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, we can generate two different datasets DSETa and DSETb.
-The union of $\{\mathcal{G}^{(k)}_{a}\}_{k=1}^{K}$ for $\mathbb{T}_{a}$ and $\{\mathcal{G}^{(k)}_{b}\}_{k=1}^{K}$ for $\mathbb{T}_{b}$ forms $\{\mathcal{G}^{(k)}_{ab}\}_{k=1}^{K}$ for the mixed-topics, denoted as the TPGs for the combined DSETa and DSETb.
+Given two topics, $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, with corresponding datasets DSETa and DSETb, the union of $\{\mathcal{G}^{(k)}_{a}\}_{k=1}^{K}$ and $\{\mathcal{G}^{(k)}_{b}\}_{k=1}^{K}$ forms the TPGs for the mixed topics $\mathbb{T}_{ab}$, denoted by $\{\mathcal{G}^{(k)}_{ab}\}_{k=1}^{K}$.
 It is clear that $\tilde{\bf W}_{a}$ trained only with DSETa will always keep topic $\mathbb{T}_{a}$.
@@
 4 Attention within mixed topics
-Let’s first understand the way in which attention models assign priority to tokens within mixed-topic setting.
+Let’s first understand how attention models assign priority to tokens within mixed-topic setting.
 For simplicity, we elaborate our results using a two-topic scenario, but it is straightforward to extend the results on multiple topics.
@@
 The embeddings in $\bf X$ corresponding to the highest entries in $\mathbb{S}({\bf XW\bar{x}})$ will receive higher priority to predict the next token, therefore we can hypothesize that models in which $\mathbb{S}({\bf XW\bar{x}})$ are ordered in a similar way will predict similar next tokens.
-This idea leads to our first main result which considers this situation within a mixed topics setting:
+This idea leads to our first main result which considers this situation within a mixed-topic setting:
 Theorem 2.
@@
 Consequently, the attention will be allocated in the same order as if the model had been trained exclusively on the original topic of $\mathbf{X}$.
-For the first input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$ from $\mathbb{T}_{a}$, as shown in Figure 3 (right), the predicted next token $\hat{y}_{{\bf w}_{ab}}$ is ${\bf e}_{5}$ and the highest probability SCC in mixed-topics is $\widehat{\mathcal{G}}^{(4)}_{ab}({\bf X})=\{{\bf e}_{5}\}$.
+For the first input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$ from $\mathbb{T}_{a}$, as shown in Figure 3 (right), the predicted next token $\hat{y}_{{\bf w}_{ab}}$ is ${\bf e}_{5}$ and the highest probability SCC in mixed topics is $\widehat{\mathcal{G}}^{(4)}_{ab}({\bf X})=\{{\bf e}_{5}\}$.
 Since $\hat{y}_{{\bf w}_{ab}}$ belongs to ${\widehat{\mathcal{G}}}^{(4)}_{ab}({\bf X})$, ${\bf W}_{ab}$ for input sequence $\bf X$ is considered as topic continuity, based on the Definition 3.
@@
 However, if $\mathbf{X}$ belongs to $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, the priority will be preserved within both topics.
-Additionally, strict equality in the attention priority holds, but strict inequalities may not, as the union of their TPGs can form new SCCs, like $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\}$ in $\mathcal{G}^{(4)}_{ab}$ (left side of Figure 3), disrupting strict priority (see details in Appendix C).
+Additionally, strict equality in the attention priority holds, but strict inequalities may not, as the union of their TPGs can form new SCCs.
+As illustrated on the left of Figure 3, $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ denote the TPGs corresponding to the last input token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively.
+In $\mathcal{G}^{(4)}_{a}$, the token priority is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}>{\bf e}_{4}$.
+In contrast, in $\mathcal{G}^{(4)}_{ab}$ for the mixed topics, the priority order is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}={\bf e}_{4}$.
+The equality ${\bf e}_{1}={\bf e}_{2}$ from $\mathcal{G}^{(4)}_{a}$ is maintained in $\mathcal{G}^{(4)}_{ab}$, whereas the strict inequality ${\bf e}_{2}>{\bf e}_{4}$ is relaxed to ${\bf e}_{2}={\bf e}_{4}$ in mixed topics, forming the new SCC, $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\}$, in $\mathcal{G}^{(4)}_{ab}$.
 Figure 3: Depiction of each scenario in next token prediction.
 Left: Taking the last token ${\bf e}_{4}$ as an example, $\mathcal{G}^{(4)}_{ab}$ for $\mathbb{T}_{ab}$ is formed by the union of $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$.
-Right: For each input sequence belonging to $\mathbb{T}_{a}$, we employ a self-attention model trained on $\mathbb{T}_{a}$ and another model trained on the mixed-topics dataset $\mathbb{T}_{ab}$ to predict the next tokens, denoted as $\hat{y}_{{\bf w}_{a}}$ and $\hat{y}_{{\bf w}_{ab}}$ (orange).
-$\widehat{\mathcal{G}}^{(4)}_{ab}$ and $\widehat{\mathcal{G}}^{(4)}_{a}$ represent the highest probability SCCs (Definition 1) in mixed-topics setting and in $\mathbb{T}_{a}$, respectively.
+The direction of edge is from output to input and the dotted square denotes the strongly-connected components (SCC) in which tokens have equal priority.
+Right: For each input sequence belonging to $\mathbb{T}_{a}$, we use a self-attention model trained on $\text{DSET}_{a}$ and another model trained on the mixed-topic dataset $\text{DSET}_{ab}$ to predict the next tokens, denoted as $\hat{y}_{{\bf w}_{a}}$ and $\hat{y}_{{\bf w}_{ab}}$, respectively.
+$\widehat{\mathcal{G}}^{(4)}_{ab}$ and $\widehat{\mathcal{G}}^{(4)}_{a}$ represent the highest probability SCCs (Definition 1) in mixed-topic setting and in $\mathbb{T}_{a}$, respectively.
 There are three scenarios, topic continuity (Definition 3), ambiguous sequence (Definition 4), and change of topic (Definition 5).
-The numeric details for each scenario are provided in Appendix D.5.
+The numeric details for each scenario are provided in Appendix C.5.
 5 Explaining topic shifts
@@
 This also suggests that topic shifts may arise from ambiguity within an input sequence rather than a straightforward change in topic.
-In our analogy on the oracle, gaining knowledge of both Topic A and Topic B might cause a conversation to be naturally followed within Topic A or also outside Topic A.
+In our oracle analogy, gaining knowledge of both Topic A and Topic B might cause a conversation to be naturally followed within Topic A or also outside Topic A.
 We introduce the following definition to characterize this phenomenon:
@@
   1. 1.
-    If $p>\max_{x\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})}\mathbb{P}\left({\bf x}_{i}^{\prime}={\bf x}\right)$,
-    then $\lim_{T\to\infty}{\mathbb{P}(\tilde{\bf W}_{ab}\text{ changes topic }\mathbb{T}_{a}\text{ for }{\bf X}^{\prime})}=0$.
+    If $p>\max_{x\not\in\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})}\mathbb{P}\left({\bf x}_{i}^{\prime}={\bf x}\right)$, then $\lim_{T\to\infty}{\mathbb{P}(\tilde{\bf W}_{ab}\text{ changes topic }\mathbb{T}_{a}\text{ for }{\bf X}^{\prime})}=0$.
   2. 2.
@@
 First, as the input sequence length increases sufficiently, the likelihood of topic changes vanishes.
-Second, notice that increasing $l$ increases the probability of overlap between topics, and it also decreases the probability of satisfying the necessary conditions for a topic change, effectively creating a bound on the frequency of topic changes.
+Second, increasing $l$ raises the probability of overlap between topics and reduces the probability of satisfying the necessary conditions for a topic change, effectively creating a bound on the frequency of topic changes.
 In practice, consider the oracle analogy: if the oracle is following a sufficiently long conversation on a specific topic, it becomes exceedingly unlikely to shift topics.
 Similarly, as topics A and B become more interconnected, this increased ambiguity does not lead to more topic changes; rather, it may reduce their occurrence.
-This contrasts with human cognition, where longer conversations and greater inter-connectivity of knowledge increase the likelihood of spontaneous shifts.
-To illustrate Theorem 5 with simulations, we generate embeddings with $K=10$ and $d=16$.
+This contrasts with human cognition, where longer conversations and greater inter-connectivity of knowledge increase the likelihood of spontaneous topic changes.
+To illustrate Theorem 4 through simulations, we generate embeddings with $K=10$ and $d=16$.
 We approximate $\tilde{\mathbf{W}}_{a}$ and $\tilde{\mathbf{W}}_{ab}$ as the results obtained after $\tau=8000$ iterations of Algo-GD.
-We quantify the proportion of test sequences in which $\tilde{{\bf W}}_{ab}$ keeps $\mathbb{T}_{a}$ (keep topic), proportion of ambiguous sequences in $\mathbb{T}_{ab}$ (ambiguity) and proportion of times in which $\tilde{{\bf W}}_{ab}$ changes topic (change topic).
+We quantify the proportion of test sequences in which $\tilde{{\bf W}}_{ab}$ keeps $\mathbb{T}_{a}$ (keep topic), proportion of ambiguous sequences in $\mathbb{T}_{ab}$ (ambiguity) and proportion in which $\tilde{{\bf W}}_{ab}$ changes topic (change topic).
 First, we explore the effect of longer sequences by varying the length $T$ of the test sequences $\bf Z$.
@@
 Figure 4 illustrates how the proportion of change topic decreases as $T$ increases.
-Second, we investigate the effect of overlapping topics by generating TPGs with an increased number of edges $L$.
+Second, we investigate the effect of topic overlap with an increasing number of edges $L$.
 Intuitively, a higher $L$ results in an increase $l$ and a greater overlap between TPGs of different topics.
 We vary $L$ from 4 to 18.
-Figure 5 demonstrates that as $L$ increases, ambiguity increases, while proportion of change topic doesn’t increase.
+Figure 4 demonstrates that as $L$ increases, ambiguity increases, while the proportion of change topic doesn’t increase.
 These two findings contrast with expectations derived from human cognition but align with the result of Theorem 4.
-Lastly, among the 85000 test sequences generated for these experiments, 99.98% satisfy Theorem 4 (i.e., topic changes occur when a low-priority token appears more frequently than high-priority tokens).
-The remaining 0.02% mismatched cases were solely due to minor approximation discrepancies in the attention softmax.
-These results validate Theorem 4 (See simulation details in Appendix D).
-(a) Input length.
+Lastly, among the 85,000 test sequences generated for these experiments, 99.98% satisfy Theorem 3 (i.e., topic changes occur when a low-priority token appears more frequently than high-priority tokens).
+The remaining 0.02% mismatched cases are solely due to minor approximation discrepancies in the attention softmax.
+These results validate Theorem 3 (see simulation details in Appendix C).
+(a) Length of sequences.
 (b) Topic ambiguity.
-Figure 4: The proportion of topic continuity, ambiguity, and topic change as (a) input length and (b) topic ambiguity increase.
-(a) Input length.
-(b) Topic ambiguity.
-Figure 5: Cosine similarity between continuations generated with single-topic and mixed-topic knowledge as (a) input length and (b) topic ambiguity vary.
+Figure 4: The proportion of topic continuity, ambiguous sequence, and change of topic as (a) input length and (b) topic ambiguity increase.
 6 Experiments in frontier LLMs
-To prove Theorem 5 we work within the simplified, single-layer self-attention model of Li et al., 2024b .
+To prove Theorem 4 we work within the simplified, single-layer self-attention model of li2024mechanics.
 Although this abstraction omits many hallmarks of contemporary LLMs (deep stacks of attention blocks, alternative cost functions, and other training heuristics), it offers a mathematically tractable setting that lets us derive interesting mathematical results.
@@
 For two distinct papers A and B, and an input prompt ($\bf X)$ from paper A, we consider a measure of topic continuity as the cosine similarity between the embeddings of the texts generated when the LLM has contextual knowledge solely from paper A ($\hat{y}_{\bf{W}_{a}}$) and when the LLM has contextual knowledge from both paper A and B ($\hat{y}_{\bf{W}_{ab}}$).
-We consider this cosine similarity as an empirical proxy for our definition of topic continuity (Definition 3): therefore the larger the similarity, the smaller the chance that the model has lead to a change of topic.
-This proxy suggests two testable consequences which become the empirical counterpart of our Theorem 5: Cosine similarity is expected to increase with the length of the input prompt, and is not expected to decrease when increasing ambiguity between paper A and paper B.
-To more closely align with our theoretical framework, where a model gains knowledge of topic A and incrementally gains knowledge of topic B, we implement a Retrieval-Augmented Generation (RAG) approach, retrieving information exclusively from paper A or jointly from papers A and B (Von Oswald et al.,, 2023).
-Based on input prompt, we retrieve the top 3 most relevant excerpts from paper A or B to form the contextual knowledge set A or set B.
+We treat this cosine similarity as an empirical proxy for our formal definition of topic continuity (Definition 3): therefore the larger the similarity, the smaller the chance that the model has led to a change of topic.
+This proxy suggests two testable consequences which become the empirical counterpart of our Theorem 4: (1) cosine similarity is expected to increase with the length of the input prompt, and (2) it is not expected to decrease with increasing ambiguity in paper A and paper B.
+To more closely align with our theoretical framework, where a model gains knowledge of topic A and incrementally gains knowledge of topic B, we implement a Retrieval-Augmented Generation (RAG) approach, retrieving information exclusively from paper A or jointly from papers A and B [von2023transformers].
+Based on the input prompt, we retrieve the top 3 most relevant excerpts from paper A or paper B to form the contextual knowledge set A or set B.
 The combined contextual knowledge set is simply the union of sets A and B.
@@
 To closely follow our greedy decoding approach in our theoretical framework, we set the temperature parameter to 0 for all LLMs.
-We designate each paper as paper A, and we randomly select 5 different papers from the remaining 99 papers as distinct paper B.
+We designate each paper as paper A and randomly select 5 different papers from the remaining 99 papers as distinct paper B.
 For each input segment, we calculate the average cosine similarity between $\hat{y}_{\bf{W}a}$ and $\hat{y}_{\bf{W}{ab}}$ across these five pairs of paper A and paper B, using each LLM.
-The results for each LLM are averaged over the 100 papers (See more experimental details in Appendix E).
-Experiment 1: Impact of the input length.
+The results for each LLM are averaged over all 100 papers.
+See additional experimental details in Appendix D.
+Experiment 1: Impact of input length.
 We use the first $10,30,\ldots,150$ words from each paper A’s abstract as the input prompt.
-Figure 5 plots, for each LLM, the average cosine similarity as a function of the input length; the shaded bands indicate 95% confidence intervals.
-Across all models, similarity tends to increase with prompt length, aligning with the behavior predicted by Theorem 5.
-Experiment 2: Impact of the topic ambiguity.
-We fix the input prompt length to the first $80$ words for each paper A’s abstract.
-We quantify paper ambiguity by the average similarity among each paper’s keywords: higher keyword similarity signifies lower paper ambiguity because topic is more specific.
-We partitioned the papers into six equal-width bins along this ambiguity spectrum.
-Figure 5 summarises the results: each boxplot shows the distribution of cosine similarities within an ambiguity bin (x-axis ordered from least to most ambiguous).
+Figure 5 shows, for each LLM, the average cosine similarity as a function of input length; shaded bands indicate 95% confidence intervals.
+Across all models, similarity tends to increase with input length, aligning with the behavior predicted by Theorem 4.
+Appendix D.3.1 presents an additional experiment in which we extend the input length to 1210 words extracted from each paper’s introduction; the results further support our conclusions.
+(a) Length of sequences.
+(b) Topic ambiguity.
+Figure 5: Similarity between continuations generated with single-topic and mixed-topic knowledge as (a) input length and (b) topic ambiguity increase.
+Experiment 2: Impact of topic ambiguity.
+We fix the input prompt length to the first $80$ words of each paper A’s abstract.
+We quantify topic ambiguity by the average similarity among each paper A’s keywords: lower keyword similarity signifies higher probability of overlap between paper A and other papers, consistent with our setup in Theorem 4.
+We partition the papers into five equal-width bins along this ambiguity spectrum.
+Figure 5 summarizes the results: each boxplot shows the distribution of cosine similarities within an ambiguity bin, with the x-axis ordered from least to most ambiguous.
 Across all LLMs the median similarity does not seem to decrease, in agreement with the prediction of Theorem 4.
-Taken together, the two experiments provide preliminary empirical support for Theorem 5, showing that its prediction, derived for a single-layer self-attention toy model, can be extended to today’s deep, multi-layer LLMs.
-Crucially, an important divergence between machine and human cognition persists in these frontier models: neither longer prompts nor greater topic ambiguity appreciably raise the likelihood of an spontaneous topic change.
+In Appendix D.3.2, we present an additional experiment using an alternative ambiguity measure based on cross-paper keyword similarity, yielding results consistent with our conclusions.
+Taken together, the two experiments provide preliminary empirical support for Theorem 4, showing that its prediction, derived from a single-layer self-attention toy model, can be extended to today’s deep, multi-layer LLMs.
+Crucially, an important divergence between machine and human cognition persists in these frontier models: neither longer prompts nor greater topic ambiguity appreciably increases the likelihood of a spontaneous topic change.
 7 Related work
@@
 The self-attention mechanism employs the softmax function to selectively emphasize different parts of the input.
-Gu et al., (2024), Goodfellow et al., (2016), and Deng et al., (2023) underscore the pivotal role of the softmax function in shaping attention distributions, influencing how models process and prioritize information within input sequences.
-Bombari and Mondelli, (2024) examined the word sensitivity of attention layers, revealing that softmax-based attention-layers are adept at capturing the significance of individual words.
-However, recent work has also pointed out limitations of the softmax function (Saratchandran et al.,, 2024; Deng et al.,, 2023).
+gu2024exploring, Goodfellow-et-al-2016, and deng2023superiority underscore the pivotal role of the softmax function in shaping attention distributions, influencing how models process and prioritize information within input sequences.
+bombari2024towards examined the word sensitivity of attention layers, revealing that softmax-based attention layers are adept at capturing the significance of individual words.
+However, recent work has also pointed out limitations of the softmax function [saratchandran2024rethinking, deng2023superiority].
 (2) Optimization in attention-based models.
-Additionally, recent researches interpret Transformer models as kernel machines, akin to support vector machines (SVMs), with self-attention layers performing maximum margin separation in the token space (Tarzanagh et al., 2023a, ; Tarzanagh et al., 2023b, ; Li et al., 2024b, ; Julistiono et al.,, 2024).
+Additionally, recent research interprets Transformer models as kernel machines, akin to support vector machines (SVMs), with self-attention layers performing maximum margin separation in the token space [tarzanagh2023transformers, ataee2023max, li2024mechanics, julistiono2024optimizing].
 (3) Chain-of-Thought (CoT) and In-Context Learning (ICL).
-Moreover, transformers exhibit remarkable abilities in generalization through ICL, where models effectively learn from contextual cues during inference (Brown et al.,, 2020; Xie et al.,, 2022; Olsson et al.,, 2022).
-CoT prompting (Wei et al.,, 2022; Zhou et al.,, 2023; Shao et al.,, 2023; Li et al., 2024a, ) enhances this by breaking down reasoning processes into intermediate steps, highlighting the emergent reasoning abilities of transformers.
+Moreover, transformers exhibit remarkable abilities in generalization through ICL, where models effectively learn from contextual cues during inference [brown2020language, xie2021explanation, olsson2022context].
+CoT prompting [wei2022chain, zhou2022least, pmlr-v202-shao23a, li2024training] enhances this by breaking down reasoning processes into intermediate steps, highlighting the emergent reasoning abilities of transformers.
 (4) Improvement efficiency of transformers.
-Recent advancements aim to improve the computational efficiency of transformers (Kitaev et al.,, 2020; Choromanski et al.,, 2021; Sukhbaatar et al.,, 2019; Wang et al.,, 2020), ensuring their viability for large-scale deployment while maintaining or enhancing their representational capabilities.
+Recent advancements aim to improve the computational efficiency of transformers [Kitaev2020Reformer, choromanski2021rethinking, sukhbaatar-etal-2019-adaptive, wang2020linformer], ensuring their viability for large-scale deployment while maintaining or enhancing their representational capabilities.
 Next token prediction in LLMs.
 (1) Theoretical and architectural innovations.
-Shannon, (1951)’s foundational work laid the groundwork for estimating the predictability of natural language sequences, providing a basis for subsequent advances in language modeling.
-Recent studies have expanded our understanding of how LLMs anticipate future tokens from internal hidden states, offering valuable insights into the efficiency and effectiveness of transformer-based architectures (He and Su,, 2024; Pal et al.,, 2023; Shlegeris et al.,, 2024).
+shannon1950prediction’s foundational work laid the groundwork for estimating the predictability of natural language sequences, providing a basis for subsequent advances in language modeling.
+Recent studies have expanded our understanding of how LLMs anticipate future tokens from internal hidden states, offering valuable insights into the efficiency and effectiveness of Transformer-based architectures [he2024law, pal2023future, shlegeris2024languagemodelsbetterhumans].
 Despite their impressive predictive capabilities, these models face fundamental limitations.
-For instance, Bachmann and Nagarajan, (2024) highlights the shortcomings of teacher-forced training, emphasizing how this approach can fail and suggesting strategies to improve model robustness.
+For instance, pmlr-v235-bachmann24a highlights the shortcomings of teacher-forced training, emphasizing how this approach can fail and suggesting strategies to improve model robustness.
 (2) Efficiency and Optimization.
-Goyal et al., (2024) introduces a novel method that incorporates a deliberate computation step before output generation, enhancing reasoning capabilities.
-Additionally, Gloeckle et al., (2024) advocates for multi-token prediction, which significantly improves both efficiency and speed.
+goyal2024think introduces a novel method that incorporates a deliberate computation step before output generation, enhancing reasoning capabilities.
+Additionally, gloeckle2024better advocates for multi-token prediction, which significantly improves both efficiency and speed.
 Self-Attention and topic dynamics.
 Advancements in self-attention research have deepened our understanding of how transformers handle evolving semantic contexts.
-Prior work has explored diverse aspects of topic modeling, such as dynamic topic structures (Miyamoto et al.,, 2023), hierarchical relationships (Lin et al.,, 2024), topic-aware attention mechanisms (Panwar et al.,, 2021), and the mechanistic underpinnings of topic representation (Li et al.,, 2023).
+Prior work has explored diverse aspects of topic modeling, such as dynamic topic structures [miyamoto-etal-2023-dynamic], hierarchical relationships [lin-etal-2024-hierarchical], topic-aware attention mechanisms [panwar-etal-2021-tan], and the mechanistic underpinnings of topic representation [li2023transformers].
 While these studies provide insights into managing static and hierarchical topic structures, our work focuses on the topic changes with the given input sequences from a specific topic.
 8 Discussion
-Our theoretical analysis on self-attention models and empirical investigations on modern LLMs reveal fundamental clues regarding the distinctions between model-based spontaneous topic changes and human spontaneous thought, a phenomenon that is critical for comparing conversational dynamics across humans and AI.
+Our theoretical analysis on self-attention models and empirical investigations on modern LLMs reveal fundamental clues regarding the distinctions between model-based spontaneous topic changes and spontaneous human thought, a phenomenon that is critical for comparing conversational dynamics across humans and AI.
 In an era of growing concern about AI’s cognitive resemblance to humans, our framework provides preliminary results differentiating these phenomena, thereby opening pathways for future interdisciplinary research at the interface of artificial and human cognition.
 Limitations.
-Our theoretical framework builds on the same simplified single-layer self-attention model with a log-loss objective from Li et al., 2024b and defining topics as TPGs.
+Our theoretical framework builds on the same simplified single-layer self-attention model with a log-loss objective from li2024mechanics and defines topics as TPGs.
 These abstractions do not fully capture the complexities of contemporary LLMs, including deep attention architectures, alternative loss functions, and diverse training objectives.
 Despite loosening these assumptions, our experiments suggest that the essence of our core theoretical conclusions holds across modern LLMs within our framework of study.
-Future work will investigate how broadly these theoretical insights generalize to complex LLM architectures.
+Future work will investigate how broadly these theoretical insights generalize to complex architectures, for example within longer context windows and/or LLM outputs.
+Code.
+Acknowledgments and Disclosure of Funding
+This work was supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) under grant DGECR-2022-04531.
+The authors thank Prof.
+Yingcong Li, the anonymous reviewers, and the session chair for their valuable feedback and insightful suggestions, which greatly improved the quality of this work.
+Supplementary Materials
 Appendix A Technical proofs
@@
 $={\bf C}\left({\bf X}^{\top}{\bf a}\right)$ (4)
-$=\begin{bmatrix}\sum_{i=1}^{T}a_{i}\,({\bf c}_{1}^{\top}\cdot{\bf x}_{i})\\ \sum_{i=1}^{T}a_{i}\,({\bf c}_{2}^{\top}\cdot{\bf x}_{i})\\ \vdots\\ \sum_{i=1}^{T}a_{i}\,({\bf c}_{K}^{\top}\cdot{\bf x}_{i})\end{bmatrix}.$ (5)
+$=\begin{bmatrix}\sum_{i=1}^{T}a_{i}\,({\bf c}_{1}^{\top}\cdot{\bf x}_{i})\\ \sum_{i=1}^{T}a_{i}\,({\bf c}_{2}^{\top}\cdot{\bf x}_{i})\\ \vdots\\ \sum_{i=1}^{T}a_{i}\,({\bf c}_{K}^{\top}\cdot{\bf x}_{i})\end{bmatrix}$.
+(5)
 Let $k_{i}$ be the number of times token ${\bf x}_{i}$ appears in $\bf X$}. Then,
@@
 For an input sequence $\bf X$} that belongs to $\mathbb{T}$ and $i,j\in[T]$,
-  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{i}\asymp x_{j})\in\mathcal{G}^{(\bar{x})}$ or $i=j$.
-  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}^{(\bar{x})}$.
-  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$.
+  • $\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{i}\asymp x_{j})\in\mathcal{G}^{(\bar{x})}$ or $i=j$.
+  • $\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}^{(\bar{x})}$.
+  • $\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$.
 Proof.
-Since $\bf X$ belongs to $\mathbb{T}$, $\forall{\bf x}\in{\bf X}$ we have $x\in\mathcal{G}^{(\bar{x})}$, therefore from the construction of TPGs by Li et al., 2024b , for every ${\bf x}_{i},{\bf x}_{j}\in{\bf X}$ we have one of the these relationships: $(x_{i}\Rightarrow x_{j})$, $(x_{j}\Rightarrow x_{i})$, $(x_{i}\asymp x_{j})$ or $x_{i}=x_{j}$.
+Since $\bf X$ belongs to $\mathbb{T}$, $\forall{\bf x}\in{\bf X}$ we have $x\in\mathcal{G}^{(\bar{x})}$, therefore from the construction of TPGs by li2024mechanics, for every ${\bf x}_{i},{\bf x}_{j}\in{\bf X}$ we have one of the these relationships: $(x_{i}\Rightarrow x_{j})$, $(x_{j}\Rightarrow x_{i})$, $(x_{i}\asymp x_{j})$ or $x_{i}=x_{j}$.
 From the constraints in Algo-GD:
-  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}=0\iff(x_{j}\asymp x_{i})\in\mathcal{G}^{(\bar{x})}\text{ or }i=j$.
-  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}>1\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}^{(\bar{x})}$.
-  • $[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}<1\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$.
+[…]
+  • $\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}>1\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}^{(\bar{x})}$.
+  • $\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}\bar{{\bf x}})]_{j}\iff({\bf x}_{i}-{\bf x}_{j})^{\top}\tilde{\bf W}{\bar{x}}<1\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}^{(\bar{x})}$.
 A.3 Proof of Lemma 3
@@
 Combining with Lemma 2:
-  • $[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{i}\asymp x_{j})\in\mathcal{G}_{a}^{(\bar{x})}$ or $i=j$, then $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $i=j\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
-  • $[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}_{a}^{(\bar{x})}$ then $(x_{j}\Rightarrow x_{i})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\leq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
-  • $[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}_{a}^{(\bar{x})}$ then $(x_{i}\Rightarrow x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\geq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
+  • $\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{i}\asymp x_{j})\in\mathcal{G}_{a}^{(\bar{x})}$ or $i=j$, then $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $i=j\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}=[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
+  • $\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}<[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{j}\Rightarrow x_{i})\in\mathcal{G}_{a}^{(\bar{x})}$ then $(x_{j}\Rightarrow x_{i})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\leq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
+  • $\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{i}>[\mathbb{S}({\bf X}\tilde{\bf W}_{a}\bar{{\bf x}})]_{j}\iff(x_{i}\Rightarrow x_{j})\in\mathcal{G}_{a}^{(\bar{x})}$ then $(x_{i}\Rightarrow x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}$ or $(x_{i}\asymp x_{j})\in\mathcal{G}_{ab}^{(\bar{x})}\iff[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{i}\geq[\mathbb{S}({\bf X}\tilde{\bf W}_{ab}\bar{{\bf x}})]_{j}$
 A.5 Proof of Theorem 3
@@
     Let ${\bf X}_{l}^{\prime}=[{\bf x}_{1,l}^{\prime}\ {\bf x}_{2,l}^{\prime}\ \cdots\ {\bf x}_{T,l}^{\prime}]^{\top}$ be a random sequences generated as described in the theorem, where the size of $\dot{\mathcal{G}}_{a}^{(\bar{x})}({\bf X})$ is $l$.
-Let $k_{i,l}$ be the number of times ${\bf x}_{i}$ is selected in ${\bf X}_{l}^{\prime}$ Let $A_{l}=\max_{1\leq i\leq l}k_{i,l}$ and $B_{l}=\max_{l+1\leq i\leq K}k_{i,l}$.
+Let $k_{i,l}$ be the number of times ${\bf x}_{i}$ is selected in ${\bf X}_{l}^{\prime}$ .
+Let $A_{l}=\max_{1\leq i\leq l}k_{i,l}$ and $B_{l}=\max_{l+1\leq i\leq K}k_{i,l}$.
 Let $P(l)=\mathbb{P}(B_{l}>A_{l})$.
@@
 Therefore $P(l+1)=\mathbb{P}(B_{l+1}>A_{l+1})\leq\mathbb{P}(B_{l}>A_{l})=P(l)$.
-Appendix B Detailed Explanation of Assumption 4
+Appendix B Explanation of Assumption 4
 As illustrated in Figure 6, the dataset for $\mathbb{T}_{a}$ and the dataset for $\mathbb{T}_{b}$ demonstrate interchangeability with $\mathcal{G}_{a}^{(4)}$ and $\mathcal{G}_{b}^{(4)}$, respectively.
@@
 A similar process applies for the $\mathcal{G}_{b}^{(4)}$.
-Appendix C Explanation of TPGs Combination
-Figure 3 presents $\mathcal{G}^{(4)}_{a}$ and $\mathcal{G}^{(4)}_{b}$ as the TPGs corresponding to the last input token ${\bf e}_{4}$ for $\mathbb{T}_{a}$ and $\mathbb{T}_{b}$, respectively.
-In $\mathcal{G}^{(4)}_{a}$, the token priority is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}>{\bf e}_{4}$.
-In $\mathcal{G}^{(4)}_{ab}$ for the mixed-topics, the priority order is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{2}={\bf e}_{4}$.
-The strict equality ${\bf e}_{1}={\bf e}_{2}$ from $\mathcal{G}^{(4)}_{a}$ is maintained in $\mathcal{G}^{(4)}_{ab}$, but the strict inequality ${\bf e}_{2}>{\bf e}_{4}$ is replaced with ${\bf e}_{2}={\bf e}_{4}$ in mixed-topics, forming the new SCC, $\{{\bf e}_{1},{\bf e}_{2},{\bf e}_{4}\}$, in $\mathcal{G}^{(4)}_{ab}$.
-Appendix D Detailed Simulation Studies with Single-layer Self-attention
-D.1 Simulation Process
+Appendix C Detailed simulation studies with single-layer self-attention
+C.1 Simulation process
 Theoretical TPGs generation.
@@
 The model is trained using the SGD optimizer with a learning rate $\eta=0.01$ for $8000$ iterations.
-The training of attention weights is divided into two stages for each instance: (1) computing ${\bf W}^{\text{svm}}$ for each topic  [^4] Note: ${\bf W}^{\text{svm}}={\bf 0}$ means the number of SCCs is $1$ for $\mathcal{G}^{k},\forall k\in[K]$.
-During the simulation, we proceed to the next instance when ${\bf W}^{\text{svm}}={\bf 0}$ until reaching a total of $100$ instances.; (2) get $\bf{W}(\tau)$ at each iteration for each topic.
+The training of attention weights is divided into two stages for each instance: (1) computing ${\bf W}^{\text{svm}}$ for each topic; [^4] Note: ${\bf W}^{\text{svm}}={\bf 0}$ means the number of SCCs is $1$ for $\mathcal{G}^{k},\forall k\in[K]$.
+During the simulation, we proceed to the next instance when ${\bf W}^{\text{svm}}={\bf 0}$ until reaching a total of $100$ instances.
+(2) get $\bf{W}(\tau)$ at each iteration for each topic.
 In Stage (1), prior to using the CVXPY package to get ${\bf W}^{\text{svm}}$, SCCs are identified for each TPG derived from the using Tarjan’s algorithm.
@@
 To reduce the potential numerical issues in the outputs, $\mathbb{S}({\bf X}{\bf W}\bar{{\bf x}})$ is rounded to three decimals, ensuring that tokens within the same SCC yield consistent softmax outputs.
-Our code for all simulations is available in ( ).
-D.2 Additional Experiments to Support Theorem 3
+C.2 Additional experiments to support Theorem 2
 To further illustrate Theorem 2 we define the attention priority similarity of weights $\bf W$$^{\prime}$ relative to $\bf W$} for a sequence $\bf X$} as: $R_{{\bf W},{\bf W}^{\prime}}({\bf X})=$
@@
 Using these TPGs, we randomly generate $\text{DSET}_{a}$ and $\text{DSET}_{b}$.
-We compute $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$, $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ and $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ using the same procedure as Li et al., 2024b .
+We compute $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$, $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ and $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ using the same procedure as li2024mechanics.
 We generate test sequences $\bf Z$} within $\mathbb{T}_{a}$, and we calculate the attention priority similarity of $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to both $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$.
-We repeat this process for multiple TPGs and input sequences (simulation details in Appendix D).
+We repeat this process for multiple TPGs and input sequences (simulation details in Appendix C).
 Figure 7 clearly demonstrates that the similarity converges to 1 after $\tau=8000$ iterations when evaluated relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue line), but fails to converge relative to $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ (orange line).
@@
 Figure 7: Convergence of attention priority similarity for $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$ relative to $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$ (blue) and $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$ (orange).
-D.3 Simulation in Section 5
+C.3 Simulation in Section 5
 In Figure 5(a), we predict next tokens for 5000 test sequences from $\mathbb{T}_{a}$ with $T_{test}=\{4,8,16,24,32,64,128,256,512\}$, while fixing $L=4$, $d=16$, $T_{train}=4$, and $K=10$.
@@
 18 98.02 $\pm$ 1.58 1.26 $\pm$ 1.14 0.72 $\pm$ 0.86
-D.4 Additional Experiments for Convergence in Mixed-topics
+C.4 Additional experiments for convergence in mixed topics
 (a) $K=6,L=4$
@@
 Figure 8: Convergence of $\frac{{\bf W}_{a}(\tau)}{\|{\bf W}_{a}(\tau)\|_{F}}$(blue), $\frac{{\bf W}_{b}(\tau)}{\|{\bf W}_{b}(\tau)\|_{F}}$(orange), and $\frac{{\bf W}_{ab}(\tau)}{\|{\bf W}_{ab}(\tau)\|_{F}}$(green) for varying $K$ and $L$, with fixed $T_{train}=4$ and $d=16$.
-Building upon the convergence experiments in Li et al., 2024b , our work demonstrates that the correlation coefficients $\langle{\bf W}_{ab}(\tau),{\bf W}^{\text{svm}}_{ab}\rangle/\langle\|{\bf W}_{ab}(\tau)\|_{F},\|{\bf W}^{\text{svm}}_{ab}\|_{F}\rangle$ (green lines) in Figure 8, measured with varying $K=\{6,10,14\}$ and $L=\{8,12,16\}$, approach to 1.
-These results indicate that Theorem 1 extends beyond individual topics to also capture the convergence dynamics of mixed-topics scenarios, albeit with relatively slower convergence.
+Building upon the convergence experiments in li2024mechanics, our work demonstrates that the correlation coefficients $\langle{\bf W}_{ab}(\tau),{\bf W}^{\text{svm}}_{ab}\rangle/\langle\|{\bf W}_{ab}(\tau)\|_{F},\|{\bf W}^{\text{svm}}_{ab}\|_{F}\rangle$ (green lines) in Figure 8, measured with varying $K=\{6,10,14\}$ and $L=\{8,12,16\}$, approach to 1.
+These results indicate that Theorem 1 extends beyond individual topics to also capture the convergence in mixed-topic scenarios, albeit with relatively slower convergence.
 In these experiments, we fix $T_{train}=4$ and $d=16$.
@@
 The shaded area around each line represents the $95\%$ confidence interval, computed over 50 epochs.
-D.5 Numerical Analysis for Each Scenario in Figure 3
+C.5 Numerical analysis for each scenario in Figure 3
 Figure 9 provides a numerical breakdown for each scenario in Figure 3.
@@
 For each input sequence $\bf X$, we apply ${\bf W}_{a}(\tau)$ and ${\bf W}_{ab}(\tau)$ with $\tau=8000$ to predict the next token, yielding $\hat{y}_{{\bf W}_{a}}$ and $\hat{y}_{{\bf W}_{ab}}$, respectively.
-Let $[\mathbb{S}({\bf X}{\bf W}_{a}(\tau){\bar{\bf x}})]_{i}=a_{i}$ and $[\mathbb{S}({\bf X}{\bf W}_{ab}(\tau){\bar{\bf x}})]_{i}=b_{i}$, for $i\in[T]$.
-Following Equation 9 and Equation 10, we compute $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{x_{i}}$ and $[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{x_{i}}$ to get the highest probability SCC and predict the next token for each input sequence.
+Let $\mathbb{S}({\bf X}{\bf W}_{a}(\tau){\bar{\bf x}})]_{i}=a_{i}$ and $\mathbb{S}({\bf X}{\bf W}_{ab}(\tau){\bar{\bf x}})]_{i}=b_{i}$, for $i\in[T]$.
+Following Equation 9 and Equation 10, we compute $Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{x_{i}}$ and $Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{x_{i}}$ to get the highest probability SCC and predict the next token for each input sequence.
 (a) Topic continuity.
@@
 Based on $\mathcal{G}^{(4)}_{a}$ in Figure 3 (left), the priority order of these tokens is ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}>{\bf e}_{4}$, with corresponding $a_{i}$ values: $0.45>0.25>0.20>0.1$.
-Since $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.45$ is the largest, $\widehat{\mathcal{G}}_{a}^{(4)}=\{{\bf e}_{5}\}$ and $\hat{y}_{{\bf W}_{a}}={\bf e}_{5}$.
-In the mixed-topics scenario, ${\bf W}_{ab}$ preserves the attention priority but ${\bf e}_{4}$ and ${\bf e}_{1}$ have the same priority: ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{4}$, with corresponding $b_{i}$ values: $0.40>0.30>0.15=0.15$.
-Token ${\bf e}_{5}$ is still with the highest probability to be chosen, as $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.40$.
+Since $Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.45$ is the largest, $\widehat{\mathcal{G}}_{a}^{(4)}=\{{\bf e}_{5}\}$ and $\hat{y}_{{\bf W}_{a}}={\bf e}_{5}$.
+In the mixed-topic scenario, ${\bf W}_{ab}$ preserves the attention priority but ${\bf e}_{4}$ and ${\bf e}_{1}$ have the same priority: ${\bf e}_{5}>{\bf e}_{3}>{\bf e}_{1}={\bf e}_{4}$, with corresponding $b_{i}$ values: $0.40>0.30>0.15=0.15$.
+Token ${\bf e}_{5}$ is still with the highest probability to be chosen, as $Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{5}}=1\times 0.40$.
 Following the Definition 3, ${\bf W}_{ab}$ keeps topic for the the input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{1},{\bf e}_{3},{\bf e}_{4}]^{\top}$.
@@
 The corresponding values are $a_{1}=a_{3}=0.3$ and $a_{2}=a_{4}=0.2$.
-Then $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{1}}=2\times 0.30$ and $[Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.20$.
+Then $Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{1}}=2\times 0.30$ and $Cf_{{\bf W}_{a}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.20$.
 Thus, $\widehat{\mathcal{G}}_{a}^{(4)}$ is $\{{\bf e}_{5}\}$ with the highest probability.
 ${\bf W}_{ab}$ makes ${\bf e}_{4}$ and ${\bf e}_{1}$ with the same priority, as indicated by $\mathcal{G}^{(4)}_{ab}$ in Figure 3 (left).
-Both ${\bf e}_{1}$ and ${\bf e}_{4}$ are within the highest probability SCC, $\widehat{\mathcal{G}}_{ab}^{(4)}$, due to $[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{1}}=[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.25$.
+Both ${\bf e}_{1}$ and ${\bf e}_{4}$ are within the highest probability SCC, $\widehat{\mathcal{G}}_{ab}^{(4)}$, due to $Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{1}}=[Cf_{{\bf W}_{ab}(\tau)}({\bf X})]_{{\bf e}_{4}}=2\times 0.25$.
 Although $\hat{y}_{{\bf W}_{ab}}\not\in\widehat{\mathcal{G}}_{a}^{(4)}$, $\widehat{\mathcal{G}}_{a}^{(4)}\in\widehat{\mathcal{G}}_{ab}^{(4)}$.
@@
 In the input sequence ${\bf X}=[{\bf e}_{5},{\bf e}_{4},{\bf e}_{4},{\bf e}_{4}]^{\top}$, the lower-priority token ${\bf e}_{4}\not\in\dot{\mathcal{G}}^{(4)}_{ab}({\bf X})$ appears more frequently than the higher-priority token ${\bf e}_{5}\in\dot{\mathcal{G}}^{(4)}_{ab}({\bf X})$, illustrating our Theorem 3.
-Appendix E Experimental Details in Section 6
+Appendix D Experimental details in Section 6
 In this section, we provide the experimental details in four LLMs: GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3.
-Here, we outline the general procedure used in each model, under identical parameter settings, to generate continuations for each segment of the abstract as follows:
+Here, we outline the general procedure used in each model, under identical parameter settings, to generate continuations for each segment of the abstract.
   1. 1.
@@
   2. 2.
-    Randomly select 5 different papers from paper $A$, as papers in $\{B_{i}\}_{i=1}^{5}$.
+    Randomly select 5 papers different from paper A, as papers B in $\{\text{B}_{i}\}_{i=1}^{5}$.
   3. 3.
-    For the input segment $\bf{X}$, apply RAG to extract top 3 relevant excerpts (chunks) from paper $A$ as the the knowledge $A$, denoted as $\text{Ref}_{A}$. Each chunk is with 800 tokens.
+    For the input segment $\bf{X}$, apply RAG to extract top 3 relevant excerpts (chunks) from paper A as the the knowledge A, denoted as $\text{Ref}_{\text{A}}$. Each chunk has 800 tokens length.
   4. 4.
-    Similarly, retrieve top 3 relevant excerpts from paper $B_{i}$ as the knowledge $B_{i}$, denoted as $\text{Ref}_{B_{i}}$.
+    Similarly, retrieve top 3 relevant excerpts from paper $\text{B}_{i}$ as the knowledge $\text{B}_{i}$, denoted as $\text{Ref}_{\text{B}_{i}}$.
   5. 5.
-    Combine the knowledge from Topic $A$ and from Topic $B_{i}$ as the knowledge $AB_{i}$ for mixed-Topic, denoted as $\text{Ref}_{AB_{i}}$.
+    Combine the knowledge from Topic A and from Topic $\text{B}_{i}$ as the knowledge $\text{AB}_{i}$ for mixed Topics, denoted as $\text{Ref}_{\text{AB}_{i}}$.
   6. 6.
-    For the input segment $\bf{X}$, each LLM with the follow prompts, $\text{Prompt}_{A}$ and $\text{Prompt}_{AB_{i}}$, to generate the continuations as $\hat{y}_{\bf{W}_{a}}$ and $\hat{y}_{\bf{W}_{ab}}$, respectively.
-Notably, the only difference between $\text{Prompt}_{A}$ and $\text{Prompt}_{AB_{i}}$ is the reference excerpts provided $\text{Ref}_{A}$ or $\text{Ref}_{AB_{i}}$.
-All LLMs were queried with a temperature of 0 to match the greedy decoding in our theoretical framework.
+    For the input segment $\bf{X}$, promot each LLM with $\text{Prompt}_{\text{A}}$ and $\text{Prompt}_{\text{AB}_{i}}$, to generate the continuations as $\hat{y}_{\bf{W}_{a}}$ and $\hat{y}_{\bf{W}_{ab}}$, respectively.
+Notably, the only difference between $\text{Prompt}_{\text{A}}$ and $\text{Prompt}_{\text{AB}_{i}}$ is the reference excerpts provided $\text{Ref}_{\text{A}}$ or $\text{Ref}_{\text{AB}_{i}}$.
+All LLMs are set with a temperature of 0 to match the greedy decoding in our theoretical framework.
 The maximum completion length was set to 1000 tokens to ensure that the generated continuations could complete the abstract.
     1. (a)
-      $\text{Prompt}_{A}$:
-      Here are some relevant excerpts from research paper(s) as reference:$\text{Ref}_{A}$.
-Below is the 1st fragment of an abstract from arXiv paper $A$: $\bf{X}$.
-Please continue the 2nd fragment of the abstract based on the relevant excerpts without including the given content in the output.
+      $\text{Prompt}_{\text{A}}$:
+[…]
     2. (b)
-      $\text{Prompt}_{AB_{i}}$:
-      Here are some relevant excerpts from research paper(s) as reference:$\text{Ref}_{AB_{i}}$.
-Below is the 1st fragment of an abstract from arXiv paper $A$: $\bf{X}$.
-Please continue the 2nd fragment of the abstract based on the relevant excerpts without including the given content in the output.
+      $\text{Prompt}_{\text{AB}_{i}}$:
+[…]
   7. 7.
-    Calculate the average cosine similarity between $\hat{y}_{\bf{W}a}$ and $\hat{y}_{\bf{W}{ab}}$ across five pairs of paper $A$ and paper $B_{i}$.
-E.1 Impact of the Input Length
-To investigate the impact of the input length, we vary $T=\{10,30,50,70,90,110,130,150\}$ for every paper as Topic $A$, increasing the length of the input segment $\bf X$, as shown on the x-axis from Figure 5.
-E.2 Impact of the Topic Ambiguity
-We quantify topic (paper) ambiguity by the average similarity among each paper’s keywords.
-As arXiv papers do not provide keywords, we use Llama-3.3 to generate four keywords for each paper prior to generating continuations with the LLMs.
+    Calculate the average cosine similarity between $\hat{y}_{\bf{W}a}$ and $\hat{y}_{\bf{W}{ab}}$ across five pairs of paper A and paper $\text{B}_{i}$.
+D.1 Impact of input length
+To investigate the impact of the input length, we vary $T=\{10,30,50,70,90,110,130,150\}$ for every paper as Topic $A$, increasing the length of the input segment $\bf X$ extracted from the abstract of paper A, as shown on the x-axis from Figure 5.
+D.2 Impact of topic ambiguity
+We quantify topic (paper) ambiguity by computing the average similarity among each paper’s keywords.
+Since arXiv papers do not provide keywords, we use Llama-3.3 to generate four keywords for each paper prior to generating continuations with the LLMs.
 To investigate the topic ambiguity, we fix the input length with $T=80$ for every paper as Topic A and order papers by the average keywords similarity, as shown on the x-axis of Figure 5.
-Higher keywords similarity corresponds to lower topic ambiguity.
-Appendix F Computational Resources for Experiments
-In our simulations based on the single-layer self-attention model, each group of parameter setting requires 7 hours to train two models separately, one for single input topic and one for mixed-topic, followed by 2 additional hours for next-token prediction.
+A higher keywords similarity corresponds to lower topic ambiguity.
+D.3 Additional experiments
+D.3.1 Extended input length
+To further examine the impact of the input length, we randomly select 50 out of 100 arXiv papers introduced in Sec.
+6 and extend the input length with the first $310,460,\ldots,1210$ words from the introduction of each paper A as the input prompt.
+We start from $T=310$ because shorter introductions provide insufficient contextual information to reliably extract relevant excerpts from the full paper when using RAG.
+As shown in Figure 10, the average cosine similarity generally increases with the more introduction content from paper A, with the exception of DeepSeek, which exhibits only a marginal improvement.
+Figure 10: Extended length of input sequences.
+D.3.2 Alternative measure of ambiguity
+As an alternative approach to validate our results, we also measure the ambiguity based on the similarity between the keywords of paper A and those of paper B.
+Following the same setup, we rank the cosine similarity for each paper A by the increasing similarity between its keywords and those of five papers B, where a higher level corresponds to a greater ambiguity.
+As shown in Figure 11, our results still hold since the cosine similarity remains relatively stable as the ambiguity level increases.
+Figure 11: Topic ambiguity measured by the similarity between the keywords of paper A and those of paper B. Higher keywords similarity indicate greater topic ambiguity between the two papers.
+Appendix E Computational resources for experiments
+In our simulations based on the single-layer self-attention model, each group of parameter setting requires 7 hours to train two models separately, one for single input topic and one for mixed topics, followed by 2 additional hours for next-token prediction.
 In our experiments on LLMs, we query GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3 through API calls.
 All experiments were conducted on a standard laptop without specialized hardware.
-For each LLM, the full process, including selecting relevant excerpts using RAG and generating continuations, requires approximately 20 hours of runtime, with a total of 13 million input tokens and 2.5 million output tokens.
-The total API usage cost for the experiments is approximately 200 USD.
-Appendix G Impact Statement
+For each LLM, the full process, including selecting relevant excerpts using RAG and generating continuations, requires approximately 80 hours of runtime, with a total of 30 million input tokens and 5 million output tokens.
+The total API usage cost for the experiments is approximately 350 USD.
+Appendix F Impact statement
 Our investigation highlights fundamental differences between spontaneous topic changes in LLMs and spontaneous human thought, informing the development of more natural and flexible AI systems in domains such as customer service and mental health support.
@@
 There are many other potential societal consequences of our work, none which we feel must be specifically highlighted here.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: The fifth paragraph and Subsection 1.1 in Introduction accurately reflect the paper’s contributions, scope, assumptions and limitations.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9.
+Justification: We introduce assumptions in Introduction and we have a subsection of limitations in Section 8, including the lack of theoretical framework for complex architectures, alternative cost functions, and diverse training objectives.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14.
+Justification: Building on Assumptions 1 and 2 from Theorem 1 in li2024mechanics, and incorporating our new Assumptions 3 and 4, we present Theorems 2, 3, and 4.
+All proofs are provided in Appendix A.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: We provide full disclosure of the implementation details to reproduce experimental results.
+Section 5 and Appendix C present the simulation details and results based on the single-layer self-attention model, while the experiments on modern LLMs are detailed in Section 6 and Appendix D.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: All code and data for reproducibility is provided as supplementary material. Upon acceptance we will also provide a public GitHub repository.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29.
+Justification: We fully disclose the experimental setup in Section 5 and the corresponding details in Appendix C.1 for the single-layer self-attention model.
+The experimental setup and details for modern LLMs are provided in Section 6 and Appendix D, respectively.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34.
+Justification: Figure 5 shows the $95\%$ intervals of the average cosine similarity as the input length increases.
+In Appendix C, Table 1 and Table 2 report the $95\%$ intervals for the percentage of each scenario.
+Additionally, Figure 7 shows the $95\%$ confidence interval for the attention priority similarity, and Figure 8 presents the convergence of $\bf W$ with corresponding $95\%$ confidence intervals.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Appendix E provides details on the computational resources used in our experiments, including execution time, API usage costs, and the total number of input and output tokens.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: Our work complies with the NeurIPS Code of Ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: In Appendix E, we discuss the social impact of our work from both potential positive and negative impacts.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: This paper poses no such risks.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59.
+Justification: Our work is conducted in Python with several open-source Python libraries.
+For generating continuations in our experiments, we access the following publicly released LLMs via API: GPT-4o[hurst2024gpt4o], Llama-3.3[grattafiori2024llama3], Claude-3.7[claude37], and DeepSeek-V3[liu2024deepseekv3].
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: We provide a detailed description of the dataset used for the single-layer self-attention model in Appendix C.1 and the dataset used for the modern LLMs in Appendix D.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: This work does not involve crowdsourcing experiments or research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: Our work doesn’t involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [Yes]
+  79.
+Justification: We use LLM APIs, including GPT-4o, Llama-3.3, Claude-3.7, and DeepSeek-V3, to generate the continuations of the input segment.
+Section 6 and Appendix D provide all experimental results and detailed process related to the frontier LLMs.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.